### v2

# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

Abstract
Using the semiformal method in combinatorics we generalize Pólya’s theorem. This theorem determines the limiting probability that walks in the grid graph $\mathbb{Z}^{d}$ visit the given vertex. We generalize it to the countable complete graph $K_{\mathbb{N}}$ with edge weights in $\mathbb{C}$. In part II we treat edge weights in $\mathbb{C}[[x_{1},\dots,x_{k}]]$.
1 Introduction
Let $d,n\in\mathbb{N}$ ($\{1,2,\dots\}$) and $\overline{v}\in\mathbb{Z}^{d}$, where $\mathbb{Z}=\{\dots,-1,0,1,\dots\}$ are the integers. Let $P_{d}(\overline{v},n)$ be the probability that a walk of length $n$ in the grid graph $\mathbb{Z}^{d}$ starting at $\overline{0}$ visits in a step $i>0$ the vertex $\overline{v}$. In 1921, G. Pólya proved in [23] a theorem, which we generalize, asserting that for every vertex $\overline{v}$,
$\lim_{n\to\infty}P_{d}(\overline{v},\,n)\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ <1&\dots&\text{if $d\geq 3$}\,.\end{array}\right$.
Pólya’s theorem is almost always cited in the special case with $\overline{v}=\overline{0}$, although G. Pólya proved it for any vertex $\overline{v}$ (see [15, Section 1] for a translated quote from [23]). A selection of articles and books discussing Pólya’s theorem is [3, 4, 6, 9, 11, 13, 17, 18, 19, 20, 24, 25, 28, 29].
We generalize it as follows. Let
$\mathbb{N}_{2}=\binom{\mathbb{N}}{2}=\{e\colon\;e\subset\mathbb{N}\wedge|e|=2\}$}
be the edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$ on the vertex set $\mathbb{N}$. For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1.\dots\}$) its cardinality. Let
$h\colon\mathbb{N}_{2}\to\mathbb{C}$
be a complex edge weight. $\mathbb{C}$ denotes the field of complex numbers, which is endowed with the usual Archimedean absolute value $|z|=\sqrt{z\cdot\overline{z}}$. For $n\in\mathbb{N}_{0}$ let $W(n)$ be the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at the vertex $1$. Thus $W(0)=\{\langle 1\rangle\}$ and for $n>0$ the set $W(n)$ is infinite and countable. For $v\in\mathbb{N}$ let $W(v,n)$ be the subset of $W(n)$ of the walks that visit the vertex $v$ in a step $i>0$. We extend the weight in the usual multiplicative way to
$h\colon\bigcup_{n=0}^{\infty}W(n)\to\mathbb{C}\,.$}
Let $W\subset\bigcup_{n=0}^{\infty}W(n)$. If the series
$\sum_{w\in W}h(w)$}
absolutely converges, we denote its sum by $h(W)$. (Results on series are reviewed in the next section.) Suppose that for every $n\in\mathbb{N}_{0}$ the series $\sum_{w\in W(n)}h(w)$ absolutely converges. Then so does $\sum_{w\in W(v,n)}h(w)$ and we consider the generating function
$A_{h,v}(x)=\sum_{n\geq 0}h(W(v,\,n))x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
Our generalizations of Pólya’s theorem are formulas for the quantities
$A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))\,.$}
We treat this series both in conditional and absolute convergence. Our formulas appear in fourteen theorems: Theorems 3.23.7 in Section 3 and Theorems 4.24.9 in Section 4. Pólya’s theorem is subsumed in five of them: Theorems 3.5 and 3.7 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}=\overline{0}$, and Theorems 4.6 (4.7) and 4.9 cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
Why fourteen generalizations of one theorem? For three reasons. First, in Pólya’s theorem weights are nonnegative real numbers and convergence of infinite series is automatically absolute. For complex weights we consider both absolute, and conditional convergence. We also consider, for nonnegative real weights, series with sum $\infty$. Second, Pólya’s theorem lives in the probabilistic setup where weights of edges incident to a fixed vertex sum up to $1$. We consider two cases—when this condition is kept (convex weights) and when it is dropped. Third, already Pólya’s theorem is in reality two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in the complexity of proofs. In our generalization we correspondingly distinguish cases $v=1$ and $v\neq 1$. Now $3\cdot 2\cdot 2=12$ and two more generalizations arise by variations of assumptions.
We explain how our approach includes Pólya’s original problem. If $f\colon A\to B$ is a map and $C$ is any set (not necessarily a subset of $A$), we call the set
$f[C]=\{f(x)\colon\;x\in A\cap C\}\ \ (\subset B)$
the image of $C$ by $f$. Let $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$. We take any bijection
$f\colon\mathbb{Z}^{d}\to\mathbb{N}$
such that $f(\overline{0})=1$, and set $v=f(\overline{v})$. The grid graph on $\mathbb{Z}^{d}$, i.e. the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
$\{\overline{a},\,\overline{b}\}\in E_{d}\iff\sum_{i=1}^{d}|a_{i}-b_{i}|=1\,$,}
determines the weight $h\colon\mathbb{N}_{2}\to\{\frac{1}{2d},0\}$ by the relation $h(e)\neq 0$ iff $e=f[e^{\prime}]$ for some $e^{\prime}\in E_{d}$. Then for every $n\in\mathbb{N}_{0}$ we have
$P_{d}(\overline{v},\,n)=h(W(v,\,n))\,.$
We structure our article as follows. In Section 2 we review infinite series with complex summands and generating functions. Sections 3 and 4 contain our main results, the mentioned fourteen generalizations of Pólya’s theorem. Section 5 contains concluding remarks. We continue our investigation in [16] where we consider edge weights
$h\colon\mathbb{N}_{2}\to\mathbb{C}[[x_{1},\,\dots,\,x_{k}]]\,.$
The domain of formal power series $\mathbb{C}[[x_{1},\dots,x_{k}]]$ is endowed with the standard non-Archimedean norm.
In Sections 3 and 4 we use what we term the semiformal method in (enumerative) combinatorics. It extends the symbolic method in combinatorics [2, 8, 10, 12, 21, 22, 26] from finite to countable sets. In our situation such extension is natural and inevitable, coefficients in generating functions $A_{h,v}(x)$ arise as sums of infinite series $\sum_{w\in W(v,n)}h(w)$. We develop the semiformal method more in [14]. Why “semiformal”? Generating “functions” in the symbolic method are formal power series with coefficients determined by finite operations. Our extension allows operations with coefficients based on limit transitions.
2 Series and generating functions
We review series and generating functions. A series is a map
$h\colon X\to\mathbb{C}$
defined on an at most countable set $X$. We write it as $\sum_{x\in X}h(x)$. We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
  1. 1.
    For some $c>0$ and every finite set $Y\subset X$ we have $\sum_{x\in Y}|h(x)|\leq c$.
  2. 2.
    If $X$ is countable, then for every bijection $f\colon\mathbb{N}\to X$, the limit
    $s=\lim_{n\to\infty}{\textstyle\sum_{i=1}^{n}h(f(i))}$
    exists and does not depend on $f$.
We call $s$ the sum of the series and denote it again by $\sum_{x\in X}h(x)$. If the set $X=\{x_{1},x_{2},\dots,x_{n}\}\neq\emptyset$ is finite, then every series $\sum_{x\in X}h(x)$ absolutely converges and has the sum
$h(x_{1})+h(x_{2})+\dots+h(x_{n})\,.$
For $X=\emptyset$ we define the sum as 0. If
$U(x)=\sum_{n\geq 0}u_{n}x^{n}\ \ (\in\mathbb{C}[[x]])$}
is a formal power series, so that $x$ is a formal variable, we denote for any $x\in\mathbb{C}$ by $F_{U}(x)$ the sum of the power series $\sum_{n\geq 0}u_{n}x^{n}$, if it absolutely converges.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. A walk $w$ in $K_{\mathbb{N}}$ is any $(n+1)$-tuple $w$ of vertices $v_{i}\in\mathbb{N}$,
$w=\langle v_{0},\,v_{1},\,\dots,\,v_{n}\rangle\,$,
such that $n\in\mathbb{N}_{0}$ and $v_{i-1}\neq v_{i}$ for every $i\in[n]$ ($\{1,2,\dots,n\}$, $0]=\emptyset$). The length $n$ of $w$ is denoted by $|w|$ ($\in\mathbb{N}_{0}$). We extend $h$ to walks. For $n>0$ we set
$h(w)=\prod_{i=1}^{n}h(\{v_{i-1},\,v_{i}\})\,$,}
and for $n=0$ we define $h(w)=1$. Recall that $W(n)$ is the set of walks $w$ in $K_{\mathbb{N}}$ such that $v_{0}=1$ and $|w|=n$.
Definition 2.1 (lightness)
A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is light if for every $n\in\mathbb{N}_{0}$ the series
$\sum_{w\in W(n)}h(w)$}
absolutely converges.
Let $S=\sum_{x\in X}h(x)$ be a series. If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$. Any subseries of an absolutely convergent series absolutely converges, but we cannot compare sizes of sums
$\big|\sum_{x\in Y}h(x)\big|\,\text{ and }\,\big|\sum_{x\in X}h(x)\big|\,.$}
Except when $h\colon X\to\mathbb{R}_{\geq 0}$. If the series absolutely converges and $Y\subset X$, then the sums satisfy
$0\leq\sum_{x\in Y}h(x)\leq\sum_{x\in X}h(x)\,.$}
If $R=\sum_{x\in X}g(x)$ and $S=\sum_{x\in X}h(x)$ are series and $\alpha,\beta\in\mathbb{C}$, the series
$\alpha R+\beta S=\sum_{x\in X}(\alpha g(x)+\beta h(x))$}
is their linear combination. We omit the easy proof of the next result.
Proposition 2.2
If $R$ and $S$ absolutely converge and have respective sums $r$ and $s$, then series $\alpha R+\beta S$ absolutely converges and has the sum $\alpha r+\beta s$.
The next approximation lemma is a key tool.
Lemma 2.3
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, then for every $\varepsilon>0$ there exists a finite set $Y\subset X$, denoted by $Y(S,\,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$ we have
$\big|s-\sum_{x\in Z}h(x)\big|\leq\varepsilon\,.$}
Proof. Let an $\varepsilon>0$ be given. For finite $X$ we set $Y=X$. For countable $X$ we take any bijection $f\colon\mathbb{N}\to X$, take an $N\in\mathbb{N}$ such that
$|\sum_{n=1}^{N}h(f(n))-s|\leq\frac{\varepsilon}{2}\,\text{ and }\,\sum_{n>N}|h(f(n))|\leq\frac{\varepsilon}{2}$}
and set Y=f[[N]]Y=f[\,[N]\,]. Then for every finite set $Z$ with $Y\subset Z\subset X$ we have
$|s-\sum_{x\in Z}h(x)|\leq|s-\sum_{x\in Y}h(x)|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$}
$\Box$
Two more operations with series are grouping and product. A partition of a set $X$ is a set $P$ of nonempty and disjoint sets such that $\bigcup P=X$. If $X$ is at most countable, then so is $P$ and every set $Z\in P$. If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every set $Z\in P$ the subseries $\sum_{x\in Z}h(x)$ absolutely converges and has sum $s_{Z}$, then the series
$S_{P}=\sum_{Z\in P}s_{Z}$}
is called the grouping of $S$.
Proposition 2.4
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, and $P$ is a partition of $X$, then $S_{P}$ is correctly defined, absolutely converges and has the same sum $s$.
Proof. For every $Z\in P$ the series $S_{Z}=\sum_{x\in Z}h(x)$ is a subseries of $S$ and absolutely converges. Thus the series $S_{P}$ is correctly defined and we show that it absolutely converges. Let
$c=\sup(\{\sum_{x\in Z}|h(x)|\colon\;\text{$Z\subset X$ and $Z$ is finite}\})\ \ (<+\infty)\,.$}
Let $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. We use Lemma 2.3 and take for each $i\in[n]$ the finite set
$Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i})\ \ (\subset Z_{i})\,.$
We form the disjoint union $Z_{0}=Z_{1}^{\prime}\cup\dots\cup Z_{n}^{\prime}$ ($\subset X$). Then
${\textstyle\sum_{i=1}^{n}\big|\sum_{x\in Z_{i}}h(x)\big|}$ $\leq$ ${\textstyle\sum_{i=1}^{n}\big|\sum_{x\in Z_{i}}h(x)-\sum_{x\in Z_{i}^{\prime}}h(x)\big|\,+}$
$+$ ${\textstyle\sum_{x\in Z_{0}}|h(x)|\leq\sum_{i=1}^{n}2^{-i}+c\leq 1+c}$
and $S_{P}$ absolutely converges.
Let $t$ be the sum of $S_{P}$. We show that $|t-s|\leq\varepsilon$ for every $\varepsilon>0$; then $t=s$. Let an $\varepsilon>0$ be given. We use Lemma 2.3 and take finite sets
$X^{\prime}=Y(S,\,\frac{\varepsilon}{3})\ \ (\subset X)\,\text{ and }\,P^{\prime}=Y(S_{P},\,\frac{\varepsilon}{3})\ \ (\subset P)\,.$}
We take a finite set $\{Z_{1},\dots,Z_{n}\}\subset P$ such that
$P^{\prime}\subset\{Z_{1},\,\dots,\,Z_{n}\}\,\text{ and }\,X^{\prime}\subset\bigcup_{i=1}^{n}Z_{i}\,.$}
We again use Lemma 2.3 and take for $i\in[n]$ finite sets
$Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i}\cdot\frac{\varepsilon}{3})\ \ (\subset Z_{i})\,.$}
For $i\in[n]$ let
$Z_{i}^{\prime\prime}=Z_{i}^{\prime}\cup(X^{\prime}\cap Z_{i})\ \ (\subset Z_{i})\,.$
We form the disjoint union
$X_{0}=\bigcup_{i=1}^{n}Z_{i}^{\prime\prime}\ \ (\subset X)\,.$}
Then $X_{0}$ is finite and $X^{\prime}\subset X_{0}$. Also, $Z_{i}^{\prime}\subset Z_{i}^{\prime\prime}$ for every $i\in[n]$. We have
$|s-t|$ $\leq$ ${\textstyle|s-\sum_{x\in X_{0}}h(x)|+\sum_{i=1}^{n}|\sum_{x\in Z_{i}^{\prime\prime}}h(x)-\sum_{x\in Z_{i}}h(x)|\,+}$
$+$ ${\textstyle|\sum_{i=1}^{n}\sum_{x\in Z_{i}}h(x)-t|\leq\frac{\varepsilon}{3}+\sum_{i=1}^{n}2^{-i}\cdot\frac{\varepsilon}{3}+\frac{\varepsilon}{3}\leq\varepsilon\,.}$
$\Box$
Let $A$ and $B$ be sets. We denote their ordered pair by $\langle A,\,B\rangle$. If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, are series, then the series
$S_{1}\cdot S_{2}=\sum_{\langle x,\,y\rangle\in X_{1}\times X_{2}}h_{1}(x)h_{2}(y)$}
is the product of $S_{1}$ and $S_{2}$.
Proposition 2.5
If $S_{1}$ and $S_{2}$ absolutely converge and have sums $s_{1}$ and $s_{2}$, respectively, then $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s_{1}s_{2}$.
Proof. We show that $S_{1}\cdot S_{2}$ absolutely converges. Let $c\geq 0$ be such that for any finite sets $Y_{i}$ with $Y_{i}\subset X_{i}$ and $i=1,2$,
$\sum_{x\in Y_{i}}|h_{i}(x)|\leq c\,.$}
Let $A\subset X_{1}\times X_{2}$ be a finite set. We take any finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
$\sum_{\langle x,\,y\rangle\in A}|h_{1}(x)h_{2}(y)|\leq\sum_{x\in Y_{1}}|h_{1}(x)|\cdot\sum_{y\in Y_{2}}|h_{2}(y)|\leq c\cdot c=c^{2}\,.$}
We show that the sum $s$ of $S_{1}\cdot S_{1}$ equals $s_{1}s_{2}$. It suffices to prove that for any $\varepsilon\in(0,1)$ we have $|s-s_{1}s_{2}|\leq\varepsilon$. So let $\varepsilon\in(0,1)$. Using Lemma 2.3 we take finite sets
$X^{\prime}_{1}=Y(S_{1},\,\frac{\varepsilon}{4(1+|s_{2}|)})\ \ (\subset X_{1}),\ \ X^{\prime}_{2}=Y(S_{2},\,\frac{\varepsilon}{4(1+|s_{1}|)})\ \ (\subset X_{2})\$}
and
$Z=Y(S_{1}\cdot S_{2},\frac{\varepsilon}{4})\ \ (\subset X_{1}\times X_{2})\,.$}
We take finite sets $X_{1}^{\prime\prime}$ and $X_{2}^{\prime\prime}$ such that $X_{1}^{\prime}\subset X_{1}^{\prime\prime}\subset X_{1}$, $X_{2}^{\prime}\subset X_{2}^{\prime\prime}\subset X_{2}$ and $Z\subset X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}$. Then $|s-s_{1}s_{2}|$ is at most
${\textstyle\big|s-\sum_{\langle x,\,y\rangle\in X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}}h_{1}(x)h_{2}(y)\big|+\big|\sum_{x\in X_{1}^{\prime\prime}}h_{1}(x)\sum_{y\in X_{2}^{\prime\prime}}h_{2}(y)-s_{1}s_{2}\big|}$
${\textstyle\leq\frac{\varepsilon}{4}+|(s_{1}+t_{1})(s_{2}+t_{2})-s_{1}s_{2}|\,,}$
where
[…]
Hence $|s-s_{1}s_{2}|\leq\frac{\varepsilon}{4}+|s_{1}t_{2}|+|s_{2}t_{1}|+|t_{1}t_{2}|\leq\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}=\varepsilon$. $\Box$
Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2,\dots,k$, be series. We define more generally their product as the series
$\prod_{i=1}^{k}S_{i}=\sum_{\overline{x}\in\overline{X}}h_{1}(x_{1})h_{2}(x_{2})\dots h_{k}(x_{k})\,$,}
where $\overline{x}=\langle x_{1},\dots,x_{k}\rangle$ is an ordered $k$-tuple of $x_{1}$, $\dots$, $x_{k}$ and $\overline{X}=X_{1}\times\dots\times X_{k}$. Iterating Proposition 2.5 we generalize it as follows.
Proposition 2.6
If the series $X_{1}$, $\dots$, $X_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has the sum $s_{1}s_{2}\dots s_{k}$.
Interestingly, a reversal of Proposition 2.5 holds.
Corollary 2.7
If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$
Proof. This follows from Proposition 2.5 if we show that $S_{1}$ and $S_{2}$ absolutely converge. Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$. $i=1,2$, and let $c>0$ be such that for every finite set $X\subset X_{1}\times X_{2}$,
$\sum_{\langle x,y\rangle\in X}|h_{1}(x)h_{2}(y)|\leq c\,.$}
We prove absolute convergence of $S_{1}$. We take any $y\in X_{2}$ with $h_{2}(y)\neq 0$. Then for every finite set $X\subset X_{1}$,
$\sum_{x\in X}|h_{1}(x)|=\frac{1}{|h_{2}(y)|}\sum_{\langle x,y\rangle\in X\times\{y\}}|h_{1}(x)h_{2}(y)|\leq\frac{c}{|h_{2}(y)|}\,.$}
Thus $S_{1}$ absolutely converges. For $S_{2}$ we proceed similarly. $\Box$
It is clear that the assumption on $S_{1}$ and $S_{2}$ cannot be omitted.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$. We obtain a sufficient condition for the lightness of $h$ in terms of the values $h(e)$. We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ that can be reached from $1$ by a walk of length at most $n$ that has edges with nonzero weights.
Definition 2.8 (slimness)
A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is slim if for every $n\in\mathbb{N}_{0}$ there is a constant $c=c(n)>0$ such that for every vertex $u\in V(h,n)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
$\sum_{v\in X}|h(\{u,\,v\})|\leq c\,.$}
Proposition 2.9
Every slim weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is light.
Proof. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be slim. We show by induction on $n\in\mathbb{N}_{0}$ that the series
$\sum_{w\in W(n)}h(w)$}
absolutely converges. For $n=0$ it is trivial because $W(0)=\{\langle 1\rangle\}$. Let $n>0$ and let $c>0$ be a constant such that for every finite set $X\subset W(n-1)$ we have
$\sum_{w\in X}|h(w)|\leq c$}
and that for every $u\in V(h,n-1)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
$\sum_{v\in X}|h(\{u,\,v\})|\leq c\,.$}
Let $X\subset W(n)$ be finite. We may assume that walks in $X$ have only edges with nonzero weights. We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in W(n-1)$ and $\ell(w)$ is the last vertex of $w$. Let $Y$ be the set of walks $w^{\prime}$. For $w^{\prime}\in Y$ let
$X(w^{\prime})=\{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}\colon\;w^{\prime}v\in X\}\,.$
Using finite versions of Propositions 2.4 and 2.2 we get the bound
${\textstyle\sum_{w\in X}|h(w)|}$ $=$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot\sum_{v\in X(w^{\prime})}|h(\{\ell(w^{\prime}),\,v\})|}$
$\leq$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot c\leq c^{2}\,.}$
$\Box$
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We define
$V(h)=\bigcup_{n\geq 0}V(h,\,n)\,.$}
It is the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk formed by edges with nonzero weights. We define the graph
$G(h)=\langle V(h),\,E(h)\rangle$
as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. It follows that if $G(h)$ has finite degrees then $h$ is slim.
Definition 2.10 (convexity)
A light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is convex if for every vertex $u\in V(h)$ the sum
$\sum_{e\in E(h),\,e\ni u}h(e)=\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,\,v\})=1\,.$}
We show that these sums exist.
Proposition 2.11
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be light and $u\in V(h)$. Then the series
$\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})$}
absolutely converges.
Proof. There is a walk $w_{0}=\langle u_{0},\dots,u_{n}\rangle$ in $G(h)$ such that $u_{0}=1$ and $u_{n}=u$. Let
$c=h(w_{0})=\prod_{i=1}^{n}h(\{u_{i-1},u_{i}\})\ \ (\neq 0)\,.$}
Let $W$ be the set of walks $\langle u_{0},\dots,u_{n},u_{n+1}\rangle$ in $K_{\mathbb{N}}$. Thus vertices $u_{0}$, $\dots$, $u_{n}$ are fixed and $u_{n+1}$ runs in $\mathbb{N}\setminus\{u_{n}\}=\mathbb{N}\setminus\{u\}$. The series $\sum_{w\in W}h(w)$ is a subseries of $\sum_{w\in W(n+1)}h(w)$ and therefore absolutely converges. By Proposition 2.2,
$\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})=\frac{1}{c}\sum_{w\in W}h(w)$}
absolutely converges. $\Box$
We survey the results on generating functions, which we use in the next two sections. We work with univariate formal power series
[…]
They form an integral domain denoted by $\mathbb{C}[[x]]$. For $U(x)$ and $V(x)$ in $\mathbb{C}[[x]]$,
$U(x)+V(x)=\sum_{n\geq 0}(u_{n}+v_{n})x^{n},\ U(x)\cdot V(x)=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$, and the units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with $a_{0}\neq 0$.
Definition 2.12 ($U(1)_{1}$ and $U(1)_{2}$)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ be in $\mathbb{C}[[x]]$. We introduce the following terminology and objects.
  1. 1.
    If the limit
    $U(1)_{1}=\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}u_{j}\ \ (\in\mathbb{C})}$
    exists, we say that $U(1)_{1}$ exists.
  2. 2.
    If the series
    $\sum_{n\geq 0}u_{n}1^{n}=\sum_{n\geq 0}u_{n}$}
    absolutely converges, we denote the sum by $U(1)_{2}$ and say that $U(1)_{2}$ exists.
If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$, but the opposite implication is, in general, not true. For example, for
$U(x)=\sum_{n\geq 1}(-1)^{n-1}\frac{1}{n}x^{n}$}
$U(1)_{1}$ exists and $U(1)_{1}=\log 2$, but $U(1)_{2}$ does not exist. But if
$U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]$}
and $U(1)_{1}$ exists, then $U(1)_{2}$ exists and $U(1)_{1}=U(1)_{2}$. The following result is well known.
Proposition 2.13
If $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ and $U(1)_{1}$ exists, then for every number $x\in\mathbb{C}$ with $|x|<1$ the power series
$\sum_{n\geq 0}u_{n}x^{n}$}
absolutely converges.
Definition 2.14 ($U(1)_{1}=+\infty$)
If $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ is in $\mathbb{R}[[x]]$ and
$\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}u_{j}=+\infty\,,}$
we say that $U(1)_{1}=+\infty$.
Proposition 2.15
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$. If $V(1)_{1}$ exists and $U(x)=V(x)\frac{1}{1-x}$, then
$\lim_{n\to\infty}u_{n}=V(1)_{1}\,.$
Proof. Since $\frac{1}{1-x}=1+x+x^{2}+\dots$, we have $u_{n}=v_{0}+v_{1}+\dots+v_{n}$. $\Box$
Quantities $U(1)_{1}$ are handled by means of the classical theorem of Abel, see [1] and [27, Chapter II.7]. We state it for the radius of convergence $R\geq 1$.
Theorem 2.16 (Abel’s 1)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. If $U(1)_{1}$ exists then the function
$F_{U}(x)=\sum_{n\geq 0}u_{n}x^{n}\colon\{x\in\mathbb{C}\colon\;|x|<1\}\to\mathbb{C}\,$,}
see Proposition 2.13, has the limit
$\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{U}(x)=U(1)_{1}\,.$
The following folklore result is in fact another variant of Abel’s theorem, incomparable with Theorem 2.16. We omit the easy proof.
Theorem 2.17 (Abel’s 2)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]$. If $U(x)$ absolutely converges for every $x\in[0,1)$, then the equality
$\lim_{x\to 1}F_{U}(x)=\lim_{n\to\infty}\sum_{j=0}^{n}u_{j}\ \ (\in[0,\,+\infty)\cup\{+\infty\})$}
always holds.
Compared to Theorem 2.16, the new case is when both limits equal $\infty$.
Proposition 2.18
Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{1}$ and $W(1)_{1}$ exist, then $U(1)_{1}$ exists and
$U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}\,.$
Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}u_{n}x^{n}$. Then
$\sum_{j=0}^{n}u_{j}=\sum_{j=0}^{n}(\alpha v_{j}+\beta w_{j})=\alpha\sum_{j=0}^{n}v_{j}+\beta\sum_{j=0}^{n}w_{j}\,.$}
By the limit transition $n\to\infty$ we get that that $U(1)_{1}$ exists and $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$. $\Box$
Iterating this result we easily generalize it to $k$-term linear combinations.
Proposition 2.19
Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$ for $n\in\mathbb{N}$. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
$U(1)_{1}=\prod_{i=1}^{n}V_{i}(1)_{1}\,.$}
Proof. By Propositions 2.4, 2.6 and 2.13, and the assumptions,
$\text{$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}$ with $|x|<1$}\,.$}
Theorem 2.16 and properties of limits of functions give
$U(1)_{1}$ $=$ $\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{U}(x)=\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}{\textstyle\prod_{i=1}^{n}F_{V_{i}}(x)}$
$=$ ${\textstyle\prod_{i=1}^{n}\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{V_{i}}(x)=\prod_{i=1}^{n}V_{i}(1)_{1}}\,.$
$\Box$
Since we have to assume that $U(1)_{1}$ exists, the case $n>2$ does not follow by iterating the case $n=2$. The example in [7] shows that the assumption of existence of $U(1)_{1}$ cannot be omitted.
Proposition 2.20
Let $U(x),V(x)\in\mathbb{C}[[x]]$ and $V(0)\neq 0$. If $U(x)=\frac{1}{V(x)}$ and if $U(1)_{1}$ and $V(1)_{1}$ exist, then $V(1)_{1}\neq 0$ and
$U(1)_{1}=\frac{1}{V(1)_{1}}\,.$}
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. Proposition 2.19 gives that $U(1)_{1}V(1)_{1}=1$. Thus $V(1)_{1}\neq 0$ and division by $V(1)_{1}$ yields the stated formula. $\Box$
Proposition 2.21
Let $U(x),V(x)\in\mathbb{R}_{\geq 0}[[x]]$ and $V(0)\neq 0$. If $\frac{1}{V(x)}, U(x)U(x)$ and $V(x)$ absolutely converge for $x\in[0,1)$, and if $V(1)_{1}=+\infty$, then
$U(1)_{1}=0\,.$
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. By Propositions 2.2, 2.4, 2.5 and 2.13, $F_{U}(x)F_{V}(x)=1$ for $x\in[0,1)$. Thus $F_{V}(x)\neq 0$ and $F_{U}(x)=\frac{1}{F_{V}(x)}$ for every $x\in[0,\,1)$. Two applications of Theorem 2.17 and properties of limits of functions give
$U(1)_{1}$ $=$ ${\textstyle\lim_{x\to 1}F_{U}(x)=\lim_{x\to 1}\frac{1}{F_{V}(x)}=\frac{1}{\lim_{x\to 1}F_{V}(x)}}$
$=$ ${\textstyle\frac{1}{\lim_{n\to\infty\sum_{j=0}^{n}v_{j}}}=\frac{1}{+\infty}=0\,.}$
$\Box$
Proposition 2.22
Let $U(x),V(x),W(x)\in\mathbb{C}[[x]]$ and $W(0)\neq 0$. If $U(x)^{2}=1-\frac{V(x)}{W(x)}$, if $U(1)_{1}$, $V(1)_{1}$ and $W(1)_{1}$ exist, and if $W(1)_{1}\neq 0$, then
$\big(U(1)_{1}\big)^{2}=1-\frac{V(1)_{1}}{W(1)_{1}}\,.$}
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Let T(x)=W(x)V(x)T(x)=W(x)-V(x). By Proposition 2.18, the sum $T(1)_{1}$ exists and $T(1)_{1}=W(1)_{1}-V(1)_{1}$. By Proposition 2.19 we have $(U(1)_{1})^{2}\cdot W(1)_{1}=T(1)_{1}=W(1)_{1}-V(1)_{1}$. Division by $W(1)_{1}$ yields the stated formula. $\Box$
Proposition 2.23
Let $U(x),V(x),W(x)\in\mathbb{R}_{\geq 0}[[x]]$ and $V(0),W(0)>0$. If $U(x)^{2}=1-\frac{V(x)}{W(x)}$, if $U(x)$, $V(x)$ and $W(x)$ absolutely converge for $x\in[0,1)$, if $F_{V}(x)$ is bounded on $[0,1)$ and if $W(1)_{1}=+\infty$, then $U(1)_{1}$ exists and
$U(1)_{1}=1\,.$
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Using the assumptions and Propositions 2.2, 2.4 and 2.6 we get that $F_{U}(x)^{2}\cdot F_{W}(x)=F_{W}(x)-F_{V}(x)$ for every $x\in[0,\,1)$. Since $F_{W}(0)>0$ and $F_{W}(x)$ is non-decreasing on $[0,1)$, we can divide by $F_{W}(x)$ and get that $F_{U}(x)^{2}=1-\frac{F_{V}(x)}{F_{W}(x)}$ for every $x\in[0,\,1)$. Note that $\lim_{x\to 1}F_{V}(x)=a>0$ because $F_{V}(0)>0$ and $F_{V}(x)$ is non- decreasing and bounded from above. Theorem 2.17 and properties of limits of functions give
$(U(1)_{1})^{2}$ $=$ ${\textstyle\lim_{x\to 1}F_{U}(x)^{2}=\lim_{x\to 1}\big(1-\frac{F_{V}(x)}{F_{W}(x)}\big)=1-\frac{\lim_{x\to 1^{-}}F_{V}(x)}{\lim_{x\to 1^{-}}F_{W}(x)}}$
$=$ ${\textstyle 1-\frac{a}{\lim_{n\to\infty}\sum_{j=0}^{n}w_{j}}=1-\frac{a}{+\infty}=1}\,.$
Hence $U(1)_{1}=1$ because $U(1)_{1}\geq 0$. $\Box$
We proceed to absolute convergence.
Proposition 2.24
Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{2}$ and $W(1)_{2}$ exist, then $U(1)_{2}$ exists and
$U(1)_{2}=\alpha V(1)_{2}+\beta W(1)_{2}\,.$
Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}u_{n}x^{n}$. Since $V(1)_{2}$ and $W(1)_{2}$ are sums of absolutely convergent series $\sum_{n\in\mathbb{N}}v_{n}$ and $\sum_{n\in\mathbb{N}}w_{n}$ and $u_{n}=\alpha v_{n}+\beta w_{n}$, the result follows from Proposition 2.2. $\Box$
Proposition 2.25
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ be in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$ and if $V(1)_{2}$ and $W(1)_{2}$ exist, then the following holds.
  1. 1.
    $U(1)_{2}$ exists and $U(1)_{2}=V(1)_{2}W(1)_{2}$.
  2. 2.
    $\sum_{n\geq 0}|u_{n}|\leq\sum_{n\geq 0}|v_{n}|\cdot\sum_{n\geq 0}|w_{n}|$.
Proof. 1. We consider the series
$S=\sum_{\langle j,\,k\rangle\in\mathbb{N}_{0}^{2}}v_{j}w_{k}=\sum_{j\in\mathbb{N}_{0}}v_{j}\cdot\sum_{k\in\mathbb{N}_{0}}w_{k}\,.$}
It absolutely converges by Proposition 2.5 and has the sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping $\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}$ (the inner $\sum$ means sum and the outer $\sum$ means series) and $\sum_{j=0}^{n}v_{j}w_{n-j}=u_{n}$. By Proposition 2.4, $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}$. Thus $U(1)_{2}=V(1)_{2}W(1)_{2}$.
2. This follows from the inequality ($n\in\mathbb{N}_{0}$)
$\sum_{m=0}^{n}|u_{m}|=\sum_{m=0}^{n}\big|\sum_{j=0}^{m}v_{j}w_{m-j}\big|\leq\sum_{j=0}^{n}|v_{j}|\cdot\sum_{j=0}^{n}|w_{j}|\,.$}
$\Box$
Propositions 2.19 for $n=2$ and item 1 of 2.25 are incomparable. The former assumes only conditional convergence, but for all three series. The latter assumes stronger absolute convergence, but only for two series. Another variant of the proposition is the theorem due to Franz (Franciszek) Mertens (1840–1927).
Theorem 2.26 (Mertens’s)
Let $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$ and if $V(1)_{1}$ and $W(1)_{2}$ exist, then $U(1)_{1}$ exists and
$U(1)_{1}=V(1)_{1}W(1)_{2}\,.$
See [7] for a proof.
Proposition 2.27
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$ and $v_{0}=1$. If $U(x)=\frac{1}{V(x)}$ and $c=|v_{1}|+|v_{2}|+\dots<1$, then the following holds.
  1. 1.
    $V(1)_{2}$ and $U(1)_{2}$ exist, $V(1)_{2}\neq 0$ and $U(1)_{2}=\frac{1}{V(1)_{2}}$.
  2. 2.
    $\sum_{n\geq 0}|u_{n}|\leq\frac{1}{1-c}$.
Proof. 1. It is clear that $V(1)_{2}$ exists. Since $|V(1)_{2}-1|\leq c<1$, we get $V(1)_{2}\neq 0$. We consider the series
$S=1+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}(-1)^{m}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\,.$}
We show that $S$ absolutely converges. It follows from the bound that the sum of absolute values of summands in any finite subseries of $S$ is at most $1+\sum_{m\geq 1}(|v_{1}|+|v_{2}|+\dots)^{m}=\frac{1}{1-c}$.
For $m\in\mathbb{N}$ we consider the series
$S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}\cdot\ldots\cdot v_{j_{m}}$}
By Proposition 2.6 it absolutely converges and has the sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has the sum $1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}$.
On the other hand, in $\mathbb{C}[[x]]$ we have the identity
$U(x)=1+\sum_{m\geq 1}(-1)^{m}(v_{1}x+v_{2}x^{2}+\dots)^{m}\,.$}
Thus $u_{0}=1$ and for every $n\in\mathbb{N}$,
$u_{n}=\sum_{m\geq 1}(-1)^{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=n\end{subarray}}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\,.$}
By Proposition 2.4, $S$ has also the sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$ and the stated formula follows.
2. Indeed,
${\textstyle\sum_{k=0}^{n}|u_{k}|=}$
${\textstyle=1+\sum_{k=1}^{n}\big|\sum_{m\geq 1}(-1)^{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=k\end{subarray}}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\big|}$
${\textstyle\leq\sum_{m\geq 0}(|v_{1}|+|v_{2}|+\dots)^{m}=\frac{1}{1-c}\,.}$
$\Box$
Again, Proposition 2.20 and item 1 of Proposition 2.27 are incomparable.
The analog of Proposition 2.22 for absolute convergence requires two lemmas.
Lemma 2.28
Let $V(x)\in\mathbb{C}[[x]]$ and $V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$ where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$ (every square in $\mathbb{C}[[x]]$ has this form). Then there exists $v\in\mathbb{C}$ such that $v^{2}=t$ and
$V(x)=vx^{l}\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}\,.$}
Proof. Let $n\in\mathbb{N}_{0}$ and a,ba,\,b be formal variables. By the Vandermonde identity,
$\sum_{j=0}^{n}\binom{a}{j}\binom{b}{n-j}=\binom{a+b}{n}\ \ (\in\mathbb{Q}[a,\,b])\,.$}
Let $V_{0}(x)=\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}$. For any $v_{0}\in\mathbb{C}$ such that $(v_{0})^{2}=t$ the instance $a=b=\frac{1}{2}$ of the identity gives
$\big(v_{0}x^{l}V_{0}(x)\big)^{2}$ $=$ ${\textstyle tx^{2l}\sum_{n\geq 0}\big(\sum_{j=0}^{n}\binom{1/2}{j}\binom{1/2}{n-j}\big)(t_{1}x+t_{2}x^{2}+\dots)^{n}}$
$=$ ${\textstyle tx^{2l}\sum_{n\geq 0}\binom{1}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}}$
$=$ $tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)=V(x)^{2}\,.$
Hence $V(x)=\pm v_{0}x^{l}V_{0}(x)$. $\Box$
Lemma 2.29
Let $U(x),V(x)\in\mathbb{C}[[x]]$ and $U(x)=V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$ where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$. If $c=|t_{1}|+|t_{2}|+\dots<1$, then $U(1)_{2}$ and $V(1)_{2}$ exist, and
$U(1)_{2}=\big(V(1)_{2}\big)^{2}\,.$
Proof. Clearly, $U(1)_{2}$ exists. By Lemma 2.28 there is $v\in\mathbb{C}$ such that $v^{2}=t$ and
$V(x)=\sum_{n\geq 0}v_{n}x^{n}=vx^{l}\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}\,.$}
Thus $v_{n}=0$ for $n<l$, $v_{l}=v$ and for $n>l$,
$v_{n}=v\sum_{m\geq 1}\binom{1/2}{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=n-l\end{subarray}}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}\,.$}
We consider the series
$S=0+0+\dots+0+v+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}v\binom{1/2}{m}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}$}
with $l$ zero summands. We show that it absolutely converges. Let $T$ be any finite subseries of $S$. Indeed, since $\big|\binom{1/2}{m}\big|\leq 1$ for every $m\in\mathbb{N}$, the sum of absolute values of summands in $T$ is at most
$|v|+|v|\sum_{m\geq 1}(|t_{1}|+|t_{2}|+\dots)^{m}=\frac{|v|}{1-c}\,.$}
By Proposition 2.4, $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.25, $U(1)_{2}$ exists and $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
Proposition 2.30
Let $U(x),V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ be in $\mathbb{C}[[x]]$. Suppose that $V(1)_{2}$ exists, $w_{0}=1$, $U(x)^{2}=1-\frac{V(x)}{W(x)}$ and $c=|w_{1}|+|w_{2}|+\dots<1$. If $V(x)=W(x)$, then $U(x)=0$. Else, if $k\in\mathbb{N}_{0}$ is minimum such that $v_{k}\neq w_{k}$ and
$\big(|w_{k}-v_{k}|\cdot(1-c)\big)^{-1}\sum_{n>k}|w_{n}-v_{n}|<1\,$,}
then $U(1)_{2}$ and $W(1)_{2}$ exist, $W(1)_{2}\neq 0$ and
$\big(U(1)_{2}\big)^{2}=1-\frac{V(1)_{2}}{W(1)_{2}}\,.$}
Proof. If $V(x)=W(x)$ then $U(x)^{2}=0$ and $U(x)=0$. Let $V(x)\neq W(x)$ and $k\in\mathbb{N}_{0}$ be as stated. We write
$U(x)^{2}=\frac{W(x)-V(x)}{W(x)}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$}
where $l\in\mathbb{N}_{0}$, $t=w_{k}-v_{k}\in\mathbb{C}^{*}$ and $t_{j}\in\mathbb{C}$. Let
$U_{0}(x)=\frac{1}{W(x)}=\sum_{n\geq 0}u_{0,n}x^{n},\ U_{1}(x)=W(x)-V(x)$}
and $U_{2}(x)=U_{0}(x)U_{1}(x)=\sum_{n\geq 0}u_{2,n}x^{n}$. By item 1 of Proposition 2.27, $U_{0}(1)_{2}$ and $W(1)_{2}$ exist, $W(1)_{2}\neq 0$ and $U_{0}(1)_{2}=\frac{1}{W(1)_{2}}$. By Proposition 2.24, $U_{1}(1)_{2}$ exists and $U_{1}(1)_{2}=W(1)_{2}-V(1)_{2}$. By item 1 of Proposition 2.25, $U_{2}(1)_{2}$ exists and $U_{2}(1)_{2}=U_{0}(1)_{2}\cdot U_{1}(1)_{2}$. By item 2 of Proposition 2.27, $\sum_{n\geq 0}|u_{0,n}|\leq\frac{1}{1-c}$. By item 2 of Proposition 2.25,
$\sum_{n\geq 0}|u_{2,n}|\leq\sum_{n\geq 0}|u_{0,n}|\cdot\sum_{n>k}|w_{n}-v_{n}|\leq\frac{1}{1-c}\sum_{n>k}|w_{n}-v_{n}|\,.$}
Since $U(x)^{2}=U_{2}(x)$, we have bound
$|t_{1}|+|t_{2}|+\dots\leq\frac{1}{|t|(1-c)}\sum_{n>k}|w_{n}-v_{n}|<1\,.$}
By Lemma 2.29, $U(1)_{2}$ exists and
$\big(U(1)_{2}\big)^{2}=U_{2}(1)_{2}=U_{0}(1)_{2}\cdot U_{1}(1)_{2}=1-\frac{V(1)_{2}}{W(1)_{2}}\,.$}
$\Box$
Finally, we consider light convex weights.
Proposition 2.31
For every convex light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ we have
$\sum_{n=0}^{\infty}h(W(n))\cdot x^{n}=\frac{1}{1-x}$}\,.
Proof. Let $h$ be a convex light weight. We prove by induction on $n\in\mathbb{N}_{0}$ that $h(W(n))=1$. For $n=0$ this holds as $W(0)=\{\langle 1\rangle\}$. Let $n>0$. Recall that $\ell(w)$ denotes the last vertex of a walk $w$. Using Propositions 2.2 and 2.4, the convexity assumption and induction, we get
$h(W(n))$ $=$ ${\textstyle\sum_{w\in W(n)}h(w)}$
$=$ ${\textstyle\sum_{\begin{subarray}{c}w^{\prime}\in W(n-1)\\ \ell(w^{\prime})\in V(h)\end{subarray}}h(w^{\prime})\sum_{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}}h(\{\ell(w^{\prime}),\,v\})}$
$=$ ${\textstyle\sum_{w^{\prime}\in W(n-1)}h(w^{\prime})\cdot 1=1\,.}$
We can add and remove the condition $\ell(w^{\prime})\in V(h)$ without affecting the sum because walks $w^{\prime}$ not satisfying it have zero weights. $\Box$
3 Extensions of Pólya’s theorem with $v=1$
In this section we establish the first group of generalizations of Pólya’s theorem, namely Theorems 3.23.7. We treat the case $v=1$ when the vertex to be visited by walks coincides with their starting vertex.
We introduce some generating functions in $\mathbb{C}[[x]]$. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the set of walks that visit $v$ at a step $i>0$. In this section, $v=1$. We are interested in the generating function
$A_{h}(x)=\sum_{n\geq 0}a_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(1,\,n))x^{n}\,.$}
We also define
$B_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h}x^{n}=\sum_{n\geq 0}h(W_{b}(n))x^{n}\,,}$
$C_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h}x^{n}=\sum_{n\geq 0}h(W_{c}(n))x^{n}\,\text{ and}}$
$D_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}d_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(n))x^{n}\,,}$
where $W_{b}(n)$ ($\subset W(n)$) are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and $W_{c}(n)$ ($\subset W_{b}(n)$) are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$ but avoid $1$ between. The coefficients of $A_{h}(x)$, $\dots$, $D_{h}(x)$ are correctly defined because they are sums of subseries of absolutely convergents series. We have $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$.
Proposition 3.1
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be as above. The following relations hold between them.
  1. 1.
    $A_{h}(x)=C_{h}(x)D_{h}(x)$.
  2. 2.
    $B_{h}(x)=\frac{1}{1-C_{h}(x)}$, equivalently, $C_{h}(x)=1-\frac{1}{B_{h}(x)}$.
Proof. 1. We show that $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}\cdot d_{n-j}^{h}$ for every $n\in\mathbb{N}_{0}$. For $n\leq 1$ it holds, $a_{n}^{h}=c_{n}^{h}=0$. For $n\geq 2$ the set $W(1,n)$ is countable and we split every walk in it at the first revisit of $1$. We get a map
$F\colon W(1,\,n)\to\bigcup_{j=0}^{n}W_{c}(j)\times W(n-j)$}
that is defined as follows. For any walk
$w=\langle u_{0},\,u_{1},\,\dots,\,u_{j},\,\dots,\,u_{n}\rangle\in W(1,\,n)\,$,
where $u_{0}=u_{j}=1$, $j>0$ and $u_{i}\neq 1$ for $i\in[j-1]$, we set
$F(w)=\langle w_{1},\,w_{2}\rangle\ \ (\in W_{c}(j)\times W(n-j))\,$,
where $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$. It is easy to see that the map $F$ is a bijection. $F$ is weight-preserving (WP): for every weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and value $F(w)=\langle w_{1},w_{2}\rangle$ we have $h(w)=h(w_{1})h(w_{2})$. If $X$ is a set of pairs of walks in $K_{\mathbb{N}}$, we write $h(X)$ for the sum of the series
$\sum_{\langle w,\,w^{\prime}\rangle\in X}h(w)h(w^{\prime})\,$,}
provided that it absolutely converges.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and $n\geq 2$. Using $F$ and above propositions, we get
$a_{n}^{h}$ $=$ $h(W(1,\,n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W(1,\,n)])$
[…]
2. We show that
$B_{h}(x)=\frac{1}{1-C_{h}(x)}=1+\sum_{j\geq 1}(C_{h}(x))^{j}\,.$}
We recall that $b_{0}^{h}=1$ and $c_{0}^{h}=0$ and show that for every $n\in\mathbb{N}$,
$b_{n}^{h}=\sum_{j=1}^{\infty}\sum_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}c_{n_{1}}^{h}c_{n_{2}}^{h}\dots c_{n_{j}}^{h}\,.$}
For odd $n$ it is true because then $b^{h}_{n}=c^{h}_{n}=0$ as $W_{b}(n)=W_{c}(n)=\emptyset$, and in every inner summand at least one index $n_{i}$ is odd. For even $n\geq 2$ the set $W_{b}(n)$ is countable. We split every walk in it at the visits of $1$ and get a map
$F\colon W_{b}(n)\to\bigcup_{j=1}^{\infty}\bigcup_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}W_{c}(n_{1})\times\cdots\times W_{c}(n_{j})$}
defined as follows. For any walk
$w=\langle u_{0},\,\dots,\,u_{m_{1}},\,\dots,\,\dots,\,u_{m_{j-1}},\,\dots,\,u_{n}\rangle\in W_{b}(n)\,$,
where $0=m_{0}<m_{1}<\dots<m_{j}=n$ and $u_{k}=1$ iff $k=0$ or $k=m_{i}$ for some $i\in[j]$, we set
$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{i}),\,n_{i}=m_{i}-m_{i-1})$
where $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i\in[j]$. It is easy to see that $F$ is a bijection and that it is weight-preserving similarly as above. If $X$ is a set of tuples of walks, we write $h(X)$ for the sum of the series
$\sum_{\langle w_{1},\,\dots,\,w_{k})\in X}h(w_{1})\dots h(w_{k})$}
if it absolutely converges.
Let $h$ be a light weight and $n\geq 2$ be even. Using $F$ and above propositions, we get
$b_{n}^{h}$ $=$ $h(W_{b}(n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W_{b}(n)])$
[…]
$\Box$
For better orientation we label our theorems by the triples
$\langle x,\,y,\,z\rangle\in\{v=1,\,v\neq 1\}\times\{\mathrm{gen},\,\mathrm{con}\}\times\{U(1)_{1},\,U(1)_{2},\,U(1)_{1}=+\infty\}\,$,
where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to convex ones. The generating functions $A_{h}(x)$, $\dots$, $D_{h}(x)$ are as above
Theorem 3.2 (1: $v=1$, gen, $U(1)_{1}$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Suppose that $D_{h}(1)_{1}$, $C_{h}(1)_{1}$, $B_{h}(1)_{1}$ and $A_{h}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
$A_{h}(1)_{1}=\big(1-\frac{1}{B_{h}(1)_{1}}\big)D_{h}(1)_{1}\,.$}
Proof. By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By Proposition 2.19, $A_{h}(1)_{1}=C_{h}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By Proposition 2.20 we have $B_{h}(1)_{1}\neq 0$ and $C_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
Theorem 3.3 (2: $v=1$, gen, $U(1)_{2}$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Suppose that $D_{h}(1)_{2}$ and $C_{h}(1)_{2}$ exist, and that $\sum_{n\geq 1}|b_{n}^{h}|<1$. Then $B_{h}(1)_{2}$ exists and is nonzero, $A_{h}(1)_{2}$ exists and
$A_{h}(1)_{2}=\big(1-\frac{1}{B_{h}(1)_{2}}\big)D_{h}(1)_{2}\,.$}
Proof. By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By item 1 of Proposition 2.25, $A_{h}(1)_{2}$ exists and $A_{h}(1)_{2}=C_{h}(1)_{2}D_{h}(1)_{2}$ . By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h}(1)_{2}$ exists and $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. The stated formula follows. $\Box$
Theorem 3.4 (3: $v=1$, gen, $U(1)_{1}=+\infty$)
Suppose that $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ is a light weight and $D_{h}(1)_{1}=+\infty$. Then
$A_{h}(1)_{1}=+\infty\,.$
Proof. Thus $d_{n}^{h}>0$ for some $n>0$ and there is a vertex $u\in\mathbb{N}\setminus\{1\}$ with $h(\{1,u\})>0$. Then the walk $\langle 1,u,1\rangle$ shows that $c_{2}^{h}>0$. Since by item 1 of Proposition 3.1 we have $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}d_{n-j}^{h}$, we get $a_{n+2}^{h}\geq c_{2}^{h}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}a_{j+2}^{h}\geq c_{2}^{h}\cdot\sum_{j=0}^{n}d_{j}^{h}\to+\infty,\ n\to\infty\,$,}
and $A_{h}(1)_{1}=+\infty$. $\Box$
We proceed to convex weights.
Theorem 3.5 (4: $v=1$, con, $U(1)_{1}$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a convex light weight. Suppose that $B_{h}(1)_{1}$ and $C_{h}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
$\lim_{n\to\infty}{\textstyle a_{n}^{h}=1-\frac{1}{B_{h}(1)_{1}}\,.}$
Proof. By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}$. By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By Proposition 2.20 we have $B_{h}(1)_{1}\neq 0$ and $C_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
This theorem generalizes the case $d\geq 3$ of Pólya’s theorem with $\overline{v}=\overline{0}$. The case $d\leq 2$ is generalized in Theorem 3.7.
Theorem 3.6 (5: $v=1$, con, $U(1)_{2}$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a convex light weight. Suppose that $\sum_{n\geq 1}|b_{n}^{h}|<1$. Then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}{\textstyle a_{n}^{h}=1-\frac{1}{B_{h}(1)_{2}}\,.}$
Proof. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h}(1)_{2}$ exists and $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}$. By Proposition 2.15,
$\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)_{1}=C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$}
$\Box$
Theorem 3.7 (6: $v=1$, con, $U(1)_{1}=+\infty$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a convex light weight and $B_{h}(1)=+\infty$. Then
$\lim_{n\to\infty}a_{n}^{h}=1\,.$
Proof. By item 1 of Proposition 3.1 and Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}$. By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. But now by Proposition 2.31 we have $0\leq b_{n}^{h},c_{n}^{h}\leq d_{n}^{h}=1$ for every $n\in\mathbb{N}_{0}$. Hence $b_{n}^{h},c_{n}^{h}\in[0,1]$ and the generating functions $B_{h}(x)$ and $C_{h}(x)$ absolutely converge for $x\in[0,1)$. By Proposition 2.21, $C_{h}(1)_{1}=1$. The stated formula follows. $\Box$
4 Extensions of Pólya’s theorem with $v\neq 1$
The case $v\neq 1$ when the vertex to be visited by walks differs from their starting vertex is more complicated than the previous case. The grid graph $\mathbb{Z}^{d}$ is vertex-transitive but so far this was irrelevant. For $\overline{v}\neq\overline{0}$ vertex-transitivity becomes relevant and we generalize it to $K_{\mathbb{N}}$ as follows. Let $v\in\mathbb{N}\setminus\{1\}$ with $v\neq 1$. We say that a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $v$-transitive if there is a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that $f(1)=v$ and $h(f[e])=h(e)$ for every $e\in\mathbb{N}_{2}$.
We again introduce some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$ with $v\neq 1$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ are the walks visiting $v$ ($\neq 1$) at a step $i>0$. We are interested in the generating function
$A_{h,v}(x)=\sum_{n\geq 0}a_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W(v,\,n))x^{n}\,.$}
The generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ are as before. We additionally introduce
$B_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{b,v}(n))x^{n}\,,}$
$C_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{c,v}(n))x^{n}\,\text{ and}}$
$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e,v}(n))x^{n}\,.}$
Here $W_{b,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$. $W_{c,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding $v$. Finally, $W_{e,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding both $1$ and $v$. We have $a_{0}^{h,v}=c_{0}^{h,v}=c_{0}^{h}=e_{0}^{h,v}=0$ and $b_{0}^{h}=b_{0}^{h,v}=d_{0}^{h}=1$.
We obtain an analog of Proposition 3.1. The proof is similar. We omit details and argue only on the high level of semiformal method.
Proposition 4.1
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and the generating functions $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ in $\mathbb{C}[[x]]$ be as above. The following relations hold between them.
  1. 1.
    $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$.
  2. 2.
    $B_{h}(x)=B_{h,v}(x)+C_{h,v}(x)^{2}B_{h}(x)$, equivalently, $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
  3. 3.
    $C_{h,v}(x)=B_{h,v}(x)E_{h,v}(x)$.
Proof. 1. Every walk $w$ in $\bigcup_{n\geq 0}W(v,n)$ splits at the first visit of $v$ in two walks as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $C_{h,v}(x)$ and $w_{2}$ are arbitrary walks starting at $v$. By the $v$-transitivity of $h$, walks $w_{2}$ are weight-counted by $D_{h}(x)$. The first relation follows.
2. It suffices to prove the first equality. $B_{h}(x)$ weight-counts walks $w$ starting and ending at $1$. Those avoiding $v$ are weight-counted by $B_{h,v}(x)$. If $w$ visits $v$, then it uniquely splits at the first and last visits of $v$ in three walks as $w=w_{1}w_{2}w_{3}$. Walks $w_{1}$ start at $1$, end at $v$ and between avoid $v$. Walks $w_{2}$ start and end at $v$. Walks $w_{3}$ start at $v$, end at $1$ and between avoid $v$. We reverse walks $w_{3}$ and see that both $w_{1}$ and $w_{3}$ are weight-counted by $C_{h,v}(x)$. By the $v$-transitivity of $h$, the middle walks $w_{2}$ are weight-counted by $B_{h}(x)$. The second relation follows.
3. We consider walks $w$ weight-counted by $C_{h,v}(x)$. They start at $1$, end at $v$, and between avoid $v$. These walks uniquely split at the last visit of $1$ in two walks as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $B_{h,v}(x)$, and $w_{2}$ by $E_{h,v}(x)$. The third relation follows. $\Box$
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $v\in\mathbb{N}\setminus\{1\}$. If $v\not\in V(h)$ then
$a_{n}^{h,\,v}=h(W(v,\,n))=\sum_{w\in W(v,\,n)}h(w)=\sum_{w\in W(v,\,n)}0=0$}
for every $n\in\mathbb{N}_{0}$ because every walk starting at $1$ and ending at $v$ contains an edge with zero weight. Henceforth we therefore assume that $v\in V(h)$. For $w,z\in\mathbb{C}$ we define the set
$w\cdot\mathrm{sqrt}(z)=\{w\alpha\colon\;\alpha\in\mathbb{C},\,\alpha^{2}=z\}\ \ (\subset\mathbb{C})\,.$
For $w=0$ or $z=0$ it equals $\{0\}$. Else it has two non-zero elements differing by sign. If $w=1$, we write just $\mathrm{sqrt}(z)$.
Theorem 4.2 (7: $v\neq 1$, gen, $U(1)_{1}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. Suppose that $D_{h}(1)_{1}$, $C_{h,v}(1)_{1}$, $C_{h}(1)_{1}$, $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$ and $A_{h,v}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
$A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,.$
Proof. By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By Proposition 2.19, $A_{h,v}(1)_{1}=C_{h,v}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By item 2 of Proposition 3.1 and by Proposition 2.20, $B_{h}(1)_{1}\neq 0$. Thus by Proposition 2.22, $C_{h,v}(1)_{1}^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
We assume the existence of $E_{h,v}(1)_{1}$ instead of $C_{h}(1)_{1}$ and get the following variant of Theorem 4.2.
Theorem 4.3 (7: $v\neq 1$, gen, $U(1)_{1}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. Suppose that $E_{h,v}(1)_{1}$, $D_{h}(1)_{1}$, $C_{h,v}(1)_{1}$, $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$ and $A_{h,v}(1)_{1}$ exist. Then
$A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,$,
where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret the fraction $\frac{0}{0}$ as $1$, so that $A_{h,v}(1)_{1}=0$.
Proof. For $B_{h}(1)_{1}\neq 0$ we use the previous proof. If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.19 and 2.18 that also $B_{h,v}(1)_{1}=0$. Item 3 of Proposition 4.1 and Proposition 2.19 give $C_{h,v}(1)_{1}=0$. Finally, item 1 of Proposition 4.1 and Proposition 2.19 give $A_{h,v}(1)_{1}=0$. $\Box$
Theorem 4.4 (8: $v\neq 1$, gen, $U(1)_{2}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. Suppose that $D_{h}(1)_{2}$ and $B_{h,v}(1)_{2}$ exist, and that $c=\sum_{n\geq 1}|b_{n}^{h}|<1$. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. Let $B_{h,v}(x)\neq B_{h}(x)$ and $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1$}
then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ and $A_{h,v}(1)_{2}$ exist, and
$A_{h,\,v}(1)_{2}\in{\textstyle D_{h}(1)_{2}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
Proof. If $B_{h,v}(x)=B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=1-1=0$ and $C_{h,v}(x)=0$. Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed inequality hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $C_{h,v}(1)_{2}^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By item 1 of Proposition 2.25, $A_{h,v}(1)_{2}$ exists and $A_{h,v}(1)_{2}=C_{h,v}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
Theorem 4.5 (9: $v\neq 1$, gen, $U(1)_{1}=+\infty$)
Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive light weight such that $v\in V(h)$. If $D_{h}(1)_{1}=+\infty$ then
$A_{h,v}(1)_{1}=+\infty\,.$
Proof. Let $m\in\mathbb{N}$ be the minimum length of a walk in $G(h)$ joining $1$ and $v$. The inner vertices of this shortest walk differ from $v$ and we get that $c^{h,v}_{m}>0$. Since by item 1 of Proposition 4.1 we have $a_{n}^{h,v}=\sum_{j=0}^{n}c_{j}^{h,v}d_{n-j}^{h}$, we conclude that $a_{n+m}^{h,v}\geq c_{m}^{h,v}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}a_{j+m}^{h,v}\geq c_{m}^{h,v}\cdot\sum_{j=0}^{n}d_{j}^{h}$}
for every $n\in\mathbb{N}_{0}$ and $A_{h,v}(1)_{1}=+\infty$. $\Box$
We proceed to convex weights.
Theorem 4.6 (10: $v\neq 1$, con, $U(1)_{1}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive convex light weight such that $v\in V(h)$. Suppose that $C_{h,v}(1)_{1}$, $C_{h}(1)_{1}$, $B_{h,v}(1)_{1}$ and $B_{h}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,.$
Proof. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$. By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h,v}=C_{h,v}(1)_{1}$. By item 2 of Proposition 3.1 and by Proposition 2.20, $B_{h}(1)_{1}\neq 0$. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.22, $C_{h,v}(1)_{1}^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
We again assume the existence of $E_{h,v}(1)_{1}$ instead of $C_{h}(1)_{1}$ and get the following variant of the theorem.
Theorem 4.7 (10: $v\neq 1$, con, $U(1)_{1}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive convex light weight such that $v\in V(h)$. Suppose that $E_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$, $B_{h,v}(1)_{1}$ and $B_{h}(1)_{1}$ exist. Then
$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\ \ (\subset\mathbb{C})\,$,
where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret the fraction $\frac{0}{0}$ as $1$, so that the limit is zero.
Proof. If $B_{h}(1)_{1}\neq 0$, the previous proof works. If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.19 and 2.18 that also $B_{h,v}(1)_{1}=0$. Item 3 of Proposition 4.1 and Proposition 2.19 give $C_{h,v}(1)_{1}=0$. Thus, as in the previous proof, $\lim_{n\to\infty}a_{n}^{h,v}=C_{h,v}(1)_{1}=0$. $\Box$
These two theorems generalize the case $d\geq 3$ of Pólya’s theorem with $\overline{v}\neq\overline{0}$. The case $d\leq 2$ is generalized in Theorem 4.9.
Theorem 4.8 (11: $v\neq 1$, con, $U(1)_{2}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive convex light weight such that $v\in V(h)$. Suppose that $B_{h,v}(1)_{2}$ exists and $c=\sum_{n\geq 1}|b_{n}^{h}|<1$. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. Let $B_{h,v}(x)\neq B_{h}(x)$ and $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1$}
then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ and $A_{h,v}(1)_{2}$ exist, and
$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle D_{h}(1)_{2}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
Proof. If $B_{h,v}(x)\neq B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=1-1=0$ and $C_{h,v}(x)=0$. Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed inequality hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $C_{h,v}(1)_{2}^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$. Proposition 2.15 gives the stated formula:
$\lim_{n\to\infty}a_{n}^{h,v}={\textstyle C_{h,\,v}(1)_{1}=C_{h,\,v}(1)_{2}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}\big)\,.}$
$\Box$
Theorem 4.9 (12: $v\neq 1$, con, $U(1)_{1}=+\infty$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive convex light weight such that $v\in V(h)$. Suppose that $B_{h}(1)_{1}=+\infty$. Then
$\lim_{n\to\infty}a_{n}^{h,v}=1\,.$
Proof. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-x}$. Hence $0\leq b_{n}^{h},b_{n}^{h,v},c_{n}^{h,v},e_{n}^{h,v}\leq d_{n}^{h}=1$ for every $n\in\mathbb{N}_{0}$ and we can use continuous and non-decreasing functions
$F_{B_{h}},\,F_{B_{h,v}},\,F_{C_{h,v}},\,F_{E_{h,v}}\colon[0,\,1)\to[0,\,+\infty)\,.$
Any shortest walk in $G(h)$ joining $1$ and $v$ shows that for some $m\in\mathbb{N}$ we have $e_{m}^{h,v}>0$. Thus the function $F_{E_{h,v}}(x)$ is on the interval $(0,1)$ positive and increasing. We deduce from it that $F_{B_{h,v}}(x)$ is bounded on $[0,1)$.
Suppose for the contrary that $\lim_{x\to 1}F_{B_{h,v}}(x)=+\infty$. By item 3 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality $F_{C_{h,v}}(x)=F_{B_{h,v}}(x)F_{E_{h,v}}(x)$. Thus $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$. From the facts that $F_{C_{h,v}}(0)=0$, $F_{C_{h,v}}(x)$ increases and is continuous and $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$, we obtain a (unique) number $x_{0}\in(0,1)$ such that $F_{C_{h,v}}(x)<1$ for $x\in[0,x_{0})$ and $\lim_{x\to x_{0}^{-}}F_{C_{h,v}}(x)=1$. By item 2 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality
$F_{B_{h}}(x)=F_{B_{h,v}}(x)+F_{C_{h,v}}(x)^{2}F_{B_{h}}(x)\,\,.$
Hence for every $x\in[0,x_{0})$ we have $F_{B_{h}}(x)=\frac{F_{B_{h,v}}(x)}{1-F_{C_{h,v}}(x)^{2}}$. Since $F_{B_{h,v}}(0)=1$, it follows that $\lim_{x\to x_{0}^{-}}F_{B_{h}}(x)=+\infty$. But this contradicts the continuity of $F_{B_{h}}(x)$ at $x_{0}$.
Using item 2 of Proposition 4.1, the boundedness of $F_{B_{h,v}}(x)$ and Proposition 2.23, we get $C_{h,v}(1)_{1}=1$. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$. By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h,v}=C_{h,v}(1)_{1}=1$. $\Box$
5 Concluding remarks
In [16] we consider the same problem of determination of formulas for the quantities
$A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))$}
for formal weights $h$. These are maps
$h\colon\mathbb{N}_{2}\to\mathbb{C}[[x_{1},\,x_{2},\,\dots,\,x_{k}]]\,$,
where the last integral domain of formal power series in $k$ variables is endowed with the usual non-Archimedean norm. In [16] we will work out some concrete examples, both for complex and formal weights.
The main novelty of our generalization is this. If we return to the equality
$P_{d}(\overline{v},\,n)=h(W(v,\,n))$
in the Introduction, the left-hand side involves, in the standard rigorous probabilistic treatment, an implicit probability space that supports the corresponding Markov chain. Construction of this probability space is far from trivial; see, for example, [6]. In our approach, the right-hand side is completely self-contained.

### v3

@@
 $\lim_{n\to\infty}P_{d}(\overline{v},\,n)\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ <1&\dots&\text{if $d\geq 3$}\,.\end{array}\right$.
-Pólya’s theorem is almost always cited in the special case with $\overline{v}=\overline{0}$, although G.
-Pólya proved it for any vertex $\overline{v}$ (see [15, Section 1] for a translated quote from [23]).
+Pólya’s theorem is almost always cited in the special case with $\overline{v}=\overline{0}$, although he proved it for any vertex $\overline{v}$ (see [16, Section 1] for a translated quote from [24]).
 A selection of articles and books discussing Pólya’s theorem is [3, 4, 6, 9, 11, 13, 18, 19, 20, 21, 25, 26, 29, 30].
@@
 $\mathbb{N}_{2}=\binom{\mathbb{N}}{2}=\{e\colon\;e\subset\mathbb{N}\wedge|e|=2\}$}
-be the edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$ on the vertex set $\mathbb{N}$.
+be edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$ on the vertex set $\mathbb{N}$.
 For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1.\dots\}$) its cardinality.
@@
 $A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))\,.$}
-We treat this series both in conditional and absolute convergence.
-Our formulas appear in fourteen theorems: Theorems 3.23.7 in Section 3 and Theorems 4.24.9 in Section 4.
-Pólya’s theorem is subsumed in five of them: Theorems 3.5 and 3.7 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}=\overline{0}$, and Theorems 4.6 (4.7) and 4.9 cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
-Why fourteen generalizations of one theorem?
-For three reasons.
+The formulas appear in fourteen theorems: Theorems 3.23.7 in Section 3 and Theorems 4.24.9 in Section 4.
+Pólya’s theorem is subsumed in five of them.
+Theorems 3.5 and 3.7 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}=\overline{0}$.
+Theorems 4.6 (4.7) and 4.9 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
+We have fourteen generalizations of one theorem for three reasons.
 First, in Pólya’s theorem weights are nonnegative real numbers and convergence of infinite series is automatically absolute.
-For complex weights we consider both absolute, and conditional convergence.
+For complex weights we consider both kinds of convergence, absolute and conditional one.
 We also consider, for nonnegative real weights, series with sum $\infty$.
 Second, Pólya’s theorem lives in the probabilistic setup where weights of edges incident to a fixed vertex sum up to $1$.
-We consider two cases—when this condition is kept (convex weights) and when it is dropped.
-Third, already Pólya’s theorem is in reality two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in the complexity of proofs.
-In our generalization we correspondingly distinguish cases $v=1$ and $v\neq 1$.
-Now $3\cdot 2\cdot 2=12$ and two more generalizations arise by variations of assumptions.
+We consider the case when this condition is kept (convex weights) and the case when it is dropped.
+Third, already Pólya’s theorem is in reality two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in complexity of proofs.
+In our generalization we distinguish cases $v=1$ and $v\neq 1$ as well.
+Now $3\cdot 2\cdot 2=12$; two more generalizations arise by variations of assumptions.
 We explain how our approach includes Pólya’s original problem. If $f\colon A\to B$ is a map and $C$ is any set (not necessarily a subset of $A$), we call the set
 $f[C]=\{f(x)\colon\;x\in A\cap C\}\ \ (\subset B)$
-the image of $C$ by $f$. Let $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$. We take any bijection
-$f\colon\mathbb{Z}^{d}\to\mathbb{N}$
-such that $f(\overline{0})=1$, and set $v=f(\overline{v})$. The grid graph on $\mathbb{Z}^{d}$, i.e. the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
+the image of $C$ by $f$.
+Let $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$.
+We take any bijection $f\colon\mathbb{Z}^{d}\to\mathbb{N}$ such that $f(\overline{0})=1$, and set $v=f(\overline{v})$.
+The grid graph on $\mathbb{Z}^{d}$, which is the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
 $\{\overline{a},\,\overline{b}\}\in E_{d}\iff\sum_{i=1}^{d}|a_{i}-b_{i}|=1\,$,}
@@
 Section 5 contains concluding remarks.
-We continue our investigation in [16] where we consider edge weights
-$h\colon\mathbb{N}_{2}\to\mathbb{C}[[x_{1},\,\dots,\,x_{k}]]\,.$
-The domain of formal power series $\mathbb{C}[[x_{1},\dots,x_{k}]]$ is endowed with the standard non-Archimedean norm.
+We continue our investigation in [17] (see Section 5 for more details).
 In Sections 3 and 4 we use what we term the semiformal method in (enumerative) combinatorics.
@@
 Generating “functions” in the symbolic method are formal power series with coefficients determined by finite operations.
-Our extension allows operations with coefficients based on limit transitions.
-2 Series and generating functions
-We review series and generating functions. A series is a map
-$h\colon X\to\mathbb{C}$
-defined on an at most countable set $X$. We write it as $\sum_{x\in X}h(x)$. We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
+Our extension allows operations based on limit transitions.
+See [15] more applications of the semiformal method.
+2 Series and generating functions in $\mathbb{C}$
+We review series and generating functions in the domain $\mathbb{C}$.
+A series (in $\mathbb{C}$) is a map $h\colon X\to\mathbb{C}$, defined on an at most countable set $X$.
+We write it as $\sum_{x\in X}h(x)$.
+We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
   1. 1.
@@
 such that $n\in\mathbb{N}_{0}$ and $v_{i-1}\neq v_{i}$ for every $i\in[n]$ ($\{1,2,\dots,n\}$, $0]=\emptyset$).
-The length $n$ of $w$ is denoted by $|w|$ ($\in\mathbb{N}_{0}$).
+The length $n$ of $w$ is is $n$ ($\in\mathbb{N}_{0}$).
 We extend $h$ to walks.
@@
 $h(w)=\prod_{i=1}^{n}h(\{v_{i-1},\,v_{i}\})\,$,}
-and for $n=0$ we define $h(w)=1$. Recall that $W(n)$ is the set of walks $w$ in $K_{\mathbb{N}}$ such that $v_{0}=1$ and $|w|=n$.
+and for $n=0$ we define $h(w)=1$. Recall that $W(n)$ is the set of walks $w$ in $K_{\mathbb{N}}$ with length $n$ such that $v_{0}=1$.
 Definition 2.1 (lightness)
@@
 If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$.
-Any subseries of an absolutely convergent series absolutely converges, but we cannot compare sizes of sums
+Any subseries of an absolutely convergent series absolutely converges, but we cannot compare sizes of their sums
 $\big|\sum_{x\in Y}h(x)\big|\,\text{ and }\,\big|\sum_{x\in X}h(x)\big|\,.$}
-Except when $h\colon X\to\mathbb{R}_{\geq 0}$. If the series absolutely converges and $Y\subset X$, then the sums satisfy
+Except when $h\colon X\to\mathbb{R}_{\geq 0}$: if $\sum_{x\in X}h(x)$ absolutely converges and $Y\subset X$, then the sums satisfy
 $0\leq\sum_{x\in Y}h(x)\leq\sum_{x\in X}h(x)\,.$}
@@
 Proposition 2.4
-If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, and $P$ is a partition of $X$, then $S_{P}$ is correctly defined, absolutely converges and has the same sum $s$.
-Proof.
-For every $Z\in P$ the series $S_{Z}=\sum_{x\in Z}h(x)$ is a subseries of $S$ and absolutely converges.
-Thus the series $S_{P}$ is correctly defined and we show that it absolutely converges.
-Let
+If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, and if $P$ is a partition of $X$, then $S_{P}$ is correctly defined, absolutely converges and has the same sum $s$.
+Proof. For every $Z\in P$ the series $S_{Z}=\sum_{x\in Z}h(x)$ is a subseries of $S$ and absolutely converges. Thus the series $S_{P}$ is correctly defined. We show that it absolutely converges. Let
 $c=\sup(\{\sum_{x\in Z}|h(x)|\colon\;\text{$Z\subset X$ and $Z$ is finite}\})\ \ (<+\infty)\,.$}
-Let $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. We use Lemma 2.3 and take for each $i\in[n]$ the finite set
+Let $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. We use Lemma 2.3 and take for $i\in[n]$ finite sets
 $Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i})\ \ (\subset Z_{i})\,.$
@@
 If $S_{1}$ and $S_{2}$ absolutely converge and have sums $s_{1}$ and $s_{2}$, respectively, then $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s_{1}s_{2}$.
-Proof. We show that $S_{1}\cdot S_{2}$ absolutely converges. Let $c\geq 0$ be such that for any finite sets $Y_{i}$ with $Y_{i}\subset X_{i}$ and $i=1,2$,
-$\sum_{x\in Y_{i}}|h_{i}(x)|\leq c\,.$}
-Let $A\subset X_{1}\times X_{2}$ be a finite set. We take any finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
+Proof.
+We show that $S_{1}\cdot S_{2}$ absolutely converges.
+Let $c\geq 0$ be such that for any finite sets $Y_{i}\subset X_{i}$, $i=1,2$, we have $\sum_{x\in Y_{i}}|h_{i}(x)|\leq c$.
+Let $A\subset X_{1}\times X_{2}$ be a finite set.
+We take any finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
 $\sum_{\langle x,\,y\rangle\in A}|h_{1}(x)h_{2}(y)|\leq\sum_{x\in Y_{1}}|h_{1}(x)|\cdot\sum_{y\in Y_{2}}|h_{2}(y)|\leq c\cdot c=c^{2}\,.$}
@@
 ${\textstyle\leq\frac{\varepsilon}{4}+|(s_{1}+t_{1})(s_{2}+t_{2})-s_{1}s_{2}|\,,}$
-where
-[…]
-Hence $|s-s_{1}s_{2}|\leq\frac{\varepsilon}{4}+|s_{1}t_{2}|+|s_{2}t_{1}|+|t_{1}t_{2}|\leq\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}=\varepsilon$. $\Box$
-Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2,\dots,k$, be series. We define more generally their product as the series
+where $|t_{1}|\leq\frac{\varepsilon}{4(1+|s_{2}|)}$ and $|t_{2}|\leq\frac{\varepsilon}{4(1+|s_{1}|)}$.
+Hence $|s-s_{1}s_{2}|\leq\frac{\varepsilon}{4}+|s_{1}t_{2}|+|s_{2}t_{1}|+|t_{1}t_{2}|\leq\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}=\varepsilon$.
+$\Box$
+Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i\in[k]$, be $k\in\mathbb{N}$ series. We define more generally their product as the series
 $\prod_{i=1}^{k}S_{i}=\sum_{\overline{x}\in\overline{X}}h_{1}(x_{1})h_{2}(x_{2})\dots h_{k}(x_{k})\,$,}
@@
 Proposition 2.6
-If the series $X_{1}$, $\dots$, $X_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has the sum $s_{1}s_{2}\dots s_{k}$.
+If the series $S_{1}$, $\dots$, $S_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
 Interestingly, a reversal of Proposition 2.5 holds.
 Corollary 2.7
-If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$
+If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if their product $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
 Proof.
 This follows from Proposition 2.5 if we show that $S_{1}$ and $S_{2}$ absolutely converge.
-Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$.
-$i=1,2$, and let $c>0$ be such that for every finite set $X\subset X_{1}\times X_{2}$,
+Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, and let $c>0$ be such that for every finite set $X\subset X_{1}\times X_{2}$ we have
 $\sum_{\langle x,y\rangle\in X}|h_{1}(x)h_{2}(y)|\leq c\,.$}
-We prove absolute convergence of $S_{1}$. We take any $y\in X_{2}$ with $h_{2}(y)\neq 0$. Then for every finite set $X\subset X_{1}$,
+We prove absolute convergence of $S_{1}$. We take any $y\in X_{2}$ with $h_{2}(y)\neq 0$. Then for every finite set $X\subset X_{1}$ we have
 $\sum_{x\in X}|h_{1}(x)|=\frac{1}{|h_{2}(y)|}\sum_{\langle x,y\rangle\in X\times\{y\}}|h_{1}(x)h_{2}(y)|\leq\frac{c}{|h_{2}(y)|}\,.$}
-Thus $S_{1}$ absolutely converges. For $S_{2}$ we proceed similarly. $\Box$
-It is clear that the assumption on $S_{1}$ and $S_{2}$ cannot be omitted.
+Thus $S_{1}$ absolutely converges. For $S_{2}$ we argue similarly. $\Box$
+It is clear that the assumptions on $S_{1}$ and $S_{2}$ cannot be omitted.
 Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$.
-We obtain a sufficient condition for the lightness of $h$ in terms of the values $h(e)$.
-We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ that can be reached from $1$ by a walk of length at most $n$ that has edges with nonzero weights.
+We obtain a sufficient condition for the lightness of the weight $h$ in terms of its values.
+We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk of length at most $n$ and with nonzero weight.
 Definition 2.8 (slimness)
@@
 We may assume that walks in $X$ have only edges with nonzero weights.
-We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in W(n-1)$ and $\ell(w)$ is the last vertex of $w$.
+We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in W(n-1)$ and $\ell(w)$ is the last vertex.
 Let $Y$ be the set of walks $w^{\prime}$.
@@
 $X(w^{\prime})=\{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}\colon\;w^{\prime}v\in X\}\,.$
-Using finite versions of Propositions 2.4 and 2.2 we get the bound
+We get the bound
 ${\textstyle\sum_{w\in X}|h(w)|}$ $=$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot\sum_{v\in X(w^{\prime})}|h(\{\ell(w^{\prime}),\,v\})|}$
@@
 $\Box$
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We define
-$V(h)=\bigcup_{n\geq 0}V(h,\,n)\,.$}
-It is the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk formed by edges with nonzero weights. We define the graph
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We define $V(h)=\bigcup_{n\geq 0}V(h,n)$. It is the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by walks with nonzero weights. We define the graph
 $G(h)=\langle V(h),\,E(h)\rangle$
-as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. It follows that if $G(h)$ has finite degrees then $h$ is slim.
+as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. It follows that if $G(h)$ has finite degrees then $h$ is slim and hence light.
 Definition 2.10 (convexity)
@@
 absolutely converges.
-Proof. There is a walk $w_{0}=\langle u_{0},\dots,u_{n}\rangle$ in $G(h)$ such that $u_{0}=1$ and $u_{n}=u$. Let
-$c=h(w_{0})=\prod_{i=1}^{n}h(\{u_{i-1},u_{i}\})\ \ (\neq 0)\,.$}
-Let $W$ be the set of walks $\langle u_{0},\dots,u_{n},u_{n+1}\rangle$ in $K_{\mathbb{N}}$.
+Proof.
+There is a walk $w_{0}=\langle u_{0},\dots,u_{n}\rangle$ in $G(h)$ such that $u_{0}=1$ and $u_{n}=u$.
+Let $c=h(w_{0})$ ($\neq 0$) and $W$ be the set of walks $\langle u_{0},\dots,u_{n},u_{n+1}\rangle$ in $K_{\mathbb{N}}$.
 Thus vertices $u_{0}$, $\dots$, $u_{n}$ are fixed and $u_{n+1}$ runs in $\mathbb{N}\setminus\{u_{n}\}=\mathbb{N}\setminus\{u\}$.
@@
 absolutely converges. $\Box$
-We survey the results on generating functions, which we use in the next two sections. We work with univariate formal power series
+We survey some results on generating functions. We work with univariate formal power series
 […]
@@
 $U(x)+V(x)=\sum_{n\geq 0}(u_{n}+v_{n})x^{n},\ U(x)\cdot V(x)=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
-Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$, and the units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with $a_{0}\neq 0$.
+Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with $a_{0}\neq 0$.
 Definition 2.12 ($U(1)_{1}$ and $U(1)_{2}$)
@@
     absolutely converges, we denote the sum by $U(1)_{2}$ and say that $U(1)_{2}$ exists.
-If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$, but the opposite implication is, in general, not true. For example, for
+If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$, but the opposite implication is, in general, not true. For example, if
 $U(x)=\sum_{n\geq 1}(-1)^{n-1}\frac{1}{n}x^{n}$}
-$U(1)_{1}$ exists and $U(1)_{1}=\log 2$, but $U(1)_{2}$ does not exist. But if
+then $U(1)_{1}$ exists and $U(1)_{1}=\log 2$, but $U(1)_{2}$ does not exist. But if
 $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]$}
@@
 Proposition 2.13
-If $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ and $U(1)_{1}$ exists, then for every number $x\in\mathbb{C}$ with $|x|<1$ the power series
-$\sum_{n\geq 0}u_{n}x^{n}$}
-absolutely converges.
+If $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ and $U(1)_{1}$ exists, then the power series $\sum_{n\geq 0}u_{n}x^{n}$ absolutely converges for every $x\in\mathbb{C}$ with $|x|<1$.
+In fact, for the absolute convergence it suffices (and is necessary) by Hadamard’s formula that $\limsup_{n\to\infty}|u_{n}|^{1/n}\leq 1$.
 Definition 2.14 ($U(1)_{1}=+\infty$)
@@
 Proof. Since $\frac{1}{1-x}=1+x+x^{2}+\dots$, we have $u_{n}=v_{0}+v_{1}+\dots+v_{n}$. $\Box$
-Quantities $U(1)_{1}$ are handled by means of the classical theorem of Abel, see [1] and [27, Chapter II.7]. We state it for the radius of convergence $R\geq 1$.
+Quantities $U(1)_{1}$ are handled by means of the classical theorem due to N. Abel, see [1] and [28, Chapter II.7]. We state it for the radius of convergence $R\geq 1$.
 Theorem 2.16 (Abel’s 1)
@@
 Proposition 2.19
-Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$ for $n\in\mathbb{N}$. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
+Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$, $n\in\mathbb{N}$. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
 $U(1)_{1}=\prod_{i=1}^{n}V_{i}(1)_{1}\,.$}
-Proof. By Propositions 2.4, 2.6 and 2.13, and the assumptions,
-$\text{$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}$ with $|x|<1$}\,.$}
-Theorem 2.16 and properties of limits of functions give
+Proof. By Propositions 2.4, 2.6 and 2.13, and by the assumptions,
+$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}$ with $|x|<1$
+(see proofs of Propositions 2.25 and 2.27 for details of similar arguments). Theorem 2.16 and properties of limits of functions give
 $U(1)_{1}$ $=$ $\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{U}(x)=\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}{\textstyle\prod_{i=1}^{n}F_{V_{i}}(x)}$
@@
 In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$.
-By Propositions 2.2, 2.4, 2.5 and 2.13, $F_{U}(x)F_{V}(x)=1$ for $x\in[0,1)$.
+By Propositions 2.4 and 2.5, $F_{U}(x)F_{V}(x)=1$ for $x\in[0,1)$.
 Thus $F_{V}(x)\neq 0$ and $F_{U}(x)=\frac{1}{F_{V}(x)}$ for every $x\in[0,\,1)$.
@@
 Let T(x)=W(x)V(x)T(x)=W(x)-V(x).
-By Proposition 2.18, the sum $T(1)_{1}$ exists and $T(1)_{1}=W(1)_{1}-V(1)_{1}$.
+By Proposition 2.18, $T(1)_{1}$ exists and $T(1)_{1}=W(1)_{1}-V(1)_{1}$.
 By Proposition 2.19 we have $(U(1)_{1})^{2}\cdot W(1)_{1}=T(1)_{1}=W(1)_{1}-V(1)_{1}$.
@@
 Theorem 2.17 and properties of limits of functions give
-$(U(1)_{1})^{2}$ $=$ ${\textstyle\lim_{x\to 1}F_{U}(x)^{2}=\lim_{x\to 1}\big(1-\frac{F_{V}(x)}{F_{W}(x)}\big)=1-\frac{\lim_{x\to 1^{-}}F_{V}(x)}{\lim_{x\to 1^{-}}F_{W}(x)}}$
-$=$ ${\textstyle 1-\frac{a}{\lim_{n\to\infty}\sum_{j=0}^{n}w_{j}}=1-\frac{a}{+\infty}=1}\,.$
+$(U(1)_{1})^{2}$ $=$ ${\textstyle\big(\lim_{x\to 1}F_{U}(x)\big)^{2}=\lim_{x\to 1}F_{U}(x)^{2}=\lim_{x\to 1}\big(1-\frac{F_{V}(x)}{F_{W}(x)}\big)}$
+$=$ ${\textstyle 1-\frac{\lim_{x\to 1}F_{V}(x)}{\lim_{x\to 1}F_{W}(x)}=1-\frac{a}{\lim_{n\to\infty}\sum_{j=0}^{n}w_{j}}=1-\frac{a}{+\infty}=1}\,.$
 Hence $U(1)_{1}=1$ because $U(1)_{1}\geq 0$. $\Box$
-We proceed to absolute convergence.
+We obtain analogs of the previous results in absolute convergence.
 Proposition 2.24
@@
 $S=\sum_{\langle j,\,k\rangle\in\mathbb{N}_{0}^{2}}v_{j}w_{k}=\sum_{j\in\mathbb{N}_{0}}v_{j}\cdot\sum_{k\in\mathbb{N}_{0}}w_{k}\,.$}
-It absolutely converges by Proposition 2.5 and has the sum $V(1)_{2}W(1)_{2}$.
+By Proposition 2.5 it absolutely converges and has sum $V(1)_{2}W(1)_{2}$.
 On the other hand, $S$ has the grouping $\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}$ (the inner $\sum$ means sum and the outer $\sum$ means series) and $\sum_{j=0}^{n}v_{j}w_{n-j}=u_{n}$.
@@
 $\Box$
-Propositions 2.19 for $n=2$ and item 1 of 2.25 are incomparable.
+Proposition 2.19 for $n=2$ and item 1 of Proposition 2.25 are incomparable.
 The former assumes only conditional convergence, but for all three series.
 The latter assumes stronger absolute convergence, but only for two series.
-Another variant of the proposition is the theorem due to Franz (Franciszek) Mertens (1840–1927).
+Another variant of the proposition is due to Franz (Franciszek) Mertens (1840–1927).
 Theorem 2.26 (Mertens’s)
-Let $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$ and if $V(1)_{1}$ and $W(1)_{2}$ exist, then $U(1)_{1}$ exists and
+Suppose that $U(x)$, $V(x)$ and $W(x)$ are in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$, and if $V(1)_{1}$ and $W(1)_{2}$ exist, then $U(1)_{1}$ exists and
 $U(1)_{1}=V(1)_{1}W(1)_{2}\,.$
@@
 For $m\in\mathbb{N}$ we consider the series
-$S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}\cdot\ldots\cdot v_{j_{m}}$}
-By Proposition 2.6 it absolutely converges and has the sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has the sum $1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}$.
+$S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}v_{j_{2}}\dots v_{j_{m}}\,.$}
+By Proposition 2.6 it absolutely converges and has sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has sum $1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}$.
 On the other hand, in $\mathbb{C}[[x]]$ we have the identity
@@
 $u_{n}=\sum_{m\geq 1}(-1)^{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=n\end{subarray}}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\,.$}
-By Proposition 2.4, $S$ has also the sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$ and the stated formula follows.
+By Proposition 2.4, $S$ has also sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$ and the stated formula follows.
 2. Indeed,
@@
 $|v|+|v|\sum_{m\geq 1}(|t_{1}|+|t_{2}|+\dots)^{m}=\frac{|v|}{1-c}\,.$}
-By Proposition 2.4, $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.25, $U(1)_{2}$ exists and $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
+By Proposition 2.4, $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.25, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
 Proposition 2.30
@@
 $\Box$
-Finally, we consider light convex weights.
+Finally, we consider convex light weights.
 Proposition 2.31
@@
 $=$ ${\textstyle\sum_{w^{\prime}\in W(n-1)}h(w^{\prime})\cdot 1=1\,.}$
-We can add and remove the condition $\ell(w^{\prime})\in V(h)$ without affecting the sum because walks $w^{\prime}$ not satisfying it have zero weights. $\Box$
+We can add and remove the condition $\ell(w^{\prime})\in V(h)$ without affecting the sum because walks $w^{\prime}$ not satisfying it have zero weight. $\Box$
 3 Extensions of Pólya’s theorem with $v=1$
-In this section we establish the first group of generalizations of Pólya’s theorem, namely Theorems 3.23.7.
-We treat the case $v=1$ when the vertex to be visited by walks coincides with their starting vertex.
-We introduce some generating functions in $\mathbb{C}[[x]]$.
+We establish the first group of generalizations of Pólya’s theorem in Theorems 3.23.7. We consider the case $v=1$ when the vertex visited by walks coincides with their starting vertex.
+We introduce some generating functions.
 Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight.
-Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the set of walks that visit $v$ at a step $i>0$.
-In this section, $v=1$.
+Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the set of walks that visit $v$ in step $i>0$.
 We are interested in the generating function
-$A_{h}(x)=\sum_{n\geq 0}a_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(1,\,n))x^{n}\,.$}
-We also define
+$A_{h}(x)=\sum_{n\geq 0}a_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(1,\,n))x^{n}$}
+and define
 $B_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h}x^{n}=\sum_{n\geq 0}h(W_{b}(n))x^{n}\,,}$
@@
 $D_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}d_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(n))x^{n}\,,}$
-where $W_{b}(n)$ ($\subset W(n)$) are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and $W_{c}(n)$ ($\subset W_{b}(n)$) are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$ but avoid $1$ between.
-The coefficients of $A_{h}(x)$, $\dots$, $D_{h}(x)$ are correctly defined because they are sums of subseries of absolutely convergents series.
+all in $\mathbb{C}[[x]]$.
+Here $W_{b}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and $W_{c}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and avoid $1$ between.
+The coefficients $a_{n}^{h}$, $\dots$, $d_{n}^{h}$ are correctly defined because they are sums of subseries of absolutely convergents series.
 We have $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$.
@@
 where $0=m_{0}<m_{1}<\dots<m_{j}=n$ and $u_{k}=1$ iff $k=0$ or $k=m_{i}$ for some $i\in[j]$, we set
-$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{i}),\,n_{i}=m_{i}-m_{i-1})$
+$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{1})\times W_{c}(n_{2})\times\dots\times W_{c}(n_{j}),\,n_{i}=m_{i}-m_{i-1})$
 where $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i\in[j]$.
@@
 $\Box$
-For better orientation we label our theorems by the triples
+For better orientation we label the fourteen theorems by the triples
 $\langle x,\,y,\,z\rangle\in\{v=1,\,v\neq 1\}\times\{\mathrm{gen},\,\mathrm{con}\}\times\{U(1)_{1},\,U(1)_{2},\,U(1)_{1}=+\infty\}\,$,
-where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to convex ones. The generating functions $A_{h}(x)$, $\dots$, $D_{h}(x)$ are as above
+where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to convex ones. The generating functions $A_{h}(x)$, $\dots$, $D_{h}(x)$ are as above.
 Theorem 3.2 (1: $v=1$, gen, $U(1)_{1}$)
@@
 Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight.
-Suppose that $D_{h}(1)_{2}$ and $C_{h}(1)_{2}$ exist, and that $\sum_{n\geq 1}|b_{n}^{h}|<1$.
-Then $B_{h}(1)_{2}$ exists and is nonzero, $A_{h}(1)_{2}$ exists and
+Suppose that $D_{h}(1)_{2}$ exists and $\sum_{n\geq 1}|b_{n}^{h}|<1$.
+Then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $A_{h}(1)_{2}$ exists and
 $A_{h}(1)_{2}=\big(1-\frac{1}{B_{h}(1)_{2}}\big)D_{h}(1)_{2}\,.$}
 Proof.
+By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$.
+By item 1 of Proposition 2.27, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h}(1)_{2}$ exists and $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$.
 By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$.
 By item 1 of Proposition 2.25, $A_{h}(1)_{2}$ exists and $A_{h}(1)_{2}=C_{h}(1)_{2}D_{h}(1)_{2}$ .
-By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$.
-By item 1 of Proposition 2.27, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h}(1)_{2}$ exists and $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$.
 The stated formula follows.
@@
 Theorem 3.4 (3: $v=1$, gen, $U(1)_{1}=+\infty$)
-Suppose that $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ is a light weight and $D_{h}(1)_{1}=+\infty$. Then
+Suppose that $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ is a light weight such that $D_{h}(1)_{1}=+\infty$. Then
 $A_{h}(1)_{1}=+\infty\,.$
@@
 Theorem 3.7 (6: $v=1$, con, $U(1)_{1}=+\infty$)
-Let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a convex light weight and $B_{h}(1)=+\infty$. Then
+Let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a convex light weight such that $B_{h}(1)=+\infty$. Then
 $\lim_{n\to\infty}a_{n}^{h}=1\,.$
@@
 4 Extensions of Pólya’s theorem with $v\neq 1$
-The case $v\neq 1$ when the vertex to be visited by walks differs from their starting vertex is more complicated than the previous case.
-The grid graph $\mathbb{Z}^{d}$ is vertex-transitive but so far this was irrelevant.
-For $\overline{v}\neq\overline{0}$ vertex-transitivity becomes relevant and we generalize it to $K_{\mathbb{N}}$ as follows.
+The case when the vertex visited by walks differs from their starting vertex is more complicated.
+The grid graph $\mathbb{Z}^{d}$ is vertex-transitive, but so far, this has been irrelevant.
+For $\overline{v}\neq\overline{0}$ vertex-transitivity becomes important, and we generalize it to $K_{\mathbb{N}}$ as follows.
 Let $v\in\mathbb{N}\setminus\{1\}$ with $v\neq 1$.
@@
 We again introduce some generating functions.
-Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$ with $v\neq 1$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight.
-Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ are the walks visiting $v$ ($\neq 1$) at a step $i>0$.
+Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight.
+Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ are the walks visiting the vertex $v$ ($\neq 1$); necessarily in a step $i>0$.
 We are interested in the generating function
 $A_{h,v}(x)=\sum_{n\geq 0}a_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W(v,\,n))x^{n}\,.$}
-The generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ are as before. We additionally introduce
+The generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ are as before. We define
 $B_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{b,v}(n))x^{n}\,,}$
 $C_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{c,v}(n))x^{n}\,\text{ and}}$
-$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e,v}(n))x^{n}\,.}$
+$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e,v}(n))x^{n}\,,}$
+all in $\mathbb{C}[[x]]$.
 Here $W_{b,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$.
@@
 Proposition 4.1
-Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and the generating functions $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ in $\mathbb{C}[[x]]$ be as above.
+Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and generating functions $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ be as above.
 The following relations hold between them.
@@
 1.
-Every walk $w$ in $\bigcup_{n\geq 0}W(v,n)$ splits at the first visit of $v$ in two walks as $w=w_{1}w_{2}$.
+Every walk $w$ weight-counted by $A_{h,v}(x)$ splits at the first visit of $v$ as $w=w_{1}w_{2}$.
 Walks $w_{1}$ are weight-counted by $C_{h,v}(x)$ and $w_{2}$ are arbitrary walks starting at $v$.
 By the $v$-transitivity of $h$, walks $w_{2}$ are weight-counted by $D_{h}(x)$.
-The first relation follows.
 2.
@@
 Those avoiding $v$ are weight-counted by $B_{h,v}(x)$.
-If $w$ visits $v$, then it uniquely splits at the first and last visits of $v$ in three walks as $w=w_{1}w_{2}w_{3}$.
+If $w$ visits $v$, then it uniquely splits at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$.
 Walks $w_{1}$ start at $1$, end at $v$ and between avoid $v$.
@@
 We reverse walks $w_{3}$ and see that both $w_{1}$ and $w_{3}$ are weight-counted by $C_{h,v}(x)$.
-By the $v$-transitivity of $h$, the middle walks $w_{2}$ are weight-counted by $B_{h}(x)$.
-The second relation follows.
+Due to the $v$-transitivity of $h$, the middle walks $w_{2}$ are weight-counted by $B_{h}(x)$.
 3.
@@
 They start at $1$, end at $v$, and between avoid $v$.
-These walks uniquely split at the last visit of $1$ in two walks as $w=w_{1}w_{2}$.
+These walks uniquely split at the last visit of $1$ as $w=w_{1}w_{2}$.
 Walks $w_{1}$ are weight-counted by $B_{h,v}(x)$, and $w_{2}$ by $E_{h,v}(x)$.
-The third relation follows.
 $\Box$
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $v\in\mathbb{N}\setminus\{1\}$. If $v\not\in V(h)$ then
-$a_{n}^{h,\,v}=h(W(v,\,n))=\sum_{w\in W(v,\,n)}h(w)=\sum_{w\in W(v,\,n)}0=0$}
-for every $n\in\mathbb{N}_{0}$ because every walk starting at $1$ and ending at $v$ contains an edge with zero weight.
-Henceforth we therefore assume that $v\in V(h)$.
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $v\in\mathbb{N}\setminus\{1\}$.
+If $v\not\in V(h)$ then $a_{n}^{h,v}=\sum_{w\in W(v,n)}h(w)=\sum_{w\in W(v,n)}0=0$ for every $n\in\mathbb{N}_{0}$ because every walk starting at $1$ and visiting $v$ contains an edge with zero weight.
+Henceforth, we therefore assume that $v\in V(h)$.
 For $w,z\in\mathbb{C}$ we define the set
@@
 $A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,$,
-where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret the fraction $\frac{0}{0}$ as $1$, so that $A_{h,v}(1)_{1}=0$.
+where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret $\frac{0}{0}$ as $1$, so that $A_{h,v}(1)_{1}=0$.
 Proof.
@@
 $\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1$}
-then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ and $A_{h,v}(1)_{2}$ exist, and
+then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $A_{h,v}(1)_{2}$ exists and
 $A_{h,\,v}(1)_{2}\in{\textstyle D_{h}(1)_{2}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
@@
 Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
-Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed inequality hold.
+Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold.
 By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
-By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $C_{h,v}(1)_{2}^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$.
+By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $(C_{h,v}(1)_{2})^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$.
 By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$.
@@
 $\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\ \ (\subset\mathbb{C})\,$,
-where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret the fraction $\frac{0}{0}$ as $1$, so that the limit is zero.
+where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret $\frac{0}{0}$ as $1$, so that the limit is zero.
 Proof.
@@
 $\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1$}
-then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ and $A_{h,v}(1)_{2}$ exist, and
-$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle D_{h}(1)_{2}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
+then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$ and
+$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
 Proof.
@@
 Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
-Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed inequality hold.
+Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold.
 By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
@@
 Theorem 4.9 (12: $v\neq 1$, con, $U(1)_{1}=+\infty$)
-Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive convex light weight such that $v\in V(h)$. Suppose that $B_{h}(1)_{1}=+\infty$. Then
+Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive convex light weight such that $v\in V(h)$. If $B_{h}(1)_{1}=+\infty$ then
 $\lim_{n\to\infty}a_{n}^{h,v}=1\,.$
@@
 Thus $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$.
-From the facts that $F_{C_{h,v}}(0)=0$, $F_{C_{h,v}}(x)$ increases and is continuous and $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$, we obtain a (unique) number $x_{0}\in(0,1)$ such that $F_{C_{h,v}}(x)<1$ for $x\in[0,x_{0})$ and $\lim_{x\to x_{0}^{-}}F_{C_{h,v}}(x)=1$.
+Since $F_{C_{h,v}}(0)=0$, $F_{C_{h,v}}(x)$ increases and is continuous and $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$, we obtain a (unique) number $x_{0}\in(0,1)$ such that $F_{C_{h,v}}(x)<1$ for $x\in[0,x_{0})$ and $\lim_{x\to x_{0}^{-}}F_{C_{h,v}}(x)=1$.
 By item 2 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality
@@
 where the last integral domain of formal power series in $k$ variables is endowed with the usual non-Archimedean norm.
-In [16] we will work out some concrete examples, both for complex and formal weights.
-The main novelty of our generalization is this. If we return to the equality
+The main novelty of our generalization is this. In an equality like
 $P_{d}(\overline{v},\,n)=h(W(v,\,n))$
-in the Introduction, the left-hand side involves, in the standard rigorous probabilistic treatment, an implicit probability space that supports the corresponding Markov chain.
+in the Introduction, the left-hand side involves, in the standard rigorous probabilistic treatment, an implicit probability space supporting the corresponding Markov chain.
 Construction of this probability space is far from trivial; see, for example, [6].

### v6

@@
 Using the semiformal method in combinatorics we generalize Pólya’s theorem.
-This theorem determines the limiting probability that walks in the grid graph $\mathbb{Z}^{d}$ visit the given vertex.
+This theorem determines limits of probabilities that walks in the grid graph $\mathbb{Z}^{d}$ starting at the origin visit the given vertex.
 We generalize it to the countable complete graph $K_{\mathbb{N}}$ with edge weights in $\mathbb{C}$.
-In part II we treat edge weights in $\mathbb{C}[[x_{1},\dots,x_{k}]]$.
+In part II we consider analogs of Pólya’s theorem for edge weights in $\mathbb{C}[[x_{1},\dots,x_{k}]]$.
+Besides the abstract, we want to help the reader understand our article by the following preamble. In it we have more space to explain what we do.
+We mention Pólya’s theorem, which is a well known result in the discipline of random walks.
+However, our article is not really about generalizing Pólya’s theorem.
+We use extensions of Pólya’s theorem to edge weighted countable complete graphs merely as a test case for the novel generalization of the well known ([3, 8, 10, 12, 22, 23, 28]) symbolic method in enumerative (analytic) combinatorics.
+This generalization is due to the first author, and we call it the semiformal method.
+It will be developed in detail in [14].
+The present article is an explanation of the functioning of the semiformal method in the form of an extensive example.
+More applications of it are given in [15].
+The idea of the semiformal method is to extend finite sums defining coefficients of generating functions in the symbolic method to infinite (countable) series.
+Finite sums in a ring (in the symbolic method, one also needs to multiply) are straightforward, but it is less clear what is a sum of an infinite series of real or complex numbers.
+We use three kinds of sums of infinite series: absolute convergence (which is used in probability where summands are nonnegative), the more general and weaker conditional convergence, and sums equal to $\pm\infty$.
+The last sums are in fact important — the heart of the combinatorial proof of the case $d\leq 2$ in Pólya’s theorem is the fact that a certain series has sum $\infty$.
+For details on the uses of these three kinds of sums we refer to the article below.
+But why confine ourselves to infinite series of real or complex numbers, when non-Archimedean infinite series are also interesting?
+This is the topic of [17].
+To summarize, the main contribution of this article is the demonstration of the working of the semiformal method in the fourteen cases that arise when Pólya’s theorem is generalized to $K_{\mathbb{N}}$ with complex edge weights.
 1 Introduction
@@
 In 1921, G.
-Pólya proved in [24] a theorem, which we generalize, asserting that for every vertex $\overline{v}$,
+Pólya proved in [24] a theorem, which we generalize, that for every vertex $\overline{v}$,
 $\lim_{n\to\infty}P_{d}(\overline{v},\,n)\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ <1&\dots&\text{if $d\geq 3$}\,.\end{array}\right$.
-Pólya’s theorem is almost always cited in the special case with $\overline{v}=\overline{0}$, although he proved it for any vertex $\overline{v}$ (see [16, Section 1] for a translated quote from [24]).
-A selection of articles and books discussing Pólya’s theorem is [3, 4, 6, 9, 11, 13, 18, 19, 20, 21, 25, 26, 29, 30].
+Pólya’s theorem is almost always mentioned in the special case with $\overline{v}=\overline{0}$, but he proved it for any vertex $\overline{v}$ (see [16, Section 1] for a quote where he states his result).
+A selection of articles and books discussing Pólya’s theorem is [4, 5, 6, 9, 11, 13, 18, 19, 20, 21, 25, 26, 27, 31, 32].
 We generalize it as follows. Let
-$\mathbb{N}_{2}=\binom{\mathbb{N}}{2}=\{e\colon\;e\subset\mathbb{N}\wedge|e|=2\}$}
-be edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$ on the vertex set $\mathbb{N}$.
-For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1.\dots\}$) its cardinality.
-Let
+$\mathbb{N}_{2}=\binom{\mathbb{N}}{2}=\{e\colon\;e\subset\mathbb{N}\text{ with }|e|=2\}$}
+be edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$. Here $|A|$ ($\in\mathbb{N}_{0}=\{0,1.\dots\}$) denotes the cardinality of a finite set $A$. Let
 $h\colon\mathbb{N}_{2}\to\mathbb{C}$
-be a complex edge weight.
-$\mathbb{C}$ denotes the field of complex numbers, which is endowed with the usual Archimedean absolute value $|z|=\sqrt{z\cdot\overline{z}}$.
-For $n\in\mathbb{N}_{0}$ let $W(n)$ be the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at the vertex $1$.
+be an edge weight, where $\mathbb{C}$ denotes the field of complex numbers, which is endowed with the usual Archimedean absolute value $|z|=\sqrt{z\cdot\overline{z}}$.
+Let $n\in\mathbb{N}_{0}$.
+We denote by $W(n)$ the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$.
 Thus $W(0)=\{\langle 1\rangle\}$ and for $n>0$ the set $W(n)$ is infinite and countable.
-For $v\in\mathbb{N}$ let $W(v,n)$ be the subset of $W(n)$ of the walks that visit the vertex $v$ in a step $i>0$.
-We extend the weight in the usual multiplicative way to
+Let $v\in\mathbb{N}$.
+We denote by $W(v,n)$ ($\subset W(n)$) the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$ and visiting $v$ in a step $i>0$.
+We extend the weight multiplicatively to
 $h\colon\bigcup_{n=0}^{\infty}W(n)\to\mathbb{C}\,.$}
@@
 (Results on series are reviewed in the next section.) Suppose that for every $n\in\mathbb{N}_{0}$ the series $\sum_{w\in W(n)}h(w)$ absolutely converges.
-Then so does $\sum_{w\in W(v,n)}h(w)$ and we consider the generating function
+Then so does the series $\sum_{w\in W(v,n)}h(w)$ and we consider the generating function
 $A_{h,v}(x)=\sum_{n\geq 0}h(W(v,\,n))x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
-Our generalizations of Pólya’s theorem are formulas for the quantities
-$A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))\,.$}
+Our generalizations of Pólya’s theorem are formulas for three quantities:
+$\lim_{x\nearrow 1}A_{h,v}(x),\ A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))\,$,}
+where $\lim_{x\nearrow 1}$ means that $x$ goes to $1$ through values $x\in[0,1)$.
+In some cases, these quantities are twisted by powers $\xi^{n}$ for $\xi\in\mathbb{C}^{*}$ ($\mathbb{C}\setminus\{0\}$).
 The formulas appear in fourteen theorems: Theorems 3.23.7 in Section 3 and Theorems 4.34.10 in Section 4.
@@
 Theorems 3.5 and 3.7 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}=\overline{0}$.
-Theorems 4.6 (4.7) and 4.9 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
-We have fourteen generalizations of one theorem for three reasons.
-First, in Pólya’s theorem weights are nonnegative real numbers and convergence of infinite series is automatically absolute.
-For complex weights we consider both kinds of convergence, absolute and conditional one.
-We also consider, for nonnegative real weights, series with sum $\infty$.
-Second, Pólya’s theorem lives in the probabilistic setup where weights of edges incident to a fixed vertex sum up to $1$.
-We consider the case when this condition is kept (convex weights) and the case when it is dropped.
-Third, already Pólya’s theorem is in reality two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in complexity of proofs.
-In our generalization we distinguish cases $v=1$ and $v\neq 1$ as well.
-Now $3\cdot 2\cdot 2=12$; two more generalizations arise by variations of assumptions.
-We explain how our approach includes Pólya’s original problem. If $f\colon A\to B$ is a map and $C$ is any set (not necessarily a subset of $A$), we call the set
+Theorems 4.7 (4.8) and 4.10 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
+The fourteen generalizations come from three sources.
+First, weights in Pólya’s theorem are nonnegative real numbers, and the convergence of infinite series is absolute.
+For our complex weights, we consider two kinds of convergence: absolute and conditional.
+We also consider real series with sums $\pm\infty$.
+Second, Pólya’s theorem is on probabilities, which means that the weights of edges incident to a fixed vertex sum up to $\xi=1$.
+We generalize it to $\xi$-convexity for any $\xi\in\mathbb{C}$; the case of $\xi\neq 0$ is normalized to $|\xi|=1$ and the case $\xi=0$ is treated separately (it corresponds to Kirchhoff’s first law).
+Thus we consider two cases, general weights and $\xi$-convex weights.
+Third, Pólya’s theorem is, in reality, two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in the complexity of proofs.
+In our generalization, we correspondingly distinguish between cases $v=1$ and $v\neq 1$.
+Now $3\cdot 2\cdot 2=12$ and two additional generalizations arise by varying the assumptions.
+We explain how our approach includes Pólya’s original problem. If $f\colon A\to B$ is a map and $C$ is any set (not necessarily a subset of $A$), we denote by
 $f[C]=\{f(x)\colon\;x\in A\cap C\}\ \ (\subset B)$
@@
 We take any bijection $f\colon\mathbb{Z}^{d}\to\mathbb{N}$ such that $f(\overline{0})=1$, and set $v=f(\overline{v})$.
-The grid graph on $\mathbb{Z}^{d}$, which is the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
+The grid graph $\mathbb{Z}^{d}$, which is the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
 $\{\overline{a},\,\overline{b}\}\in E_{d}\iff\sum_{i=1}^{d}|a_{i}-b_{i}|=1\,$,}
-determines the weight $h\colon\mathbb{N}_{2}\to\{\frac{1}{2d},0\}$ by the relation $h(e)\neq 0$ iff $e=f[e^{\prime}]$ for some $e^{\prime}\in E_{d}$. Then for every $n\in\mathbb{N}_{0}$ we have
+determines a weight $h\colon\mathbb{N}_{2}\to\{\frac{1}{2d},0\}$ by the relation $h(e)\neq 0$ iff $e=f[e^{\prime}]$ for $e^{\prime}\in E_{d}$. Then for every $n\in\mathbb{N}_{0}$ we have
 $P_{d}(\overline{v},\,n)=h(W(v,\,n))\,.$
+See Section 5 for the importance of (the general form of) this equality.
 We structure our article as follows.
-In Section 2 we review infinite series with complex summands and generating functions.
+In Section 2 we review infinite series with complex summands and the domain $\mathbb{C}[[x]]$ of generating functions.
 Sections 3 and 4 contain our main results, the mentioned fourteen generalizations of Pólya’s theorem.
 Section 5 contains concluding remarks.
-We continue our investigation in [17] (see Section 5 for more details).
+We continue our investigation in [17]; see Section 5 for more details on [17].
 In Sections 3 and 4 we use what we term the semiformal method in (enumerative) combinatorics.
@@
 In our situation such extension is natural and inevitable, coefficients in generating functions $A_{h,v}(x)$ arise as sums of infinite series $\sum_{w\in W(v,n)}h(w)$.
-We develop the semiformal method more in [14].
+We develop this method more in [14].
 Why “semiformal”?
-Generating “functions” in the symbolic method are formal power series with coefficients determined by finite operations.
-Our extension allows operations based on limit transitions.
-See [15] more applications of the semiformal method.
+Generating “functions” in the symbolic method are formal power series and their coefficients are determined by finite operations.
+We also use formal power series, but our extension allows operations based on limit transitions.
+See [15] for more applications of the semiformal method.
 2 Series and generating functions in $\mathbb{C}$
-We review series and generating functions in the domain $\mathbb{C}$.
-A series (in $\mathbb{C}$) is a map $h\colon X\to\mathbb{C}$, defined on an at most countable set $X$.
-We write it as $\sum_{x\in X}h(x)$.
+We review series and generating functions in $\mathbb{C}$.
+A series is a map $h\colon X\to\mathbb{C}$ defined on a finite or countable set $X$; we write it as $\sum_{x\in X}h(x)$.
 We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
   1. 1.
-    For some $c>0$ and every finite set $Y\subset X$ we have $\sum_{x\in Y}|h(x)|\leq c$.
+    For some constant $c\geq 0$ and every finite set $Y\subset X$ we have
+    $\sum_{x\in Y}|h(x)|\leq c\,.$}
   2. 2.
-    If $X$ is countable, then for every bijection $f\colon\mathbb{N}\to X$, the limit
-    $s=\lim_{n\to\infty}{\textstyle\sum_{i=1}^{n}h(f(i))}$
+    If the set $X$ is countable, then for every bijection $f\colon\mathbb{N}\to X$ the proper limit
+    $s=\lim_{n\to\infty}{\textstyle\sum_{i=1}^{n}h(f(i))\ \ (\in\mathbb{C})}$
     exists and does not depend on $f$.
-We call $s$ the sum of the series and denote it again by $\sum_{x\in X}h(x)$.
-If the set $X=\{x_{1},x_{2},\dots,x_{n}\}\neq\emptyset$ is finite, then every series $\sum_{x\in X}h(x)$ absolutely converges and has the sum
+(In fact, the independence of $s$ on $f$ may be omitted from the formulation.) We call $s$ the sum of the series and denote it again by $\sum_{x\in X}h(x)$.
+If the set $X$ is finite and nonempty, $X=\{x_{1},x_{2},\dots,x_{n}\}$, then every series $\sum_{x\in X}h(x)$ absolutely converges and has the sum
 $h(x_{1})+h(x_{2})+\dots+h(x_{n})\,.$
-For $X=\emptyset$ we define the sum as 0. If
-$U(x)=\sum_{n\geq 0}u_{n}x^{n}\ \ (\in\mathbb{C}[[x]])$}
-is a formal power series, so that $x$ is a formal variable, we denote for any $x\in\mathbb{C}$ by $F_{U}(x)$ the sum of the power series $\sum_{n\geq 0}u_{n}x^{n}$, if it absolutely converges.
+If $X=\emptyset$, we define the sum as 0. If
+$U=U(x)=\sum_{n\geq 0}u_{n}x^{n}\ \ (\in\mathbb{C}[[x]])$}
+is a formal power series, so that $x$ is a formal variable, we denote for any $c\in\mathbb{C}$ by $F_{U}(c)$ the sum of the power series $\sum_{n\geq 0}u_{n}c^{n}$, if it absolutely converges. Let
+$H_{1}={\textstyle\{U(x)\in\mathbb{C}[[x]]\colon\;\text{$U(c)$ absolutely converges for every $c\in\mathbb{C}_{<1}$}\}\,,}$
+where $\mathbb{C}_{<1}=\{c\in\mathbb{C}\colon\;|c|<1\}$. Thus for any $U\in H_{1}$ the sums form the function $F_{U}\colon\mathbb{C}_{<1}\to\mathbb{C}$.
 Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. A walk $w$ in $K_{\mathbb{N}}$ is any $(n+1)$-tuple $w$ of vertices $v_{i}\in\mathbb{N}$,
 $w=\langle v_{0},\,v_{1},\,\dots,\,v_{n}\rangle\,$,
-such that $n\in\mathbb{N}_{0}$ and $v_{i-1}\neq v_{i}$ for every $i\in[n]$ ($\{1,2,\dots,n\}$, $0]=\emptyset$).
-The length $n$ of $w$ is is $n$ ($\in\mathbb{N}_{0}$).
-We extend $h$ to walks.
-For $n>0$ we set
+such that $n\in\mathbb{N}_{0}$ and $v_{i-1}\neq v_{i}$ for every $i\in[n]$ ($\{1,2,\dots,n\}$, $0]=\emptyset$). The length of $w$ is $n$ ($\in\mathbb{N}_{0}$). We extend $h$ to walks. For $n>0$ we set
 $h(w)=\prod_{i=1}^{n}h(\{v_{i-1},\,v_{i}\})\,$,}
@@
 Definition 2.1 (lightness)
-A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is light if for every $n\in\mathbb{N}_{0}$ the series
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We call the weight $h$ light if for every $n\in\mathbb{N}_{0}$, the series
 $\sum_{w\in W(n)}h(w)$}
@@
 If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$.
-Any subseries of an absolutely convergent series absolutely converges, but we cannot compare sizes of their sums
+Any subseries of an absolutely convergent series absolutely converges.
+However, we cannot compare the sizes of sums
 $\big|\sum_{x\in Y}h(x)\big|\,\text{ and }\,\big|\sum_{x\in X}h(x)\big|\,.$}
 Except when $h\colon X\to\mathbb{R}_{\geq 0}$: if $\sum_{x\in X}h(x)$ absolutely converges and $Y\subset X$, then the sums satisfy
-$0\leq\sum_{x\in Y}h(x)\leq\sum_{x\in X}h(x)\,.$}
+$0\leq\sum_{x\in Y}h(x)\leq\sum_{x\in X}h(x)$}
+— we call it the subseries bound principle.
 If $R=\sum_{x\in X}g(x)$ and $S=\sum_{x\in X}h(x)$ are series and $\alpha,\beta\in\mathbb{C}$, the series
 $\alpha R+\beta S=\sum_{x\in X}(\alpha g(x)+\beta h(x))$}
-is their linear combination. We omit the easy proof of the next result.
+is the linear combination of $R$ and $S$. We omit the easy proof of the next result.
 Proposition 2.2
-If $R$ and $S$ absolutely converge and have respective sums $r$ and $s$, then series $\alpha R+\beta S$ absolutely converges and has the sum $\alpha r+\beta s$.
-The next approximation lemma is a key tool.
+If $R$ and $S$ absolutely converge and have sums $r$ and $s$, respectively, then $\alpha R+\beta S$ absolutely converges and has sum $\alpha r+\beta s$.
+The next approximation lemma is useful.
 Lemma 2.3
-If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, then for every $\varepsilon>0$ there exists a finite set $Y\subset X$, denoted by $Y(S,\,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$ we have
+If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, then for every $\varepsilon>0$ there exists a finite set $Y\subset X$, denoted by $Y(S,\,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$,
 $\big|s-\sum_{x\in Z}h(x)\big|\leq\varepsilon\,.$}
 Proof. Let an $\varepsilon>0$ be given. For finite $X$ we set $Y=X$. For countable $X$ we take any bijection $f\colon\mathbb{N}\to X$, take an $N\in\mathbb{N}$ such that
-$|\sum_{n=1}^{N}h(f(n))-s|\leq\frac{\varepsilon}{2}\,\text{ and }\,\sum_{n>N}|h(f(n))|\leq\frac{\varepsilon}{2}$}
-and set Y=f[[N]]Y=f[\,[N]\,]. Then for every finite set $Z$ with $Y\subset Z\subset X$ we have
-$|s-\sum_{x\in Z}h(x)|\leq|s-\sum_{x\in Y}h(x)|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$}
+$\big|\sum_{n=1}^{N}h(f(n))-s\big|\leq\frac{\varepsilon}{2}\,\text{ and }\,\sum_{n>N}|h(f(n))|\leq\frac{\varepsilon}{2}\,$,}
+and set $Y=\{f(1),f(2),\dots,f(N)\}$. Then for every finite set $Z$ with $Y\subset Z\subset X$ we have
+$\big|s-\sum_{x\in Z}h(x)\big|\leq\big|s-\sum_{x\in Y}h(x)\big|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$}
 $\Box$
-Two more operations with series are grouping and product.
+Two operations on series are grouping and product.
 A partition of a set $X$ is a set $P$ of nonempty and disjoint sets such that $\bigcup P=X$.
 If $X$ is at most countable, then so is $P$ and every set $Z\in P$.
-If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every set $Z\in P$ the subseries $\sum_{x\in Z}h(x)$ absolutely converges and has sum $s_{Z}$, then the series
+If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every $Z\in P$ the subseries $\sum_{x\in Z}h(x)$ absolutely converges and has sum $s_{Z}$, then the series
 $S_{P}=\sum_{Z\in P}s_{Z}$}
@@
 $\Box$
-Let $A$ and $B$ be sets. We denote their ordered pair by $\langle A,\,B\rangle$. If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, are series, then the series
+Let $X_{1}$, $X_{2}$ be sets and $X_{1}\times X_{2}$ be their Cartesian product. If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, are series, then the series
 $S_{1}\cdot S_{2}=\sum_{\langle x,\,y\rangle\in X_{1}\times X_{2}}h_{1}(x)h_{2}(y)$}
@@
 $\Box$
-Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i\in[k]$, be $k\in\mathbb{N}$ series. We define more generally their product as the series
+Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i\in[k]$, be $k\in\mathbb{N}$ series. We define their product as the series
 $\prod_{i=1}^{k}S_{i}=\sum_{\overline{x}\in\overline{X}}h_{1}(x_{1})h_{2}(x_{2})\dots h_{k}(x_{k})\,$,}
-where $\overline{x}=\langle x_{1},\dots,x_{k}\rangle$ is an ordered $k$-tuple of $x_{1}$, $\dots$, $x_{k}$ and $\overline{X}=X_{1}\times\dots\times X_{k}$.
-Iterating Proposition 2.5 we generalize it as follows.
+where $\overline{x}=\langle x_{1},\dots,x_{k}\rangle\in X_{1}\times\dots\times X_{k}=\overline{X}$. Iterating Proposition 2.5 we obtain the following result.
 Proposition 2.6
-If the series $S_{1}$, $\dots$, $S_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
+If series $S_{1}$, $\dots$, $S_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
 Interestingly, a reversal of Proposition 2.5 holds.
 Corollary 2.7
-If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if their product $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
+If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if their product $S_{1}\cdot S_{2}$ absolutely converges and has sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
 Proof.
@@
 Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$.
-We obtain a sufficient condition for the lightness of the weight $h$ in terms of its values.
-We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk of length at most $n$ and with nonzero weight.
+We obtain a sufficient condition for lightness of the weight $h$ in terms of its values.
+We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk with nonzero weight and length at most $n$.
 Definition 2.8 (slimness)
@@
 $\sum_{w\in W(n)}h(w)$}
-absolutely converges. For $n=0$ it is trivial because $W(0)=\{\langle 1\rangle\}$. Let $n>0$ and let $c>0$ be a constant such that for every finite set $X\subset W(n-1)$ we have
-$\sum_{w\in X}|h(w)|\leq c$}
-and that for every $u\in V(h,n-1)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
-$\sum_{v\in X}|h(\{u,\,v\})|\leq c\,.$}
+absolutely converges. For $n=0$ it is trivial because $W(0)=\{\langle 1\rangle\}$. Let $n>0$ and let $c>0$ be a constant such that
+$\sum_{w\in X}|h(w)|\leq c\,\text{ and }\,\sum_{v\in X^{\prime}}|h(\{u,\,v\})|\leq c$}
+for every finite set $X\subset W(n-1)$, every $u\in V(h,n-1)$ and every finite set $X^{\prime}\subset\mathbb{N}\setminus\{u\}$.
 Let $X\subset W(n)$ be finite.
-We may assume that walks in $X$ have only edges with nonzero weights.
+We may assume that all walks in $X$ have nonzero weights.
 We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in W(n-1)$ and $\ell(w)$ is the last vertex.
@@
 as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. It follows that if $G(h)$ has finite degrees then $h$ is slim and hence light.
-Definition 2.10 (convexity)
-A light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is convex if for every vertex $u\in V(h)$ the sum
-$\sum_{e\in E(h),\,e\ni u}h(e)=\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,\,v\})=1\,.$}
-We show that these sums exist.
+Definition 2.10 ($\xi$-convexity)
+Let $\xi\in\mathbb{C}$. We call a light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ $\xi$-convex if for every vertex $u\in V(h)$ the sum
+$\sum_{e\in E(h),\,e\ni u}h(e)=\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,\,v\})=\xi\,.$}
+We say that $h$ is convex if it is $1$-convex.
+We show that the sums in the definition exist.
 Proposition 2.11
@@
 absolutely converges. $\Box$
-We survey some results on generating functions. We work with univariate formal power series
-[…]
-They form an integral domain denoted by $\mathbb{C}[[x]]$. For $U(x)$ and $V(x)$ in $\mathbb{C}[[x]]$,
-$U(x)+V(x)=\sum_{n\geq 0}(u_{n}+v_{n})x^{n},\ U(x)\cdot V(x)=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
-Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with $a_{0}\neq 0$.
+We survey results on generating functions.
+We use univariate formal power series $U=U(x)=\sum_{n\geq 0}u_{n}x^{n}$ with complex coefficients.
+They form an integral domain denoted by $\mathbb{C}[[x]]$.
+For $U$ and $V$ in $\mathbb{C}[[x]]$,
+$U+V=\sum_{n\geq 0}(u_{n}+v_{n})x^{n}\,\text{ and }\,U\cdot V=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
+Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}u_{n}x^{n}$ with $u_{0}\neq 0$.
+Let $\mathbb{C}_{1}=\{x\in\mathbb{C}\colon\;|x|=1\}$ and $\xi\in\mathbb{C}_{1}$. We define
+$\xi\mathbb{R}_{\geq 0}=\{\xi x\colon\;x\in\mathbb{R}_{\geq 0}\}\ \ (\subset\mathbb{C})\,.$
+It is the ray in $\mathbb{C}$ going from the origin in the direction $\xi$.
+Clearly, $\xi\mathbb{R}_{\geq 0}$ is closed under addition.
+If $\alpha_{j}\in\xi_{j}\mathbb{R}_{\geq 0}$, $j=1,2$, then $\alpha_{1}\alpha_{2}\in(\xi_{1}\xi_{2})\mathbb{R}_{\geq 0}$.
+If
+$h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$
+is a light weight and $n\in\mathbb{N}_{0}$ then $h(W(n))\in\xi^{n}\mathbb{R}_{\geq 0}$.
+The lines $\xi\mathbb{R}$ in $\mathbb{C}$ going through the origin in the direction $\xi$ have the same properties.
+For $\alpha\in\mathbb{C}$ and $U=U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ we define
+$(U)_{\alpha}=(U)_{\alpha}(x):=U(\alpha x)=\sum_{n\geq 0}u_{n}\alpha^{n}x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
+Note that $(U+V)_{\alpha}=(U)_{\alpha}+(V)_{\alpha}$ and $(U\cdot V)_{\alpha}=(U)_{\alpha}\cdot(V)_{\alpha}$.
 Definition 2.12 ($U(1)_{1}$ and $U(1)_{2}$)
@@
   1. 1.
-    If the limit
-    $U(1)_{1}=\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}u_{j}\ \ (\in\mathbb{C})}$
+    If the (proper) limit
+    $U(1)_{1}:=\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}u_{j}\ \ (\in\mathbb{C})}$
     exists, we say that $U(1)_{1}$ exists.
@@
     absolutely converges, we denote the sum by $U(1)_{2}$ and say that $U(1)_{2}$ exists.
-If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$, but the opposite implication is, in general, not true. For example, if
+If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$. The opposite implication is, in general, not true. For example, if
 $U(x)=\sum_{n\geq 1}(-1)^{n-1}\frac{1}{n}x^{n}$}
-then $U(1)_{1}$ exists and $U(1)_{1}=\log 2$, but $U(1)_{2}$ does not exist. But if
+then $U(1)_{1}=\log 2$, but $U(1)_{2}$ does not exist. But if
 $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]$}
-and $U(1)_{1}$ exists, then $U(1)_{2}$ exists and $U(1)_{1}=U(1)_{2}$. The following result is well known.
+and $U(1)_{1}$ exists, then $U(1)_{2}$ exists and $U(1)_{1}=U(1)_{2}$. The following is a corollary of the well known Hadamard’s formula for the radius of convergence of a formal power series.
 Proposition 2.13
-If $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ and $U(1)_{1}$ exists, then the power series $\sum_{n\geq 0}u_{n}x^{n}$ absolutely converges for every $x\in\mathbb{C}$ with $|x|<1$.
-In fact, for the absolute convergence it suffices (and is necessary) by Hadamard’s formula that $\limsup_{n\to\infty}|u_{n}|^{1/n}\leq 1$.
-Definition 2.14 ($U(1)_{1}=+\infty$)
+Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. Then $U(x)\in H_{1}$ iff
+$\limsup_{n\to\infty}|u_{n}|^{1/n}\leq 1\,.$
+Thus, for example, the existence of $U(1)_{1}$ is a sufficient (but not necessary) condition for $U(x)\in H_{1}$.
+Note that if $U\in\mathbb{C}[[x]]$ then $(U)_{\xi}\in H_{1}$ either for all or no $\xi\in\mathbb{C}_{1}$.
+Definition 2.14 ($U(1)_{1}=\pm\infty$)
 If $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ is in $\mathbb{R}[[x]]$ and
-$\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}u_{j}=+\infty\,,}$
-we say that $U(1)_{1}=+\infty$.
+$\lim_{n\to\infty}\sum_{j=0}^{n}u_{j}=+\infty\,$,}
+we say that $U(1)_{1}=+\infty$. We similarly define that $U(1)_{1}=$-$\infty$.
 Proposition 2.15
-Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$. If $V(1)_{1}$ exists and $U(x)=V(x)\frac{1}{1-x}$, then
-$\lim_{n\to\infty}u_{n}=V(1)_{1}\,.$
-Proof. Since $\frac{1}{1-x}=1+x+x^{2}+\dots$, we have $u_{n}=v_{0}+v_{1}+\dots+v_{n}$. $\Box$
-Quantities $U(1)_{1}$ are handled by means of the classical theorem due to N. Abel, see [1] and [28, Chapter II.7]. We state it for the radius of convergence $R\geq 1$.
-Theorem 2.16 (Abel’s 1)
-Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. If $U(1)_{1}$ exists then the function
-$F_{U}(x)=\sum_{n\geq 0}u_{n}x^{n}\colon\{x\in\mathbb{C}\colon\;|x|<1\}\to\mathbb{C}\,$,}
-see Proposition 2.13, has the limit
-$\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{U}(x)=U(1)_{1}\,.$
-The following folklore result is in fact another variant of Abel’s theorem, incomparable with Theorem 2.16. We omit the easy proof.
-Theorem 2.17 (Abel’s 2)
-Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]$. If $U(x)$ absolutely converges for every $x\in[0,1)$, then the equality
-$\lim_{x\to 1}F_{U}(x)=\lim_{n\to\infty}\sum_{j=0}^{n}u_{j}\ \ (\in[0,\,+\infty)\cup\{+\infty\})$}
-always holds.
-Compared to Theorem 2.16, the new case is when both limits equal $\infty$.
+Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$, and let $\xi\in\mathbb{C}^{*}$. If $(V)_{1/\xi}(1)_{1}$ exists and $U(x)=V(x)\frac{1}{1-\xi x}$, then
+$\lim_{n\to\infty}\xi^{-n}u_{n}=(V)_{1/\xi}(1)_{1}\,.$
+Proof. Let $n\in\mathbb{N}_{0}$. Since $\frac{1}{1-\xi x}=1+\xi x+\xi^{2}x^{2}+\dots$, we have
+$u_{n}=v_{0}\xi^{n}+v_{1}\xi^{n-1}+\dots+v_{n}\xi^{0}=v_{0}\cdot\xi^{n}+v_{1}\xi^{-1}\cdot\xi^{n}+\dots+v_{n}\xi^{-n}\cdot\xi^{n}\,.$
+Dividing by $\xi^{n}$ and taking the limit for $n\to\infty$, we get the stated formula. $\Box$
+We manipulate the quantities $U(1)_{1}$ by means of three Abelian theorems, which we state only for the radius of convergence $R\geq 1$.
+The first theorem is classical, see [1, 2] and [29, Chapter II.7].
+Theorem 2.16 (Abel 1)
+Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. If $U(1)_{1}$ exists, so that $U(x)\in H_{1}$ by Proposition 2.13, then
+$\lim_{x\nearrow 1}F_{U}(x)=U(1)_{1}\ \ (\in\mathbb{C})\,.$
+The second Abelian theorem is rarely mentioned, but see [2] for a simple proof. It is given only for the case $U(1)_{1}=+\infty$, but it works for the sum and limit -$\infty$ as well.
+Theorem 2.17 (Abel 2)
+Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}[[x]]\cap H_{1}$. If $U(1)_{1}=\pm\infty$ then
+$\lim_{x\nearrow 1}F_{U}(x)=\pm\infty\,$,
+with equal signs.
+The third Abelian theorem allows only nonnegative real coefficients.
+It has the advantage over the previous two theorems that it is an equivalence.
+Now one can go not only from the sum to the functional limit, but also in the reverse.
+We omit the easy proof.
+Theorem 2.18 (Abel 3)
+Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]\cap H_{1}$. Then the following limit and sum always exist (and are possibly $\infty$) and are equal,
+$\lim_{x\nearrow 1}F_{U}(x)=U(1)_{1}\ \ (\in[0,\,+\infty)\cup\{+\infty\})\,.$
 Proposition 2.19
-Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{1}$ and $W(1)_{1}$ exist, then $U(1)_{1}$ exists and
+Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{1}$ and $W(1)_{1}$ exist, then
 $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}\,.$
-Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}u_{n}x^{n}$. Then
+Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$. Then
 $\sum_{j=0}^{n}u_{j}=\sum_{j=0}^{n}(\alpha v_{j}+\beta w_{j})=\alpha\sum_{j=0}^{n}v_{j}+\beta\sum_{j=0}^{n}w_{j}\,.$}
-By the limit transition $n\to\infty$ we get that that $U(1)_{1}$ exists and $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$. $\Box$
-Iterating this result we easily generalize it to $k$-term linear combinations.
+Taking the limit for $n\to\infty$ we get that that $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$. $\Box$
+Here the formulation “then $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$” abbreviates “then $U(1)_{1}$ exists and $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$”.
+We will use such abbreviated formulations.
+This result can be easily generalized to linear combinations with more than two terms.
+Next we consider products of formal power series.
 Proposition 2.20
-Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$, $n\in\mathbb{N}$. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
+Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$ be $n+1$, $n\in\mathbb{N}$, formal power series.
+If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
 $U(1)_{1}=\prod_{i=1}^{n}V_{i}(1)_{1}\,.$}
-Proof. By Propositions 2.4, 2.6 and 2.13, and by the assumptions,
-$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}$ with $|x|<1$
-(see proofs of Propositions 2.25 and 2.27 for details of similar arguments). Theorem 2.16 and properties of limits of functions give
-$U(1)_{1}$ $=$ $\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{U}(x)=\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}{\textstyle\prod_{i=1}^{n}F_{V_{i}}(x)}$
-$=$ ${\textstyle\prod_{i=1}^{n}\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{V_{i}}(x)=\prod_{i=1}^{n}V_{i}(1)_{1}}\,.$
+Proof. By Propositions 2.4, 2.6 and 2.13, and by the assumptions we have
+$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}_{<1}$
+(see proofs of Propositions 2.24 and 2.27 for details of similar arguments). Using Theorem 2.16 and properties of limits of functions we get
+$U(1)_{1}$ $=$ $\lim_{x\nearrow 1}F_{U}(x)=\lim_{x\nearrow 1}{\textstyle\prod_{i=1}^{n}F_{V_{i}}(x)}$
+$=$ ${\textstyle\prod_{i=1}^{n}\lim_{x\nearrow 1}F_{V_{i}}(x)=\prod_{i=1}^{n}V_{i}(1)_{1}}\,.$
 $\Box$
@@
 Proposition 2.21
-Let $U(x),V(x)\in\mathbb{C}[[x]]$ and $V(0)\neq 0$. If $U(x)=\frac{1}{V(x)}$ and if $U(1)_{1}$ and $V(1)_{1}$ exist, then $V(1)_{1}\neq 0$ and
+Let $U(x),V(x)\in\mathbb{C}[[x]]$. If $U(x)=\frac{1}{V(x)}$, and if $U(1)_{1}$ and $V(1)_{1}$ exist, then $V(1)_{1}\neq 0$ and
 $U(1)_{1}=\frac{1}{V(1)_{1}}\,.$
-Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. Proposition 2.19 gives that $U(1)_{1}V(1)_{1}=1$. Thus $V(1)_{1}\neq 0$ and division by $V(1)_{1}$ yields the stated formula. $\Box$
+Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. Proposition 2.20 gives that $U(1)_{1}V(1)_{1}=1$. Thus $V(1)_{1}\neq 0$ and dividing by $V(1)_{1}$ gives the stated formula. $\Box$
 Proposition 2.22
-Let $U(x),V(x)\in\mathbb{R}_{\geq 0}[[x]]$ and $V(0)\neq 0$. If $\frac{1}{V(x)}, U(x)U(x)$ and $V(x)$ absolutely converge for $x\in[0,1)$, and if $V(1)_{1}=+\infty$, then
-$U(1)_{1}=0\,.$
-Proof.
-In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$.
-By Propositions 2.4 and 2.5, $F_{U}(x)F_{V}(x)=1$ for $x\in[0,1)$.
-Thus $F_{V}(x)\neq 0$ and $F_{U}(x)=\frac{1}{F_{V}(x)}$ for every $x\in[0,\,1)$.
-Two applications of Theorem 2.17 and properties of limits of functions give
-$U(1)_{1}$ $=$ ${\textstyle\lim_{x\to 1}F_{U}(x)=\lim_{x\to 1}\frac{1}{F_{V}(x)}=\frac{1}{\lim_{x\to 1}F_{V}(x)}}$
-$=$ ${\textstyle\frac{1}{\lim_{n\to\infty\sum_{j=0}^{n}v_{j}}}=\frac{1}{+\infty}=0\,.}$
-$\Box$
-Proposition 2.22
-Let $U(x),V(x),W(x)\in\mathbb{C}[[x]]$ and $W(0)\neq 0$. If $U(x)^{2}=1-\frac{V(x)}{W(x)}$, if $U(1)_{1}$, $V(1)_{1}$ and $W(1)_{1}$ exist, and if $W(1)_{1}\neq 0$, then
+Let $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)^{2}=1-\frac{V(x)}{W(x)}$, if $U(1)_{1}$, $V(1)_{1}$ and $W(1)_{1}$ exist, and if $W(1)_{1}\neq 0$, then
 $\big(U(1)_{1}\big)^{2}=1-\frac{V(1)_{1}}{W(1)_{1}}\,.$
@@
 By Proposition 2.20 we have $(U(1)_{1})^{2}\cdot W(1)_{1}=T(1)_{1}=W(1)_{1}-V(1)_{1}$.
-Division by $W(1)_{1}$ yields the stated formula.
+Dividing by $W(1)_{1}$ we get the stated formula.
 $\Box$
-Proposition 2.23
-Let $U(x),V(x),W(x)\in\mathbb{R}_{\geq 0}[[x]]$ and $V(0),W(0)>0$.
-If $U(x)^{2}=1-\frac{V(x)}{W(x)}$, if $U(x)$, $V(x)$ and $W(x)$ absolutely converge for $x\in[0,1)$, if $F_{V}(x)$ is bounded on $[0,1)$ and if $W(1)_{1}=+\infty$, then $U(1)_{1}$ exists and
-$U(1)_{1}=1\,.$
-Proof.
-In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$.
-Using the assumptions and Propositions 2.2, 2.4 and 2.6 we get that $F_{U}(x)^{2}\cdot F_{W}(x)=F_{W}(x)-F_{V}(x)$ for every $x\in[0,\,1)$.
-Since $F_{W}(0)>0$ and $F_{W}(x)$ is non-decreasing on $[0,1)$, we can divide by $F_{W}(x)$ and get that $F_{U}(x)^{2}=1-\frac{F_{V}(x)}{F_{W}(x)}$ for every $x\in[0,\,1)$.
-Note that $\lim_{x\to 1}F_{V}(x)=a>0$ because $F_{V}(0)>0$ and $F_{V}(x)$ is non-decreasing and bounded from above.
-Theorem 2.17 and properties of limits of functions give
-$(U(1)_{1})^{2}$ $=$ ${\textstyle\big(\lim_{x\to 1}F_{U}(x)\big)^{2}=\lim_{x\to 1}F_{U}(x)^{2}=\lim_{x\to 1}\big(1-\frac{F_{V}(x)}{F_{W}(x)}\big)}$
-$=$ ${\textstyle 1-\frac{\lim_{x\to 1}F_{V}(x)}{\lim_{x\to 1}F_{W}(x)}=1-\frac{a}{\lim_{n\to\infty}\sum_{j=0}^{n}w_{j}}=1-\frac{a}{+\infty}=1}\,.$
-Hence $U(1)_{1}=1$ because $U(1)_{1}\geq 0$. $\Box$
 We obtain analogs of the previous results in absolute convergence.
 Proposition 2.23
-Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{2}$ and $W(1)_{2}$ exist, then $U(1)_{2}$ exists and
+Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{2}$ and $W(1)_{2}$ exist, then
 $U(1)_{2}=\alpha V(1)_{2}+\beta W(1)_{2}\,.$
 Proof.
-Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}u_{n}x^{n}$.
-Since $V(1)_{2}$ and $W(1)_{2}$ are sums of absolutely convergent series $\sum_{n\in\mathbb{N}}v_{n}$ and $\sum_{n\in\mathbb{N}}w_{n}$ and $u_{n}=\alpha v_{n}+\beta w_{n}$, the result follows from Proposition 2.2.
+Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$.
+Since $V(1)_{2}$ and $W(1)_{2}$ are sums of absolutely convergent series $\sum_{n\in\mathbb{N}_{0}}v_{n}$ and $\sum_{n\in\mathbb{N}_{0}}w_{n}$, respectively, and $u_{n}=\alpha v_{n}+\beta w_{n}$, the result follows from Proposition 2.2.
 $\Box$
@@
   2. 2.
-    $\sum_{n\geq 0}|u_{n}|\leq\sum_{n\geq 0}|v_{n}|\cdot\sum_{n\geq 0}|w_{n}|$.
+    The sum $\sum_{n\geq 0}|u_{n}|\leq\sum_{n\geq 0}|v_{n}|\cdot\sum_{n\geq 0}|w_{n}|$.
 Proof. 1. We consider the series
 $S=\sum_{\langle j,\,k\rangle\in\mathbb{N}_{0}^{2}}v_{j}w_{k}=\sum_{j\in\mathbb{N}_{0}}v_{j}\cdot\sum_{k\in\mathbb{N}_{0}}w_{k}\,.$}
-By Proposition 2.5 it absolutely converges and has sum $V(1)_{2}W(1)_{2}$.
-On the other hand, $S$ has the grouping $\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}$ (the inner $\sum$ means sum and the outer $\sum$ means series) and $\sum_{j=0}^{n}v_{j}w_{n-j}=u_{n}$.
+By Proposition 2.5 it absolutely converges and has sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping
+$\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}\,$,}
+where the inner $\sum$ means sum, the outer $\sum$ means series, and $\sum_{j=0}^{n}v_{j}w_{n-j}=u_{n}$.
 By Proposition 2.4, $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}$.
@@
 $\Box$
+For nonnegative real coefficients we can go in the reverse, from the product to the factors.
+Proposition 2.25
+Let $U(x)$, $V(x)$ and $W(x)$ be in $\mathbb{R}_{\geq 0}[[x]]$.
+Suppose that $U(x)=V(x)W(x)$, that $U(x)$, $V(x)$ and $W(x)$ are in $H_{1}$, $V(x)W(x)\neq 0$ and that $U(1)_{2}$ exists.
+Then $V(1)_{2}$ and $W(1)_{2}$ exist, and
+$U(1)_{2}=V(1)_{2}W(1)_{2}\,.$
+Proof. By Theorem 2.18 and the assumption $V(x)W(x)\neq 0$ we have $V(1)_{2}=V(1)_{1}=\lim_{x\nearrow 1}F_{V}(x)>0$ and the same for $W(x)$. Since $F_{U}=F_{V}F_{W}$ on $\mathbb{C}_{<1}$, we have
+$U(1)_{2}$ $=$ $U(1)_{1}=\lim_{x\nearrow 1}F_{U}(x)$
+$=$ $\lim_{x\nearrow 1}F_{V}(x)\cdot\lim_{x\nearrow 1}F_{W}(x)=V(1)_{1}W(1)_{1}=V(1)_{2}W(1)_{2}\,.$
+In the second equality we used Theorem 2.18. The third equality holds because the product of limits is neither $0\cdot(+\infty)$ nor $(+\infty)\cdot 0$. $\Box$
 Proposition 2.20 for $n=2$ and item 1 of Proposition 2.24 are incomparable.
@@
 The latter assumes stronger absolute convergence, but only for two series.
-Another variant of the proposition is due to Franz (Franciszek) Mertens (1840–1927).
+Another variant of the proposition is due to F.
+Mertens.
 Theorem 2.26 (Mertens’s)
-Suppose that $U(x)$, $V(x)$ and $W(x)$ are in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$, and if $V(1)_{1}$ and $W(1)_{2}$ exist, then $U(1)_{1}$ exists and
+Suppose that $U(x)$, $V(x)$ and $W(x)$ are in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$, and if $V(1)_{1}$ and $W(1)_{2}$ exist, then
 $U(1)_{1}=V(1)_{1}W(1)_{2}\,.$
@@
   2. 2.
-    $\sum_{n\geq 0}|u_{n}|\leq\frac{1}{1-c}$.
-Proof. 1. It is clear that $V(1)_{2}$ exists. Since $|V(1)_{2}-1|\leq c<1$, we get $V(1)_{2}\neq 0$. We consider the series
+    The sum $\sum_{n\geq 0}|u_{n}|\leq\frac{1}{1-c}$.
+Proof. 1. From the assumption on $c$ it follows that $V(1)_{2}$ exists. From
+$|V(1)_{2}-1|\leq c<1$
+we get $V(1)_{2}\neq 0$. We consider the series
 $S=1+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}(-1)^{m}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\,.$}
-We show that $S$ absolutely converges.
-It follows from the bound that the sum of absolute values of summands in any finite subseries of $S$ is at most $1+\sum_{m\geq 1}(|v_{1}|+|v_{2}|+\dots)^{m}=\frac{1}{1-c}$.
-For $m\in\mathbb{N}$ we consider the series
+We show that $S$ absolutely converges. This follows from the bound that the sum of absolute values of summands in any finite subseries of $S$ is at most
+$1+\sum_{m\geq 1}(|v_{1}|+|v_{2}|+\dots)^{m}=\frac{1}{1-c}\,.$}
+Let $m\in\mathbb{N}$. We consider the series
 $S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}v_{j_{2}}\dots v_{j_{m}}\,.$}
-By Proposition 2.6 it absolutely converges and has sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has sum $1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}$.
+By Proposition 2.6, it absolutely converges and has sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has sum
+$1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}\,.$}
 On the other hand, in $\mathbb{C}[[x]]$ we have the identity
@@
 $u_{n}=\sum_{m\geq 1}(-1)^{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=n\end{subarray}}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\,.$}
-By Proposition 2.4, $S$ has also sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$ and the stated formula follows.
-2. Indeed,
+Proposition 2.4 shows that $S$ also has sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$. The stated formula follows.
+2. Indeed, the sum
 ${\textstyle\sum_{k=0}^{n}|u_{k}|=}$
@@
 Lemma 2.28
-Let $V(x)\in\mathbb{C}[[x]]$ and $V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$ where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$ (every square in $\mathbb{C}[[x]]$ has this form).
-Then there exists $v\in\mathbb{C}$ such that $v^{2}=t$ and
+Let $V(x)\in\mathbb{C}[[x]]$ and
+$V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$
+where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$ — every square in $\mathbb{C}[[x]]$ has this form. Then there exists a number $v\in\mathbb{C}$ such that $v^{2}=t$ and
 $V(x)=vx^{l}\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}\,.$}
-Proof. Let $n\in\mathbb{N}_{0}$ and a,ba,\,b be formal variables. By the Vandermonde identity,
+Proof. Let $n\in\mathbb{N}_{0}$ and a,ba,\,b be formal variables. By Vandermonde’s identity ([30]),
 $\sum_{j=0}^{n}\binom{a}{j}\binom{b}{n-j}=\binom{a+b}{n}\ \ (\in\mathbb{Q}[a,\,b])\,.$}
@@
 Lemma 2.29
-Let $U(x),V(x)\in\mathbb{C}[[x]]$ and $U(x)=V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$ where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$.
-If $c=|t_{1}|+|t_{2}|+\dots<1$, then $U(1)_{2}$ and $V(1)_{2}$ exist, and
+Let $U(x),V(x)\in\mathbb{C}[[x]]$ and
+$U(x)=V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$
+where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$. If $c=|t_{1}|+|t_{2}|+\dots<1$, then $U(1)_{2}$ and $V(1)_{2}$ exist, and
 $U(1)_{2}=\big(V(1)_{2}\big)^{2}\,.$
-Proof. Clearly, $U(1)_{2}$ exists. By Lemma 2.28 there is $v\in\mathbb{C}$ such that $v^{2}=t$ and
+Proof. It follows from the assumption on $c$ that $U(1)_{2}$ exists. By Lemma 2.28 there is a number $v\in\mathbb{C}$ such that $v^{2}=t$ and
 $V(x)=\sum_{n\geq 0}v_{n}x^{n}=vx^{l}\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}\,.$}
@@
 $S=0+0+\dots+0+v+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}v\binom{1/2}{m}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}$}
-with $l$ zero summands.
+beginning with $l$ summands 0.
 We show that it absolutely converges.
-Let $T$ be any finite subseries of $S$.
-Indeed, since $\big|\binom{1/2}{m}\big|\leq 1$ for every $m\in\mathbb{N}$, the sum of absolute values of summands in $T$ is at most
+Let $T$ be a finite subseries of $S$.
+Since $\big|\binom{1/2}{m}\big|\leq 1$ for every $m\in\mathbb{N}$, the sum of absolute values of summands in $T$ is at most
 $|v|+|v|\sum_{m\geq 1}(|t_{1}|+|t_{2}|+\dots)^{m}=\frac{|v|}{1-c}\,.$}
-By Proposition 2.4, $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.25, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
+Thus $S$ absolutely converges. By Proposition 2.4, the series $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.24, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
 Proposition 2.30
@@
 $\Box$
-Finally, we consider convex light weights.
+In conclusion, we consider $\xi$-convex light weights.
 Proposition 2.31
-For every convex light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ we have
-$\sum_{n=0}^{\infty}h(W(n))\cdot x^{n}=\frac{1}{1-x}$}\,.
+Let $\xi\in\mathbb{C}$. For every $\xi$-convex light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ we have
+$\sum_{n=0}^{\infty}h(W(n))x^{n}=\frac{1}{1-\xi x}\,.$}
+In particular, for $\xi=0$ we have $\sum_{n=0}^{\infty}h(W(n))x^{n}=1$.
 Proof.
-Let $h$ be a convex light weight.
-We prove by induction on $n\in\mathbb{N}_{0}$ that $h(W(n))=1$.
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be light and $\xi$-convex.
+We prove by induction on $n\in\mathbb{N}_{0}$ that $h(W(n))=\xi^{n}$.
 For $n=0$ this holds as $W(0)=\{\langle 1\rangle\}$.
@@
 Recall that $\ell(w)$ denotes the last vertex of a walk $w$.
-Using Propositions 2.2 and 2.4, the convexity assumption and induction, we get
+Using Propositions 2.2 and 2.4, the $\xi$-convexity assumption and induction, we get
 $h(W(n))$ $=$ ${\textstyle\sum_{w\in W(n)}h(w)}$
 $=$ ${\textstyle\sum_{\begin{subarray}{c}w^{\prime}\in W(n-1)\\ \ell(w^{\prime})\in V(h)\end{subarray}}h(w^{\prime})\sum_{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}}h(\{\ell(w^{\prime}),\,v\})}$
-$=$ ${\textstyle\sum_{w^{\prime}\in W(n-1)}h(w^{\prime})\cdot 1=1\,.}$
+$=$ ${\textstyle\sum_{w^{\prime}\in W(n-1)}h(w^{\prime})\cdot\xi=\xi^{n-1}\cdot\xi=\xi^{n}\,.}$
 We can add and remove the condition $\ell(w^{\prime})\in V(h)$ without affecting the sum because walks $w^{\prime}$ not satisfying it have zero weight. $\Box$
 3 Extensions of Pólya’s theorem with $v=1$
-We establish the first group of generalizations of Pólya’s theorem in Theorems 3.23.7. We consider the case $v=1$ when the vertex visited by walks coincides with their starting vertex.
-We introduce some generating functions.
-Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight.
-Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the set of walks that visit $v$ in step $i>0$.
+We begin with the case $v=1$ when the vertex to be visited by walks coincides with their starting vertex. In Theorems 3.23.7 we establish the first group of generalizations of Pólya’s theorem.
+For that end, we need some generating functions.
+Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$, and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight.
+Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the subset of walks visiting $v$ in some step $i>0$.
 We are interested in the generating function
 $A_{h}(x)=\sum_{n\geq 0}a_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(1,\,n))x^{n}$}
-and define
+and define three more generating functions
 $B_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h}x^{n}=\sum_{n\geq 0}h(W_{b}(n))x^{n}\,,}$
@@
 Here $W_{b}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and $W_{c}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and avoid $1$ between.
-The coefficients $a_{n}^{h}$, $\dots$, $d_{n}^{h}$ are correctly defined because they are sums of subseries of absolutely convergents series.
-We have $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$.
+The coefficients $a_{n}^{h}$, $\dots$, $d_{n}^{h}$ are correctly defined because they are sums of subseries of the absolutely convergent series
+$d_{n}^{h}=\sum_{w\in W(n)}h(w)\,.$}
+The constant coefficients are $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$.
 Proposition 3.1
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be as above. The following relations hold between them.
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be the above generating functions. The following relations hold.
   1. 1.
@@
 We show that $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}\cdot d_{n-j}^{h}$ for every $n\in\mathbb{N}_{0}$.
-For $n\leq 1$ it holds, $a_{n}^{h}=c_{n}^{h}=0$.
+For $n\leq 1$ it holds, then $a_{n}^{h}=c_{n}^{h}=0$.
 For $n\geq 2$ the set $W(1,n)$ is countable and we split every walk in it at the first revisit of $1$.
@@
 $F(w)=\langle w_{1},\,w_{2}\rangle\ \ (\in W_{c}(j)\times W(n-j))\,$,
-where $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$.
-It is easy to see that the map $F$ is a bijection.
-$F$ is weight-preserving (WP): for every weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and value $F(w)=\langle w_{1},w_{2}\rangle$ we have $h(w)=h(w_{1})h(w_{2})$.
+with $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$.
+It is easy to see that $F$ is a bijection.
+$F$ is also weight-preserving (WP): for every weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and value $F(w)=\langle w_{1},w_{2}\rangle$, we have $h(w)=h(w_{1})h(w_{2})$.
 If $X$ is a set of pairs of walks in $K_{\mathbb{N}}$, we write $h(X)$ for the sum of the series
 $\sum_{\langle w,\,w^{\prime}\rangle\in X}h(w)h(w^{\prime})\,$,}
-provided that it absolutely converges.
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and $n\geq 2$. Using $F$ and above propositions, we get
+if it absolutely converges.
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and let $n\geq 2$. Using the map $F$, we get
 $a_{n}^{h}$ $=$ $h(W(1,\,n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W(1,\,n)])$
@@
 $b_{n}^{h}=\sum_{j=1}^{\infty}\sum_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}c_{n_{1}}^{h}c_{n_{2}}^{h}\dots c_{n_{j}}^{h}\,.$}
-For odd $n$ it is true because then $b^{h}_{n}=c^{h}_{n}=0$ as $W_{b}(n)=W_{c}(n)=\emptyset$, and in every inner summand at least one index $n_{i}$ is odd.
+For odd $n$ it is true: for odd $n$ we have $b^{h}_{n}=c^{h}_{n}=0$, and in every inner summand at least one index $n_{i}$ is odd.
 For even $n\geq 2$ the set $W_{b}(n)$ is countable.
@@
 $F\colon W_{b}(n)\to\bigcup_{j=1}^{\infty}\bigcup_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}W_{c}(n_{1})\times\cdots\times W_{c}(n_{j})$}
-defined as follows. For any walk
+that is defined as follows. For any walk
 $w=\langle u_{0},\,\dots,\,u_{m_{1}},\,\dots,\,\dots,\,u_{m_{j-1}},\,\dots,\,u_{n}\rangle\in W_{b}(n)\,$,
 where $0=m_{0}<m_{1}<\dots<m_{j}=n$ and $u_{k}=1$ iff $k=0$ or $k=m_{i}$ for some $i\in[j]$, we set
-$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{1})\times W_{c}(n_{2})\times\dots\times W_{c}(n_{j}),\,n_{i}=m_{i}-m_{i-1})$
-where $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i\in[j]$.
-It is easy to see that $F$ is a bijection and that it is weight-preserving similarly as above.
+$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{1})\times W_{c}(n_{2})\times\dots\times W_{c}(n_{j}),\,n_{i}=m_{i}-m_{i-1})\,$,
+with $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i\in[j]$.
+It is easy to see that $F$ is a bijection and that it is weight-preserving, as in part 1.
 If $X$ is a set of tuples of walks, we write $h(X)$ for the sum of the series
-$\sum_{\langle w_{1},\,\dots,\,w_{k})\in X}h(w_{1})\dots h(w_{k})$}
+$\sum_{\langle w_{1},\,\dots,\,w_{k})\in X}h(w_{1})\dots h(w_{k})\,$,}
 if it absolutely converges.
-Let $h$ be a light weight and $n\geq 2$ be even. Using $F$ and above propositions, we get
+Let $h$ be a light weight and $n\geq 2$ be even. Using the map $F$, we get
 $b_{n}^{h}$ $=$ $h(W_{b}(n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W_{b}(n)])$
@@
 $\Box$
-For better orientation we label the fourteen theorems by the triples
-$\langle x,\,y,\,z\rangle\in\{v=1,\,v\neq 1\}\times\{\mathrm{gen},\,\mathrm{con}\}\times\{U(1)_{1},\,U(1)_{2},\,U(1)_{1}=+\infty\}\,$,
-where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to convex ones. The generating functions $A_{h}(x)$, $\dots$, $D_{h}(x)$ are as above.
+We label our fourteen generalizations of Pólya’s theorem for better orientation by the triples
+$\langle x,\,y,\,z\rangle\in\{v=1,\,v\neq 1\}\times\{\mathrm{gen},\,\mathrm{con}\}\times\{U(1)_{1},\,U(1)_{2},\,U(1)_{1}=\infty\}\,$,
+where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to $\xi$-convex ones.
+For $U(1)_{1}$ and $U(1)_{2}$ see Definition 2.12, and for $U(1)_{1}=\infty$ see Definition 2.14.
+The generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$, and $D_{h}(x)$ are defined at the beginning of this section.
 Theorem 3.2 (1: $v=1$, gen, $U(1)_{1}$)
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Suppose that $D_{h}(1)_{1}$, $C_{h}(1)_{1}$, $B_{h}(1)_{1}$ and $A_{h}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
+If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a light weight such that $A_{h}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
 $A_{h}(1)_{1}=\big(1-\frac{1}{B_{h}(1)_{1}}\big)D_{h}(1)_{1}\,.$}
@@
 By Proposition 2.20, $A_{h}(1)_{1}=C_{h}(1)_{1}D_{h}(1)_{1}$.
-By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$.
-By Proposition 2.20 we have $B_{h}(1)_{1}\neq 0$ and $C_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}$.
+By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$.
+By Proposition 2.21 we have $B_{h}(1)_{1}\neq 0$ and
+$(1-C_{h})(1)_{1}=\frac{1}{B_{h}(1)_{1}},\,\text{ so that }\,C_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}\,.$}
+The stated formula follows. $\Box$
+Theorem 3.3 (2: $v=1$, gen, $U(1)_{2}$)
+If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a light weight such that $\sum_{n\geq 1}|b_{n}^{h}|<1$ and $D_{h}(1)_{2}$ exists, then $B_{h}(1)_{2}\neq 0$, and
+$A_{h}(1)_{2}=\big(1-\frac{1}{B_{h}(1)_{2}}\big)D_{h}(1)_{2}\,.$}
+Proof. By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}\neq 0$ and
+$(1-C_{h}(1))_{2}=\frac{1}{B_{h}(1)_{2}},\,\text{ so that }\,C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$}
+By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{h}(1)_{2}=C_{h}(1)_{2}D_{h}(1)_{2}$ . The stated formula follows. $\Box$
+Theorem 3.4 (3: $v=1$, gen, $U(1)_{1}=\infty$)
+If $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a light weight such that $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
+$(A_{h})_{1/\xi}(1)_{1}=+\infty\,.$
+Proof.
+From $(D_{h})_{1/\xi}(1)_{1}=+\infty$ it follows that $\xi^{-n}d_{n}^{h}>0$ for some $n\in\mathbb{N}$.
+Hence there is a vertex $u\in\mathbb{N}\setminus\{1\}$ with $\xi^{-1}h(\{1,u\})>0$.
+The walk $\langle 1,u,1\rangle$ shows by the subseries bound principle that $\xi^{-2}c_{2}^{h}>0$.
+Since by item 1 of Proposition 3.1 we have $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}d_{n-j}^{h}$, we get that
+$\xi^{-2-n}a_{n+2}^{h}\geq\xi^{-2}c_{2}\cdot\xi^{-n}d_{n}^{h}$
+for every $n\in\mathbb{N}_{0}$. Hence
+$\sum_{j=0}^{n}\xi^{-2-j}a_{j+2}^{h}\geq\xi^{-2}c_{2}^{h}\cdot\sum_{j=0}^{n}\xi^{-j}d_{j}^{h}\to+\infty$}
+for $n\to\infty$, and $(A_{h})_{1/\xi}(1)_{1}=+\infty$. $\Box$
+We proceed to $\xi$-convex weights.
+Theorem 3.5 (4: $v=1$, con, $U(1)_{1}$)
+1.
+If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $\xi$-convex light weight and if $(B_{h})_{1/\xi}(1)_{1}$ and $(C_{h})_{1/\xi}(1)_{1}$ exist, then $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}\,.$
+2. If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a 0-convex light weight, and if $B_{h}(1)_{1}$ and $C_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
+$A_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}\,.$
+Proof. 1. By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}\,.$
+By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By Proposition 2.21 we have $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
+[…]
 The stated formula follows.
-$\Box$
-Theorem 3.3 (2: $v=1$, gen, $U(1)_{2}$)
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight.
-Suppose that $D_{h}(1)_{2}$ exists and $\sum_{n\geq 1}|b_{n}^{h}|<1$.
-Then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $A_{h}(1)_{2}$ exists and
-$A_{h}(1)_{2}=\big(1-\frac{1}{B_{h}(1)_{2}}\big)D_{h}(1)_{2}\,.$}
-Proof.
-By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$.
-By item 1 of Proposition 2.27, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h}(1)_{2}$ exists and $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$.
-By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$.
-By item 1 of Proposition 2.25, $A_{h}(1)_{2}$ exists and $A_{h}(1)_{2}=C_{h}(1)_{2}D_{h}(1)_{2}$ .
-The stated formula follows.
-$\Box$
-Theorem 3.4 (3: $v=1$, gen, $U(1)_{1}=+\infty$)
-Suppose that $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ is a light weight such that $D_{h}(1)_{1}=+\infty$. Then
-$A_{h}(1)_{1}=+\infty\,.$
-Proof.
-Thus $d_{n}^{h}>0$ for some $n>0$ and there is a vertex $u\in\mathbb{N}\setminus\{1\}$ with $h(\{1,u\})>0$.
-Then the walk $\langle 1,u,1\rangle$ shows that $c_{2}^{h}>0$.
-Since by item 1 of Proposition 3.1 we have $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}d_{n-j}^{h}$, we get $a_{n+2}^{h}\geq c_{2}^{h}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$.
-Hence
-$\sum_{j=0}^{n}a_{j+2}^{h}\geq c_{2}^{h}\cdot\sum_{j=0}^{n}d_{j}^{h}\to+\infty,\ n\to\infty\,$,}
-and $A_{h}(1)_{1}=+\infty$. $\Box$
-We proceed to convex weights.
-Theorem 3.5 (4: $v=1$, con, $U(1)_{1}$)
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a convex light weight. Suppose that $B_{h}(1)_{1}$ and $C_{h}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
-$\lim_{n\to\infty}{\textstyle a_{n}^{h}=1-\frac{1}{B_{h}(1)_{1}}\,.}$
-Proof.
-By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}$.
-By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)_{1}$.
-By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$.
-By Proposition 2.20 we have $B_{h}(1)_{1}\neq 0$ and $C_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}$.
-The stated formula follows.
-$\Box$
+2. Now, by item 1 of Proposition 3.1 and by Proposition 2.31, we have
+$A_{h}(x)=C_{h}(x)$
+and argue as in the previous case for $\xi=1$. $\Box$
 This theorem generalizes the case $d\geq 3$ of Pólya’s theorem with $\overline{v}=\overline{0}$. The case $d\leq 2$ is generalized in Theorem 3.7.
 Theorem 3.6 (5: $v=1$, con, $U(1)_{2}$)
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a convex light weight. Suppose that $\sum_{n\geq 1}|b_{n}^{h}|<1$. Then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$ and
-$\lim_{n\to\infty}{\textstyle a_{n}^{h}=1-\frac{1}{B_{h}(1)_{2}}\,.}$
-Proof.
-By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$.
-By item 1 of Proposition 2.27, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h}(1)_{2}$ exists and $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$.
-By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}$.
+1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $\xi$-convex light weight and $\sum_{n\geq 1}\big|b_{n}^{h}\big|<1$, then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$
+2. If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a 0-convex light weight and $\sum_{n\geq 1}\big|b_{n}^{h}\big|<1$, then $B_{h}(1)_{2}\neq 0$ and
+$A_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$
+Proof. 1. By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
+[…]
+By item 1 of Proposition 3.1 and Proposition 2.31, $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}$. By Proposition 2.15
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}=(C_{h})_{1/\xi}(1)_{2}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
+2. In the previous case we have for $\xi=1$ that $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. By item 1 of Proposition 3.1 and by Proposition 2.31, $A_{h}(x)=C_{h}(x)$. The stated formula follows. $\Box$
+Theorem 3.7 (6: $v=1$, con, $U(1)_{1}=\infty$)
+1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $\xi$-convex light weight and if $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1\,.$
+2. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ is a 0-convex light weight, $B_{h},C_{h}\in H_{1}$, and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
+$\lim_{x\nearrow 1}F_{(A_{h})_{1/\xi}}(x)=1\,.$
+Proof. 1. By item 1 of Proposition 3.1 and by Proposition 2.31 we have
+$A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}\,.$}
 By Proposition 2.15,
-$\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)_{1}=C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$}
-$\Box$
-Theorem 3.7 (6: $v=1$, con, $U(1)_{1}=+\infty$)
-Let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a convex light weight such that $B_{h}(1)=+\infty$. Then
-$\lim_{n\to\infty}a_{n}^{h}=1\,.$
-Proof.
-By item 1 of Proposition 3.1 and Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}$.
-By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)_{1}$.
-By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$.
-But now by Proposition 2.31 we have $0\leq b_{n}^{h},c_{n}^{h}\leq d_{n}^{h}=1$ for every $n\in\mathbb{N}_{0}$.
-Hence $b_{n}^{h},c_{n}^{h}\in[0,1]$ and the generating functions $B_{h}(x)$ and $C_{h}(x)$ absolutely converge for $x\in[0,1)$.
-By Proposition 2.21, $C_{h}(1)_{1}=1$.
-The stated formula follows.
-$\Box$
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}\,.$
+By Proposition 2.31 and the subseries bound principle,
+$0\leq\xi^{-n}b_{n}^{h},\,\xi^{-n}c_{n}^{h}\leq\xi^{-n}d_{n}^{h}=1$
+for every $n\in\mathbb{N}_{0}$. Hence $B_{h},C_{h}\in H_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. Thus
+$(C_{h})_{1/\xi}(1)_{1}=\lim_{x\nearrow 1}F_{(C_{h})_{1/\xi}}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1\,,}$
+where in the first and third equality we used Theorem 2.18. The stated formula follows.
+2. By item 1 of Proposition 3.1 and by Proposition 2.31, we have $A_{h}(x)=C_{h}(x)$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. Thus
+$\lim_{x\nearrow 1}F_{(C_{h})_{1/\xi}}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1-\frac{1}{\pm\infty}=1\,,}$
+where in the second equality we used Theorem 2.17. The stated formula follows at once. $\Box$
 4 Extensions of Pólya’s theorem with $v\neq 1$
-The case when the vertex visited by walks differs from their starting vertex is more complicated.
-The grid graph $\mathbb{Z}^{d}$ is vertex-transitive, but so far, this has been irrelevant.
-For $\overline{v}\neq\overline{0}$ vertex-transitivity becomes important, and we generalize it to $K_{\mathbb{N}}$ as follows.
-Let $v\in\mathbb{N}\setminus\{1\}$ with $v\neq 1$.
+In Theorems 4.34.10 we establish the second group of generalizations of Pólya’s theorem, in the more complicated case when the vertex to be visited by walks differs from their starting vertex.
+The grid graph $\mathbb{Z}^{d}$ is vertex-transitive, which has so far been irrelevant.
+For $\overline{v}\neq\overline{0}$, vertex-transitivity becomes relevant.
+We generalize it to $K_{\mathbb{N}}$ as follows.
+Let $v\in\mathbb{N}\setminus\{1\}$.
 We say that a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $v$-transitive if there is a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that $f(1)=v$ and $h(f[e])=h(e)$ for every $e\in\mathbb{N}_{2}$.
@@
 Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight.
-Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ are the walks visiting the vertex $v$ ($\neq 1$); necessarily in a step $i>0$.
+Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the subset of walks visiting the vertex $v$ ($\neq 1$); necessarily in a step $i>0$.
 We are interested in the generating function
 $A_{h,v}(x)=\sum_{n\geq 0}a_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W(v,\,n))x^{n}\,.$}
-The generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ are as before. We define
-$B_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{b,v}(n))x^{n}\,,}$
-$C_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{c,v}(n))x^{n}\,\text{ and}}$
-$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e,v}(n))x^{n}\,,}$
+In the previous section we introduced the generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$. Now we define
+$B_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{b}(v,\,n))x^{n}\,,}$
+$C_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{c}(v,\,n))x^{n}\,\text{ and}}$
+$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e}(v,\,n))x^{n}\,,}$
 all in $\mathbb{C}[[x]]$.
-Here $W_{b,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$.
-$W_{c,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding $v$.
-Finally, $W_{e,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding both $1$ and $v$.
-We have $a_{0}^{h,v}=c_{0}^{h,v}=c_{0}^{h}=e_{0}^{h,v}=0$ and $b_{0}^{h}=b_{0}^{h,v}=d_{0}^{h}=1$.
-We obtain an analog of Proposition 3.1. The proof is similar. We omit details and argue only on the high level of semiformal method.
+Here $W_{b}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$.
+$W_{c}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding $v$.
+Finally, $W_{e}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding both $1$ and $v$.
+The constant coefficients are $a_{0}^{h,v}=c_{0}^{h,v}=c_{0}^{h}=e_{0}^{h,v}=0$ and $b_{0}^{h}=b_{0}^{h,v}=d_{0}^{h}=1$.
+We obtain an analog of Proposition 3.1. Since the proof is similar, we omit details and argue only on the level of semiformal method.
 Proposition 4.1
-Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and generating functions $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ be as above.
-The following relations hold between them.
+Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ be the above generating functions.
+The following relations hold.
   1. 1.
@@
 1.
-Every walk $w$ weight-counted by $A_{h,v}(x)$ splits at the first visit of $v$ as $w=w_{1}w_{2}$.
+Walks $w$ weight-counted by $A_{h,v}(x)$ split at the first visit of $v$ as $w=w_{1}w_{2}$.
 Walks $w_{1}$ are weight-counted by $C_{h,v}(x)$ and $w_{2}$ are arbitrary walks starting at $v$.
@@
 Those avoiding $v$ are weight-counted by $B_{h,v}(x)$.
-If $w$ visits $v$, then it uniquely splits at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$.
+If $w$ visits $v$, it uniquely splits at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$.
 Walks $w_{1}$ start at $1$, end at $v$ and between avoid $v$.
@@
 $w\cdot\mathrm{sqrt}(z)=\{w\alpha\colon\;\alpha\in\mathbb{C},\,\alpha^{2}=z\}\ \ (\subset\mathbb{C})\,.$
-For $w=0$ or $z=0$ it equals $\{0\}$. Else it has two non-zero elements differing by sign. If $w=1$, we write just $\mathrm{sqrt}(z)$.
+For $w=0$ or $z=0$ it equals $\{0\}$. Else it has two non-zero elements differing only by sign. If $w=1$, we write just $\mathrm{sqrt}(z)$.
+Before we turn to generalizations of Pólya’s theorem in the case $v\neq 1$, we utilize the square in item 2 of Proposition 4.1.
+If $h\colon X\to\mathbb{R}$ is an absolutely convergent real series and $Y\subset X$, then in general the sum $\sum_{x\in Y}h(x)$ of the subseries need not be smaller than the sum $\sum_{x\in X}h(x)$.
+However, due to the mentioned square, we obtain the next monotonicity result for the sums
+$b_{n}^{h,v}=\sum_{w\in W_{b}(v,\,n)}h(w)\,\text{ and }\,b_{n}^{h}=\sum_{w\in W_{b}(n)}h(w)\,.$}
+It is expressed in terms of the generating functions $B_{h,v}(x)=\sum_{n\geq 0}b_{n}^{h,v}x^{n}$ and $B_{h}(x)=\sum_{n\geq 0}b_{n}^{h}x^{n}$.
+Theorem 4.2
+Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{R}$ be a $v$-transitive light weight and let $B_{h,v},B_{h},C_{h,v}\in H_{1}$.
+Then for every real $c\in(-1,1)$ the following holds.
+  1. 1.
+    If $B_{h}(c)=0$ then also $B_{h,v}(c)=0$.
+  2. 2.
+    If $B_{h}(c)\neq 0$ then
+    $\frac{B_{h,v}(c)}{B_{h}(c)}\leq 1\,.$
+Proof. This follows from relations 2 in Proposition 4.1. $\Box$
+We proceed to the remaining generalizations of Pólya’s theorem.
 Theorem 4.3 (7: $v\neq 1$, gen, $U(1)_{1}$)
-Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$.
-Suppose that $D_{h}(1)_{1}$, $C_{h,v}(1)_{1}$, $C_{h}(1)_{1}$, $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$ and $A_{h,v}(1)_{1}$ exist.
-Then $B_{h}(1)_{1}\neq 0$ and
+If $v$ is in $\mathbb{N}\setminus\{1\}$, if $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$, and if $A_{h,v}(1)_{1}$, $B_{h}(1)_{1}\neq 0$, $B_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then
 $A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,.$
@@
 By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
-By item 2 of Proposition 3.1 and by Proposition 2.20, $B_{h}(1)_{1}\neq 0$.
-Thus by Proposition 2.22, $C_{h,v}(1)_{1}^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$.
+Thus by Proposition 2.22, $\big(C_{h,v}(1)_{1}\big)^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$.
 The stated formula follows.
 $\Box$
-We assume the existence of $E_{h,v}(1)_{1}$ instead of $C_{h}(1)_{1}$ and get the following variant of Theorem 4.2.
+If we replace the assumption of non-vanishing of $B_{h}(1)_{1}$ with the existence of $E_{h,v}(1)_{1}$, we obtain the following variant of the theorem.
 Theorem 4.4 (7: $v\neq 1$, gen, $U(1)_{1}$)
-Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$.
-Suppose that $E_{h,v}(1)_{1}$, $D_{h}(1)_{1}$, $C_{h,v}(1)_{1}$, $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$ and $A_{h,v}(1)_{1}$ exist.
-Then
+If $v$ is in $\mathbb{N}\setminus\{1\}$, if $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$, and if $A_{h,v}(1)_{1}$, $B_{h}(1)_{1}$, $B_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$, $D_{h}(1)_{1}$ and $E_{h,v}(1)_{1}$ exist, then
 $A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,$,
-where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret $\frac{0}{0}$ as $1$, so that $A_{h,v}(1)_{1}=0$.
+where for $B_{h}(1)_{1}=0$ we have $B_{h,v}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then $A_{h,v}(1)_{1}=0$.
 Proof.
-For $B_{h}(1)_{1}\neq 0$ we use the previous proof.
-If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.19 and 2.18 that also $B_{h,v}(1)_{1}=0$.
+If $B_{h}(1)_{1}\neq 0$, we use the previous proof.
+If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1, and Propositions 2.20 and 2.19 that also $B_{h,v}(1)_{1}=0$.
 Item 3 of Proposition 4.1 and Proposition 2.20 give $C_{h,v}(1)_{1}=0$.
@@
 Theorem 4.5 (8: $v\neq 1$, gen, $U(1)_{2}$)
-Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$.
-Suppose that $D_{h}(1)_{2}$ and $B_{h,v}(1)_{2}$ exist, and that $c=\sum_{n\geq 1}|b_{n}^{h}|<1$.
+Suppose that $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$,
+$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
+and that $B_{h,v}(1)_{2}$ and $D_{h}(1)_{2}$ exist.
 If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$.
-Let $B_{h,v}(x)\neq B_{h}(x)$ and $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$.
+For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$.
 If
-$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1$}
-then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $A_{h,v}(1)_{2}$ exists and
+$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
+then $B_{h}(1)_{2}\neq 0$ and
 $A_{h,\,v}(1)_{2}\in{\textstyle D_{h}(1)_{2}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
@@
 By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
-By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $(C_{h,v}(1)_{2})^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$.
-By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$.
-By item 1 of Proposition 2.25, $A_{h,v}(1)_{2}$ exists and $A_{h,v}(1)_{2}=C_{h,v}(1)_{2}D_{h}(1)_{2}$.
+By Proposition 2.30, $B_{h}(1)_{2}\neq 0$ and
+$(C_{h,v}(1)_{2})^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}\,.$}
+By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{h,v}(1)_{2}=C_{h,v}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
+Theorem 4.6 (9: $v\neq 1$, gen, $U(1)_{1}=\infty$)
+If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $v$-transitive light weight with $v\in V(h)$ and if $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
+$(A_{h,v})_{1/\xi}(1)_{1}=+\infty\,.$
+Proof.
+We denote the minimum length of a walk joining $1$ and $v$ in $G(h)$ by $m$.
+The inner vertices of it differ from $v$.
+By the subseries bound principle we get $\xi^{-m}c^{h,v}_{m}>0$.
+By item 1 of Proposition 4.1 we have $a_{n}^{h,v}=\sum_{j=0}^{n}c_{j}^{h,v}d_{n-j}^{h}$ and we conclude that $\xi^{-n-m}a_{n+m}^{h,v}\geq\xi^{-m}c_{m}^{h,v}\cdot\xi^{-n}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$.
+Hence
+$\sum_{j=0}^{n}\xi^{-j-m}a_{j+m}^{h,v}\geq\xi^{-m}c_{m}^{h,v}\cdot\sum_{j=0}^{n}\xi^{-j}d_{j}^{h}\to+\infty$}
+for $n\to\infty$, and $(A_{h,v})_{1/\xi}(1)_{1}=+\infty$. $\Box$
+We proceed to $\xi$-convex weights.
+Theorem 4.7 (10: $v\neq 1$, con, $U(1)_{1}$)
+1.
+If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h,v})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and $(C_{h,v})_{1/\xi}(1)_{1}$ exist, then
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\big)\,.$}
+2.
+If $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, and if $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}\neq 0$ and $C_{h,v}(1)_{1}$ exist, then
+$A_{h,v}(1)_{1}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}\big)\,.$}
+Proof. 1. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}\,.$
+By item 2 of Proposition 4.1 we have $(C_{h,v})_{1/\xi}(x)^{2}=1-\frac{(B_{h,v})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}$.
+By Proposition 2.22, $\big((C_{h,v})_{1/\xi}(1)_{1}\big)^{2}=1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}$.
 The stated formula follows.
-$\Box$
-Theorem 4.5 (9: $v\neq 1$, gen, $U(1)_{1}=+\infty$)
-Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive light weight such that $v\in V(h)$. If $D_{h}(1)_{1}=+\infty$ then
-$A_{h,v}(1)_{1}=+\infty\,.$
+2. Now, by item 1 of Proposition 4.1 and by Proposition 2.31, we have
+$A_{h,v}(x)=C_{h,v}(x)\,.$
+The rest of the argument is the same as in the previous case for $\xi=1$. $\Box$
+Again, if we replace the assumption of non-vanishing of $B_{h,v}(1)_{1}$ with the existence of $E_{h,v}(1)_{1}$, we obtain the following variant of the theorem.
+Theorem 4.8 (10: $v\neq 1$, con, $U(1)_{1}$)
+1.
+If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h,v})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}$, $(C_{h,v})_{1/\xi}(1)_{1}$ and $(E_{h,v})_{1/\xi}(1)_{1}$ exist, then
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\big)\,$,}
+where for $(B_{h})_{1/\xi}(1)_{1}=0$ we have $(B_{h,v})_{1/\xi}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then the limit is zero.
+2.
+If $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, and if $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h,v}(1)_{1}$ and $E_{h,v}(1)_{1}$ exist, then
+$A_{h,v}(1)_{1}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}\big)\,$,}
+where for $B_{h}(1)_{1}=0$ we have $B_{h,v}(1)_{1}=0$ and interpret the $\frac{0}{0}$ as $1$, so that then $A_{h,v}(1)_{1}=0$.
 Proof.
-Let $m\in\mathbb{N}$ be the minimum length of a walk in $G(h)$ joining $1$ and $v$.
-The inner vertices of this shortest walk differ from $v$ and we get that $c^{h,v}_{m}>0$.
-Since by item 1 of Proposition 4.1 we have $a_{n}^{h,v}=\sum_{j=0}^{n}c_{j}^{h,v}d_{n-j}^{h}$, we conclude that $a_{n+m}^{h,v}\geq c_{m}^{h,v}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$.
-Hence
-$\sum_{j=0}^{n}a_{j+m}^{h,v}\geq c_{m}^{h,v}\cdot\sum_{j=0}^{n}d_{j}^{h}$}
-for every $n\in\mathbb{N}_{0}$ and $A_{h,v}(1)_{1}=+\infty$. $\Box$
-We proceed to convex weights.
-Theorem 4.6 (10: $v\neq 1$, con, $U(1)_{1}$)
-Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive convex light weight such that $v\in V(h)$.
-Suppose that $C_{h,v}(1)_{1}$, $C_{h}(1)_{1}$, $B_{h,v}(1)_{1}$ and $B_{h}(1)_{1}$ exist.
-Then $B_{h}(1)_{1}\neq 0$ and
-$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,.$
-Proof.
-By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$.
-By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h,v}=C_{h,v}(1)_{1}$.
-By item 2 of Proposition 3.1 and by Proposition 2.20, $B_{h}(1)_{1}\neq 0$.
-By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
-By Proposition 2.22, $C_{h,v}(1)_{1}^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$.
-The stated formula follows.
-$\Box$
-We again assume the existence of $E_{h,v}(1)_{1}$ instead of $C_{h}(1)_{1}$ and get the following variant of the theorem.
-Theorem 4.7 (10: $v\neq 1$, con, $U(1)_{1}$)
-Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive convex light weight such that $v\in V(h)$.
-Suppose that $E_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$, $B_{h,v}(1)_{1}$ and $B_{h}(1)_{1}$ exist.
-Then
-$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\ \ (\subset\mathbb{C})\,$,
-where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret $\frac{0}{0}$ as $1$, so that the limit is zero.
-Proof.
-If $B_{h}(1)_{1}\neq 0$, the previous proof works.
-If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.19 and 2.18 that also $B_{h,v}(1)_{1}=0$.
-Item 3 of Proposition 4.1 and Proposition 2.19 give $C_{h,v}(1)_{1}=0$.
-Thus, as in the previous proof, $\lim_{n\to\infty}a_{n}^{h,v}=C_{h,v}(1)_{1}=0$.
-$\Box$
+1.
+If $(B_{h})_{1/\xi}(1)_{1}\neq 0$, we use the previous proof.
+If $(B_{h})_{1/\xi}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.20 and 2.19 that also $(B_{h,v})_{1/\xi}(1)_{1}=0$.
+Item 3 of Proposition 4.1 and Proposition 2.20 give $(C_{h,v})_{\xi}(1)_{1}=0$.
+Thus, as in the previous theorem,
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}=0\,.$
+2. Now, by item 1 of Proposition 4.1 and by Proposition 2.31, we have
+$A_{h,v}(x)=C_{h,v}(x)\,.$
+The rest of the argument is the same as in the previous case for $\xi=1$. $\Box$
 These two theorems generalize the case $d\geq 3$ of Pólya’s theorem with $\overline{v}\neq\overline{0}$. The case $d\leq 2$ is generalized in Theorem 4.10.
 Theorem 4.9 (11: $v\neq 1$, con, $U(1)_{2}$)
-Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive convex light weight such that $v\in V(h)$.
-Suppose that $B_{h,v}(1)_{2}$ exists and $c=\sum_{n\geq 1}|b_{n}^{h}|<1$.
-If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$.
-Let $B_{h,v}(x)\neq B_{h}(x)$ and $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$.
-If
-$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1$}
-then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$ and
-$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
+1. Suppose that $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$,
+$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
+and that $(B_{h,v})_{1/\xi}(1)_{2}$ exists. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
+$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
+then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}}\big)\,.$
+2. Suppose that $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$,
+$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
+and that $B_{h,v}(1)_{2}$ exists. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
+$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
+then $(B_{h})(1)_{2}\neq 0$ and
+$A_{h,v}(1)_{2}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
 Proof.
-If $B_{h,v}(x)\neq B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=1-1=0$ and $C_{h,v}(x)=0$.
+If $B_{h,v}(x)=B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=0$ and $C_{h,v}(x)=0$.
 Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
-Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold.
+1.
+Let $h$ be $\xi$-converx with $\xi\in\mathbb{C}_{1}$, $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold.
 By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
-By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $C_{h,v}(1)_{2}^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$.
-By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$.
-Proposition 2.15 gives the stated formula:
-$\lim_{n\to\infty}a_{n}^{h,v}={\textstyle C_{h,\,v}(1)_{1}=C_{h,\,v}(1)_{2}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}\big)\,.}$
+By Proposition 2.30, $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
+$\big((C_{h,v})_{1/\xi}(1)_{2}\big)^{2}=1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
+By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$. Proposition 2.15 gives the stated formula:
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}={\textstyle(C_{h,\,v})_{1/\xi}(1)_{1}=(C_{h,\,v})_{1/\xi}(1)_{2}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\big)\,.}$
+2.
+Let $h$ be 0-convex, $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold.
+The beginning of the previous case gives for $\xi=1$ that $\big(C_{h,v}(1)_{2}\big)^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$.
+Since by item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)$, the stated formula follows.
 $\Box$
-Theorem 4.9 (12: $v\neq 1$, con, $U(1)_{1}=+\infty$)
-Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive convex light weight such that $v\in V(h)$. If $B_{h}(1)_{1}=+\infty$ then
-$\lim_{n\to\infty}a_{n}^{h,v}=1\,.$
-Proof.
-By Proposition 2.31, $D_{h}(x)=\frac{1}{1-x}$.
-Hence $0\leq b_{n}^{h},b_{n}^{h,v},c_{n}^{h,v},e_{n}^{h,v}\leq d_{n}^{h}=1$ for every $n\in\mathbb{N}_{0}$ and we can use continuous and non-decreasing functions
-$F_{B_{h}},\,F_{B_{h,v}},\,F_{C_{h,v}},\,F_{E_{h,v}}\colon[0,\,1)\to[0,\,+\infty)\,.$
-Any shortest walk in $G(h)$ joining $1$ and $v$ shows that for some $m\in\mathbb{N}$ we have $e_{m}^{h,v}>0$.
-Thus the function $F_{E_{h,v}}(x)$ is on the interval $(0,1)$ positive and increasing.
-We deduce from it that $F_{B_{h,v}}(x)$ is bounded on $[0,1)$.
-Suppose for the contrary that $\lim_{x\to 1}F_{B_{h,v}}(x)=+\infty$.
-By item 3 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality $F_{C_{h,v}}(x)=F_{B_{h,v}}(x)F_{E_{h,v}}(x)$.
-Thus $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$.
-Since $F_{C_{h,v}}(0)=0$, $F_{C_{h,v}}(x)$ increases and is continuous and $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$, we obtain a (unique) number $x_{0}\in(0,1)$ such that $F_{C_{h,v}}(x)<1$ for $x\in[0,x_{0})$ and $\lim_{x\to x_{0}^{-}}F_{C_{h,v}}(x)=1$.
-By item 2 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality
-$F_{B_{h}}(x)=F_{B_{h,v}}(x)+F_{C_{h,v}}(x)^{2}F_{B_{h}}(x)\,\,.$
-Hence for every $x\in[0,x_{0})$ we have $F_{B_{h}}(x)=\frac{F_{B_{h,v}}(x)}{1-F_{C_{h,v}}(x)^{2}}$.
-Since $F_{B_{h,v}}(0)=1$, it follows that $\lim_{x\to x_{0}^{-}}F_{B_{h}}(x)=+\infty$.
-But this contradicts the continuity of $F_{B_{h}}(x)$ at $x_{0}$.
-Using item 2 of Proposition 4.1, the boundedness of $F_{B_{h,v}}(x)$ and Proposition 2.23, we get $C_{h,v}(1)_{1}=1$.
-By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$.
-By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h,v}=C_{h,v}(1)_{1}=1$.
+Theorem 4.10 (12: $v\neq 1$, con, $U(1)_{1}=\infty$)
+1.
+If $\xi\in\mathbb{C}_{1}$, $v$ is in $\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=1\,.$
+2.
+If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, $B_{h},C_{h,v}\in H_{1}$, $(B_{h,v})_{1/\xi}(1)_{1}$ exists and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
+$\lim_{x\nearrow 1}F_{(A_{h,v})_{1/\xi}}(x)\in\{-1,\,1\}\,.$
+Proof. 1. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-\xi x}$. By the subseries bound principle we have
+$0\leq\xi^{-n}b_{n}^{h},\,\xi^{-n}b_{n}^{h,v},\,\xi^{-n}c_{n}^{h,v},\,\xi^{-n}e_{n}^{h,v}\leq\xi^{-n}d_{n}^{h}=1$
+for every $n\in\mathbb{N}_{0}$.
+Hence $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$ and $E_{h,v}(x)$ are in $H_{1}$.
+Any shortest walk joining $1$ and $v$ in $G(h)$ shows that for some $m\in\mathbb{N}$ we have $\xi^{-m}e_{m}^{h,v}>0$.
+Thus the function $F_{(E_{h,v})_{1/\xi}}(x)$ is positive and increasing on $(0,1)$.
+Since $F_{(C_{h,v})_{1/\xi}}(x)$ is bounded on $[0,1)$ by item 2 of Proposition 4.1, $F_{(B_{h,v})_{1/\xi}}(x)$ is bounded by item 3.
+Since by item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$, Proposition 2.15 gives
+$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}=(C_{h,v})_{1/\xi}(1)_{2}\,.$
+By item 2 of Proposition 4.1, Propositions 2.2, 2.4 and 2.5, and by Theorem 2.18 we have
+$(C_{h,v}^{2})_{1/\xi}(1)_{1}=\lim_{x\nearrow 1}F_{(C_{h,v}^{2})_{1/\xi}}(x)=1-\frac{\lim_{x\nearrow 1}F_{(B_{h,v})_{1/\xi}}(x)}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1\,$,}
+because $F_{(B_{h,v})_{1/\xi}}(x)$ is bounded on $[0,1)$ and $\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)=B_{h}(1)_{1}=+\infty$. By Proposition 2.25, $(C_{h,v})_{1/\xi}(1)_{2}$ exists and
+$\big((C_{h,v})_{1/\xi}(1)_{2}\big)^{2}=\big((C_{h,v})_{1/\xi}\big)^{2}(1)_{2}=(C_{h,v}^{2})_{1/\xi}(1)_{2}=(C_{h,v}^{2})_{1/\xi}(1)_{1}=1\,.$
+Since $(C_{h,v})_{1/\xi}(1)_{2}\geq 0$, we have $(C_{h,v})_{1/\xi}(1)_{2}=1$ and the stated formula follows.
+2. By item 1 of Proposition 4.1 and by Proposition 2.31, $A_{h,v}(x)=C_{h,v}(x)$. By item 2 of Proposition 4.1 and by Theorems 2.16 and 2.17,
+$\lim_{x\nearrow 1}F_{(C_{h,v}^{2})_{1/\xi}}(x)=1-\frac{\lim_{x\nearrow 1}F_{(B_{h,v})_{1/\xi}}(x)}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{\pm\infty}=1\,.$}
+By the continuity of the functions $F_{\dots}$,
+$\lim_{x\nearrow 1}F_{(A_{h,v})_{1/\xi}}(x)=\lim_{x\nearrow 1}F_{(C_{h,v})_{1/\xi}}(x)\in\{-1,\,1\}\,.$
 $\Box$
 5 Concluding remarks
-In [17] we consider the same problem of determination of formulas for the quantities
-$A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))$}
-for formal weights $h$. These are maps
+In [17] we consider the same problem of determining formulas for quantities such as $A_{h,v}(1)=\sum_{n\geq 0}h(W(v,n))$ in the case of formal weights $h$. These are maps
 $h\colon\mathbb{N}_{2}\to\mathbb{C}[[x_{1},\,x_{2},\,\dots,\,x_{k}]]\,$,
 where the last integral domain of formal power series in $k$ variables is endowed with the usual non-Archimedean norm.
-The main novelty of our generalization is this. In an equality like
+The main novelty of our generalization is this. In equalities like
 $P_{d}(\overline{v},\,n)=h(W(v,\,n))$
-in the Introduction, the left-hand side involves, in the standard rigorous probabilistic treatment, an implicit probability space supporting the corresponding Markov chain.
+that was stated in the Introduction, the left-hand side involves, in the standard rigorous probabilistic treatment, an implicit probability space supporting the corresponding Markov chain.
 Construction of this probability space is far from trivial; see, for example, [6].
-In our approach, the right-hand side is completely self-contained.
+In our generalization, the right-hand side is completely self-contained.

### v8

@@
-Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights
+Extending the symbolic method in enumerative combinatorics. I
 Abstract
-Using the semiformal method in combinatorics we generalize Pólya’s theorem.
-This theorem determines limits of probabilities that walks in the grid graph $\mathbb{Z}^{d}$ starting at the origin visit the given vertex.
-We generalize it to the countable complete graph $K_{\mathbb{N}}$ with edge weights in $\mathbb{C}$.
-In part II we consider analogs of Pólya’s theorem for edge weights in $\mathbb{C}[[x_{1},\dots,x_{k}]]$.
-Besides the abstract, we want to help the reader understand our article by the following preamble. In it we have more space to explain what we do.
-We mention Pólya’s theorem, which is a well known result in the discipline of random walks.
-However, our article is not really about generalizing Pólya’s theorem.
-We use extensions of Pólya’s theorem to edge weighted countable complete graphs merely as a test case for the novel generalization of the well known ([3, 8, 10, 12, 22, 23, 28]) symbolic method in enumerative (analytic) combinatorics.
-This generalization is due to the first author, and we call it the semiformal method.
-It will be developed in detail in [14].
-The present article is an explanation of the functioning of the semiformal method in the form of an extensive example.
-More applications of it are given in [15].
-The idea of the semiformal method is to extend finite sums defining coefficients of generating functions in the symbolic method to infinite (countable) series.
-Finite sums in a ring (in the symbolic method, one also needs to multiply) are straightforward, but it is less clear what is a sum of an infinite series of real or complex numbers.
-We use three kinds of sums of infinite series: absolute convergence (which is used in probability where summands are nonnegative), the more general and weaker conditional convergence, and sums equal to $\pm\infty$.
-The last sums are in fact important — the heart of the combinatorial proof of the case $d\leq 2$ in Pólya’s theorem is the fact that a certain series has sum $\infty$.
-For details on the uses of these three kinds of sums we refer to the article below.
-But why confine ourselves to infinite series of real or complex numbers, when non-Archimedean infinite series are also interesting?
-This is the topic of [17].
-To summarize, the main contribution of this article is the demonstration of the working of the semiformal method in the fourteen cases that arise when Pólya’s theorem is generalized to $K_{\mathbb{N}}$ with complex edge weights.
+We use our extension of the symbolic method in enumerative combinatorics (we extend finite sums determining coefficients in generating functions to infinite series) to generalize Pólya’s theorem.
+This theorem finds the limits of probabilities that walks in the grid graph $\mathbb{Z}^{d}$, starting at the origin, visit the given vertex.
+We replace the grid graph with weighted countable complete graphs $K_{\mathbb{N}}$ such that the edge weights have values in $\mathbb{C}$ and satisfy a convergence condition.
+Thus general Pólya’s theorems are results in enumerative combinatorics.
 1 Introduction
+We recall in general terms the fundamental symbolic method in enumerative combinatorics, briefly the symbolic method, and explain how we extend it.
+Later in the article, we exemplify our extension on Pólya’s theorem.
+Our selection of references on the symbolic method is [3, 8, 11, 13, 29].
+1.1 The symbolic method and its extension
+Using this method, we enumerate—by exact counting formulas and by approximate asymptotic formulas—an infinite (countable) set $A$ of combinatorial objects.
+$A$ is equipped with a size function $s\colon A\to\mathbb{N}_{0}$ ($\{0,1,\dots\}$).
+It is assumed that for every $n\in\mathbb{N}_{0}$ the set of size $n$ objects in $A$,
+$A_{n}=\{a\in A\colon\;s(a)=n\}\,$,
+is finite; see [3, p. 1], [8, p. ix], [11, Definition I.1 on p. 16] and [13, Remark 2.2.3 on p. 32, implicitly]. We associate with $A$ and $s$ the generating function
+$A(x)=\sum_{a\in A}x^{s(a)}=\sum_{n\geq 0}|A_{n}|x^{n}\ \ (\in\mathbb{C}[[x]])\,$,}
+where $|X|$ denotes the cardinality of a finite set $X$.
+It is common, and crucial in the approach here, to have a weight function $h\colon A\to R$, where $R$ is a ring.
+The weighted version of $A(x)$ is then
+$A_{h}(x)=\sum_{n\geq 0}(\sum_{a\in A_{n}}h(a))x^{n}\ \ (\in R[[x]])\,.$}
+Note that every coefficient $\sum_{a\in A_{n}}h(a)$ in $A_{h}(x)$ is a finite sum.
+The symbolic method works in two phases.
+The first phase translates combinatorial relations between objects in $A$ into algebraic, differential, or functional equations for $A_{h}(x)$ and the auxiliary generating functions.
+Multivariate generating functions can also be used.
+The second phase deduces from these equations asymptotic formulas for the coefficients of generating functions.
+The second analytic phase accounts for the usefulness and popularity of the method (see [11]), but the first formal phase is the primary one (the memoir [3] is an impressive automatization of this phase).
+We give two examples to illustrate the symbolic method: one short now and another, detailed in the next subsection.
+Let $A$ be the set of all proper bracketings $b$, let $s(b)$ be the number of pairs of brackets in $b$ and let $h(b)=1$.
+For example,
+$A_{3}=\{()()(),\,()(()),\,(())(),\,(()()),\,((()))\}\,.$
+The numbers $(|A_{n}|)_{n\geq 0}=(1,1,2,5,14,42,\dots)$ are the Catalan numbers [28]. The two phases of the symbolic method yield the following.
+  1. 1.
+    We consider the first pair of brackets in $b$ and get the algebraic equation $A(x)=1+xA(x)^{2}$ (we omit the index h because the weight is trivial).
+It leads to the following exact counting formulas.
+    1. (a)
+      $|A_{0}|=1$ and $|A_{n+1}|=\sum_{j=0}^{n}|A_{j}|\cdot|A_{n-j}|$ for $n\in\mathbb{N}_{0}$.
+    2. (b)
+      $|A_{n}|=\frac{1}{n+1}\binom{2n}{n}$ for $n\in\mathbb{N}_{0}$.
+    3. (c)
+      $|A_{0}|=1$ and $|A_{n+1}|=\frac{4n+2}{n+2}\cdot|A_{n}|$ for $n\in\mathbb{N}_{0}$.
+    For $m\in\mathbb{N}_{0}$, let $v(m)=\log(m+2)$; up to multiplicative constant, $v(m)$ is the number of binary (or $q$-ary, for any $q\geq 2$) digits in $m$.
+Each of (a)–(c) is an efficient closed formula for the counting function $C:\;n\mapsto|A_{n}|$ because it provides an algorithm that computes $C$ in
+    $O(\max(v(n),\,v(|A_{n}|))^{c})=O((n+1)^{c}),\ n\in\mathbb{N}_{0},\,$
+    ($c>0$ is a constant) bit operations. See the next phase for the asymptotics of $|A_{n}|$.
+  2. 2.
+    By [11, p. 38], the equation $xA(x)^{2}-A(x)+1=0$ (or (b)) produces the approximate asymptotic formula
+    $|A_{n}|=(1+o(1))\pi^{-1/2}n^{-3/2}4^{n}\ \ (n\to\infty)\,.$}
+    See [11, Chapter VII.7] for the theory of asymptotics of coefficients of algebraic formal power series.
+At the end of phase 1 we remind the definition of an efficient closed formula for a counting function, which was given by the first author in [15, p.
+10, Wilfian formulas].
+This example of application of the symbolic method is standard, maybe even archetypal.
+Our next example in the next subsection is less standard
+Our idea of extension of the symbolic method is simple and natural.
+We drop the assumption that the sets $A_{n}$ are finite, and replace it with the assumption that the ring $R$ is endowed with a complete norm, for example, $R=\mathbb{R}$ or $R=\mathbb{C}$, and that all possibly infinite series
+$\sum_{a\in A_{n}}h(a)$}
+determining coefficients in $A_{h}(x)$ absolutely converge.
+This extension relies on our semiformal approach to formal power series.
+In this approach, we allow operations with formal power series in which coefficients arise from limit transitions.
+In this article, $R=\mathbb{C}$ and we work out an extensive example in which we apply the extended symbolic method to generalize Pólya’s theorem, as explained below.
+In [17] we develop the semiformal approach in general.
+In [19] we obtain analogs of the results in this article for the ring $R=\mathbb{C}[[x_{1},\dots,x_{k}]]$ which is endowed with a complete non-Archimedean norm.
+In [16] we use the semiformal approach in number theory and present two proofs of the transcendence of the number $\mathrm{e}$ based on formal power series.
+1.2 Pólya’s theorem for $\overline{v}=\overline{0}$ via the symbolic method
 Let $d,n\in\mathbb{N}$ ($\{1,2,\dots\}$) and $\overline{v}\in\mathbb{Z}^{d}$, where $\mathbb{Z}=\{\dots,-1,0,1,\dots\}$ are the integers.
-Let $P_{d}(\overline{v},n)$ be the probability that a walk of length $n$ in the grid graph $\mathbb{Z}^{d}$ starting at $\overline{0}$ visits in a step $i>0$ the vertex $\overline{v}$.
+Let $P_{d}(\overline{v},n)$ be the probability that a walk of length $n$ in the grid graph $\mathbb{Z}^{d}$, starting at the origin $\overline{0}=\langle 0,0,\dots,0\rangle$ ($\in\mathbb{Z}^{d}$), later visits the vertex $\overline{v}$.
 In 1921, G.
-Pólya proved in [24] a theorem, which we generalize, that for every vertex $\overline{v}$,
-$\lim_{n\to\infty}P_{d}(\overline{v},\,n)\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ <1&\dots&\text{if $d\geq 3$}\,.\end{array}\right$.
-Pólya’s theorem is almost always mentioned in the special case with $\overline{v}=\overline{0}$, but he proved it for any vertex $\overline{v}$ (see [16, Section 1] for a quote where he states his result).
-A selection of articles and books discussing Pólya’s theorem is [4, 5, 6, 9, 11, 13, 18, 19, 20, 21, 25, 26, 27, 31, 32].
-We generalize it as follows. Let
-$\mathbb{N}_{2}=\binom{\mathbb{N}}{2}=\{e\colon\;e\subset\mathbb{N}\text{ with }|e|=2\}$}
-be edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$. Here $|A|$ ($\in\mathbb{N}_{0}=\{0,1.\dots\}$) denotes the cardinality of a finite set $A$. Let
-$h\colon\mathbb{N}_{2}\to\mathbb{C}$
-be an edge weight, where $\mathbb{C}$ denotes the field of complex numbers, which is endowed with the usual Archimedean absolute value $|z|=\sqrt{z\cdot\overline{z}}$.
-Let $n\in\mathbb{N}_{0}$.
-We denote by $W(n)$ the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$.
-Thus $W(0)=\{\langle 1\rangle\}$ and for $n>0$ the set $W(n)$ is infinite and countable.
+Pólya proved in [24] a theorem, which we generalize in our article, that for every dimension $d\in\mathbb{N}$ and every vertex $\overline{v}$,
+$\lim_{n\to\infty}P_{d}(\overline{v},\,n)\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ \in(0,\,1)&\dots&\text{if $d\geq 3$}\,.\end{array}\right$. (1)
+Pólya’s theorem is almost always mentioned in the special case when $\overline{v}=\overline{0}$, but Pólya proved it for every vertex $\overline{v}$ (see [18, Section 1] for a quote in which he states his result).
+A selection of articles and books discussing Pólya’s theorem is [4, 5, 6, 9, 10, 12, 14, 20, 21, 22, 23, 25, 26, 27, 32, 33].
+Now we derive (1) for $\overline{v}=\overline{0}$ by the symbolic method. The grid graph $G_{d}$ is the $2d$-regular graph $G_{d}=\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
+$\{\overline{a},\,\overline{b}\}\in E_{d}\iff\sum_{i=1}^{d}|\overline{a}_{i}-\overline{b}_{i}|=1\,.$}
+We take the constant map $h\colon E_{d}\to\{\frac{1}{2d}\}$ and extend it to walks in $G_{d}$ by $h(w)=(2d)^{-n}$ where $n$ ($\in\mathbb{N}_{0}$) is the length of the walk $w$.
+Let $A$ be the set of (finite) walks in $G_{d}$ that start at the origin $\overline{0}$ and later revisit it.
+The size function $s$ is defined on every walk $w$ in $G_{d}$ and $s(w)$ is the length of $w$.
+Then for every $d\in\mathbb{N}$ and $n\in\mathbb{N}_{0}$,
+$P_{d}(\overline{0},\,n)=\sum_{w\in A_{n}}h(w)=:h(A_{n})$}
+(recall that $A_{n}=\{w\in A\colon\;s(w)=n\}$). We are interested in the main generating function
+$A_{h}(x)=\sum_{n\geq 0}h(A_{n})x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
+We obtain a formula for $\lim_{n\to\infty}h(A_{n})$ with the help of three auxiliary generating functions:
+$B_{h}(x)=\sum_{n\geq 0}h(B_{n})x^{n},\ C_{h}(x)=\sum_{n\geq 0}h(C_{n})x^{n},\ D_{h}(x)=\sum_{n\geq 0}h(D_{n})x^{n}\,.$}
+Here $B$ is the set of closed walks in $G_{d}$ that begin and end at $\overline{0}$, $C\subset B$ is the subset of closed walks with inner vertices different from $\overline{0}$ and $D$ is the set of all walks in $G_{d}$ starting at $\overline{0}$.
+We have $h(A_{0})=h(C_{0})=0$ and $h(B_{0})=h(D_{0})=1$.
+The first formal phase of the symbolic method yields two formal identities,
+$A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}\,\text{ and }\,C_{h}(x)=1-\frac{1}{B_{h}(x)}\ \ (\in\mathbb{C}[[x]])\,.$}
+The second equality reflects the fact that $h(D_{n})=1$ for every $n\in\mathbb{N}_{0}$.
+The other two equalities are also easy to prove; in Proposition 3.1 we prove them in a much more general setting.
+The first identity implies that for every $n\in\mathbb{N}_{0}$,
+$h(A_{n})=\sum_{j=0}^{n}h(C_{j})\,.$}
+We proceed to the second analytic phase. We consider the sums
+$C_{h}(1)=\sum_{n\geq 0}h(C_{n}):=\lim_{n\to\infty}\sum_{j=0}^{n}h(C_{j})\ \ (\in\mathbb{C})$}
+and $B_{h}(1)$ which is defined similarly.
+Since the formal power series $B_{h}(x)$ and $C_{h}(x)$ have coefficients in $[0,1]$, the formal identity $C_{h}(x)=1-\frac{1}{B_{h}(x)}$ in $\mathbb{C}[[x]]$ converts to the identity
+$C_{h}(1)=1-\frac{1}{B_{h}(1)}\ \ (\in[0,\,1])\,$,}
+where $B_{h}(1)\in[1,\,+\infty)\cup\{+\infty\}$.
+We rigorously prove it.
+We formally have $B_{h}(x)C_{h}(x)=B_{h}(x)-1$.
+Since $0\leq h(B_{n}),h(C_{n})\leq h(D_{n})=1$ for every $n\in\mathbb{N}_{0}$, the formal power series $B_{h}(x)$ and $C_{h}(x)$ absolutely converge for every $x\in[0,1)$.
+Using the operations of linear combination, grouping, and product for complex series in Propositions 2.2, 2.4, and 2.5, respectively, we get that
+$\text{$B_{h}(x)C_{h}(x)=B_{h}(x)-1$ for every number $x\in[0,\,1)$}\,.$
+Since $B_{h}(x)\geq 1$ for every $x\in[0,1)$, we can divide by it and get that
+$\text{$C_{h}(x)=1-\frac{1}{B_{h}(x)}$ for every $x\in[0,\,1)$}\,.$}
+Using the arithmetic of functional limits and Theorem 2.18 (Abel’s theorem for nonnegative real coefficients), we get that indeed
+$C_{h}(1)=\lim_{x\nearrow 1}C_{h}(x)=\lim_{x\nearrow 1}\big(1-\frac{1}{B_{h}(x)}\big)=1-\frac{1}{\lim_{x\nearrow 1}B_{h}(x)}=1-\frac{1}{B_{h}(1)}\,.$}
+Here $\lim_{x\nearrow 1}$ denotes that $x$ approaches $1$ through the values $x\in[0,1)$.
+Thus
+$\lim_{n\to\infty}P_{d}(\overline{0},\,n)=\lim_{n\to\infty}h(A_{n})=\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}h(C_{j})=C_{h}(1)=1-\frac{1}{B_{h}(1)}\,.}$
+Since $B_{h}(1)=+\infty$ for $d\leq 2$ and $1<B_{h}(1)<+\infty$ for $d\geq 3$ (this requires some work), Pólya’s theorem for $\overline{v}=\overline{0}$ follows. $\Box$
+See Theorem 3.5 for a generalization of this result. We have not seen the formula
+$\lim_{n\to\infty}P_{d}(\overline{0},\,n)=\lim_{n\to\infty}h(A_{n})=1-\frac{1}{B_{h}(1)}\,$,} (2)
+which unites the cases $d\leq 2$ and $d\geq 3$ of Pólya’s theorem for $\overline{v}=\overline{0}$, in the literature.
+The use of generating functions in the proof of Pólya’s theorem is well known, but the theorem itself is mostly viewed as a result in probability or random walks; see, for example, [6] for a probabilistic proof.
+One of our goals is to convince the reader that Pólya’s theorem can also be naturally viewed as a result in enumerative combinatorics.
+The latter interpretation does not require tools of probability theory, such as Markov chains.
+1.3 Pólya’s theorem for $\overline{v}\neq\overline{0}$ via the symbolic method
+Let $\overline{v}\neq\overline{0}$ and $A_{\overline{v}}$ be the set of walks in $G_{d}$ starting at $\overline{0}$ and visiting $\overline{v}$.
+In the case $d\leq 2$ when $B_{h}(1)=+\infty$ we have, as a special case of item 1 in Theorem 4.13, that
+$\lim_{n\to\infty}P_{d}(\overline{v},\,n)=\lim_{n\to\infty}h(A_{\overline{v},\,n})=1\,.$ (3)
+In the case $d\geq 3$ when $1<B_{h}(1)<+\infty$, we can obtain, by a more involved version of the argument in Section 1.2, the formula
+$\lim_{n\to\infty}P_{d}(\overline{v},\,n)=\lim_{n\to\infty}h(A_{\overline{v},\,n})={\textstyle\sqrt{1-\frac{B_{\overline{v},\,h}(1)}{B_{h}(1)}}}\,.$ (4)
+Here $B_{\overline{v}}$ ($\subset B$) is the set of walks in $G_{d}$ that begin and end at $\overline{0}$ and avoid $\overline{v}$. It is easy to see that
+$1\leq B_{\overline{v},\,h}(1)<B_{h}(1)\,$,
+which implies the second case in (1). This result is generalized in Theorem 4.9. We have not seen the formula (4) in the literature.
+1.4 General Pólya’s theorems
+The idea of generalizing Pólya’s theorem is simple. If we extend the grid graph $G_{d}$ to the countable complete graph
+$K=\big\langle\mathbb{Z}^{d},\,\binom{\mathbb{Z}^{d}}{2}\big\rangle\,$,}
+extend the weight $h\colon E_{d}\to\{\frac{1}{2d}\}$ to the weight $h\colon\binom{\mathbb{Z}^{d}}{2}\to\{0,\frac{1}{2d}\}$ by setting $h=0$ on $\binom{\mathbb{Z}^{d}}{2}\setminus E_{d}$, if we then extend $h$ multiplicatively to walks in $K$ (see below), and if $A_{\overline{v}}$ is the set of walks in $K$ starting at $\overline{0}$ and later visiting the given vertex $\overline{v}$, then nothing changes in the above discussion of Pólya’s theorem in Sections 1.2 and 1.3.
+So we generalize it by allowing any complex weight $h\colon\binom{\mathbb{Z}^{d}}{2}\to\mathbb{C}$, assuming that a convergence condition holds so that the coefficients $h(A_{\overline{v},n})$ and coefficients in auxiliary generating functions, which are now sums of (possibly) infinite series, exist.
+For simplicity we replace $\mathbb{Z}^{d}$ with the set $\mathbb{N}$ of natural numbers and work with the countable complete graph
+$K_{\mathbb{N}}=\langle\mathbb{N},\,\mathbb{N}_{2}\rangle,\ \mathbb{N}_{2}=\binom{\mathbb{N}}{2}\,$.}
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a weight.
+$\mathbb{C}$ denotes the field of complex numbers, with the usual complete Archimedean norm (absolute value) $|z|=\sqrt{z\cdot\overline{z}}$.
+A walk $w$ in $K_{\mathbb{N}}$ is an $n+1$-tuple of vertices
+$w=\langle u_{0},\,u_{1},\,\dots,\,u_{n}\rangle,\,\text{$n\in\mathbb{N}_{0}$ and $u_{i}\in\mathbb{N}$}\,$,
+such that $u_{i-1}\neq u_{i}$ for every $i=1,2,\dots,n$.
+Let $W$ be the set of all walks and $s\colon W\to\mathbb{N}_{0}$ be the size function given by $s(w)=n$, that is, $s(w)$ is the length of $w$.
+We standardly extend the weight $h$ to $W$: $h(w)=1$ if $s(w)=0$, and
+$h(w)=\prod_{i=1}^{n}h(\{u_{i-1},\,u_{i}\})$}
+if $s(w)=n\geq 1$. By the reversal of walks we mean the bijection (even involution) $\varphi\colon W\to W$,
+$\varphi(\langle u_{0},\,u_{1},\,\dots,\,u_{n}\rangle)=\langle u_{n},\,u_{n-1},\,\dots,\,u_{0}\rangle\,.$
+The map $\varphi$ is length-preserving and weight-preserving: always $s(w)=s(\varphi(w))$ and $h(w)=h(\varphi(w))$.
+For $n\in\mathbb{N}_{0}$ we denote by $D_{n}$ the set of walks of length (size) $n$ in $K_{\mathbb{N}}$ starting at $1$.
+Thus $D_{0}=\{\langle 1\rangle\}$ and for $n\geq 1$ the set $D_{n}$ is infinite and countable.
 Let $v\in\mathbb{N}$.
-We denote by $W(v,n)$ ($\subset W(n)$) the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$ and visiting $v$ in a step $i>0$.
-We extend the weight multiplicatively to
-$h\colon\bigcup_{n=0}^{\infty}W(n)\to\mathbb{C}\,.$}
-Let $W\subset\bigcup_{n=0}^{\infty}W(n)$. If the series
-$\sum_{w\in W}h(w)$}
-absolutely converges, we denote its sum by $h(W)$.
-(Results on series are reviewed in the next section.) Suppose that for every $n\in\mathbb{N}_{0}$ the series $\sum_{w\in W(n)}h(w)$ absolutely converges.
-Then so does the series $\sum_{w\in W(v,n)}h(w)$ and we consider the generating function
-$A_{h,v}(x)=\sum_{n\geq 0}h(W(v,\,n))x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
-Our generalizations of Pólya’s theorem are formulas for three quantities:
-$\lim_{x\nearrow 1}A_{h,v}(x),\ A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))\,$,}
-where $\lim_{x\nearrow 1}$ means that $x$ goes to $1$ through values $x\in[0,1)$.
-In some cases, these quantities are twisted by powers $\xi^{n}$ for $\xi\in\mathbb{C}^{*}$ ($\mathbb{C}\setminus\{0\}$).
-The formulas appear in fourteen theorems: Theorems 3.23.7 in Section 3 and Theorems 4.34.10 in Section 4.
-Pólya’s theorem is subsumed in five of them.
-Theorems 3.5 and 3.7 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}=\overline{0}$.
-Theorems 4.7 (4.8) and 4.10 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
-The fourteen generalizations come from three sources.
-First, weights in Pólya’s theorem are nonnegative real numbers, and the convergence of infinite series is absolute.
-For our complex weights, we consider two kinds of convergence: absolute and conditional.
-We also consider real series with sums $\pm\infty$.
-Second, Pólya’s theorem is on probabilities, which means that the weights of edges incident to a fixed vertex sum up to $\xi=1$.
-We generalize it to $\xi$-convexity for any $\xi\in\mathbb{C}$; the case of $\xi\neq 0$ is normalized to $|\xi|=1$ and the case $\xi=0$ is treated separately (it corresponds to Kirchhoff’s first law).
-Thus we consider two cases, general weights and $\xi$-convex weights.
-Third, Pólya’s theorem is, in reality, two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in the complexity of proofs.
-In our generalization, we correspondingly distinguish between cases $v=1$ and $v\neq 1$.
-Now $3\cdot 2\cdot 2=12$ and two additional generalizations arise by varying the assumptions.
-We explain how our approach includes Pólya’s original problem. If $f\colon A\to B$ is a map and $C$ is any set (not necessarily a subset of $A$), we denote by
-$f[C]=\{f(x)\colon\;x\in A\cap C\}\ \ (\subset B)$
-the image of $C$ by $f$.
-Let $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$.
-We take any bijection $f\colon\mathbb{Z}^{d}\to\mathbb{N}$ such that $f(\overline{0})=1$, and set $v=f(\overline{v})$.
-The grid graph $\mathbb{Z}^{d}$, which is the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
-$\{\overline{a},\,\overline{b}\}\in E_{d}\iff\sum_{i=1}^{d}|a_{i}-b_{i}|=1\,$,}
-determines a weight $h\colon\mathbb{N}_{2}\to\{\frac{1}{2d},0\}$ by the relation $h(e)\neq 0$ iff $e=f[e^{\prime}]$ for $e^{\prime}\in E_{d}$. Then for every $n\in\mathbb{N}_{0}$ we have
-$P_{d}(\overline{v},\,n)=h(W(v,\,n))\,.$
-See Section 5 for the importance of (the general form of) this equality.
-We structure our article as follows.
-In Section 2 we review infinite series with complex summands and the domain $\mathbb{C}[[x]]$ of generating functions.
-Sections 3 and 4 contain our main results, the mentioned fourteen generalizations of Pólya’s theorem.
-Section 5 contains concluding remarks.
-We continue our investigation in [17]; see Section 5 for more details on [17].
-In Sections 3 and 4 we use what we term the semiformal method in (enumerative) combinatorics.
-It extends the symbolic method in combinatorics [3, 8, 10, 12, 22, 23, 28] from finite to countable sets.
-In our situation such extension is natural and inevitable, coefficients in generating functions $A_{h,v}(x)$ arise as sums of infinite series $\sum_{w\in W(v,n)}h(w)$.
-We develop this method more in [14].
-Why “semiformal”?
-Generating “functions” in the symbolic method are formal power series and their coefficients are determined by finite operations.
-We also use formal power series, but our extension allows operations based on limit transitions.
-See [15] for more applications of the semiformal method.
+We denote by $A_{v}$ the set of walks in $K_{\mathbb{N}}$ starting at $1$ and later visiting $v$.
+Then
+$A_{v,n}=(A_{v})_{n}=\{\langle u_{0},\,\dots,\,u_{n}\rangle\in W\colon\;\text{$u_{0}=1$ and $\exists\,i>0$ such that $u_{i}=v$}\}\,.$
+The said convergence condition is Definition 2.1: every series
+$\sum_{w\in D_{n}}h(w),\ n\in\mathbb{N}_{0}\,$,}
+absolutely converges.
+Absolute convergence is defined at the beginning of the next section.
+Then every series $\sum_{w\in A_{v,n}}h(w)$, $n\in\mathbb{N}_{0}$, absolutely converges as well; we denote its sum by $h(A_{v,n})$.
+We consider the main generating function
+$A_{v,\,h}(x)=(A_{v})_{h}(x)=\sum_{n\geq 0}h(A_{v,\,n})x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
+The classical theorem of Pólya is the formula (1).
+Many other graph families besides $G(d)$ have been investigated.
+What should the general Pólya’s theorems be about?
+It is not so clear, weights in them are general, $h\colon\mathbb{N}_{2}\to\mathbb{C}$, and the original probabilistic interpretation disappears in the distance.
+We investigate a variety of weights: general weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$, $\xi$-convex weights with $\xi\in\mathbb{C}$ and $|\xi|=1$, 0-convex weights (see Definition 2.10), weights $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ and weights $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$.
+Our answer is that general Pólya’s theorems are identities of the forms (2)–(4) for the walks $A_{v}$ in terms of quantities of four kinds, which we present below in the order of decreasing precision.
+  1. 1.
+    The limit $\lim_{n\to\infty}h(A_{v,n})$ of total weights of walks in $A_{v,n}$.
+  2. 2.
+    The sum $A_{v,h}(1)_{2}$ of absolutely convergent series
+    $h(A_{v,0})+h(A_{v,1})+\dots$
+    of those total weights.
+  3. 3.
+    The sum $A_{v,h}(1)_{1}$ of conditionally convergent series
+    $h(A_{v,0})+h(A_{v,1})+\dots\$
+    of those total weights.
+  4. 4.
+    The generalized sum
+    $\lim_{x\nearrow 1}A_{v,h}(x)=\lim_{x\nearrow 1}{\textstyle\sum_{n\geq 0}h(A_{v,\,n})x^{n}}\,.$
+In accordance with the ranges of weights, these quantities are usually twisted by powers of $\xi$ for $\xi\in\mathbb{C}^{*}$ ($\mathbb{C}\setminus\{0\}$).
+In Sections 3 and 4 we obtain along these lines a total of sixteen general Pólya’s theorems.
+Some questions remain.
+In Theorem 3.2 we obtain the formula $A_{h}(1)_{1}=(1-\frac{1}{B(1)_{1}})D_{h}(1)_{1}$ which we deduce from the (simpler!) formula $A_{h}(1)_{1}=C_{h}(1)_{1}D_{h}(1)_{1}$.
+Why do we prefer the former formula over the latter?
+First, the walks $B$ (starting and ending at $1$) have simpler definition than the walks $C$ (starting and ending at $1$ and avoiding $1$ between).
+Second, (1) follows from formulas of the former type, but it is not clear how to deduce it from the latter.
+See also the discussion after Proposition 4.3.
+1.5 Plan of the article
+In this section, we have hopefully explained the main ideas of our article and its context.
+To work with generating functions whose coefficients are sums of infinite series, we need a theory of complex series defined on arbitrary countable sets, such as the sets $D_{n}$ of walks in $K_{\mathbb{N}}$ of length $n$ starting at $1$.
+We could not find such a theory in the literature, at least not in the form we need, and therefore we develop it in Section 2.
+Here is one example.
+The product rule (formula) in the symbolic method, the identity that
+$|A\times B|=|A|\cdot|B|$
+for any two finite sets $A$ and $B$, is so simple and obvious that one encounters difficulties when asked to provide a detailed proof: first, axioms of set theory and natural numbers have to be specified, and so on.
+The actual product formula operates on the level of generating functions, ordinary or exponential ones, and is less immediate, but it has at its heart the displayed set-theoretic identity; see [3, pp.
+22, 40], [8, Theorem on pp.
+3–4], [11, Definition I.6 on p.
+23] and [13, 2.2.14 Lemma on pp.
+36–37].
+In our extension of the symbolic method, we lift the product rule to the nontrivial Proposition 2.5 on products of absolutely convergent series.
+Proposition 2.4 concerns another important operation on series, the grouping.
+Other useful results in Section 2 are three Abel’s Theorems 2.162.18 (stated without proofs) and various conversions of formal identities to identities at $x=1$ (we saw one example in Section 1.2), such as Proposition 2.30.
+In Section 3 we present generalizations of Pólya’s theorem in the simple case $v=1$ when the vertex to be visited by walks in $K_{\mathbb{N}}$ coincides with their starting vertex.
+In Proposition 3.1, we give detailed proofs of formal identities
+$A_{h}(x)=C_{h}(x)D_{h}(x)\,\text{ and }\,C_{h}(x)=1-\frac{1}{B_{h}(x)}\ \ (\in\mathbb{C}[[x]])$}
+whose instances for the graph $G_{d}$ we saw in Section 1.2.
+At first glance, they look like typical identities produced in the first phase of the symbolic method, but they are in fact completely new.
+The coefficients in the generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$, and $D_{h}(x)$ arise as sums of infinite (absolutely convergent) series.
+In the symbolic method, such coefficients always arise as finite sums or finite linear combinations.
+The proof of Proposition 3.1 therefore transcends analogous proofs in the standard symbolic method.
+Our seven general Pólya’s theorems for $v=1$ are given in Theorems 3.23.8.
+As we already mentioned, these theorems show that formulas (2)–(4) hold for weights $h$ far more general than probabilistic ones and they explore the logical space of conditional and absolute convergence assumptions on the values at $x=1$ of generating functions.
+See Section 5 for the summary of these theorems.
+Results in Section 4 in Theorems 4.54.13 are analogous to Section 3 but they are obtained in the more complicated case $v\neq 1$ of general Pólya’s theorems; see Section 5 for the summary of these theorems..
+Proofs of the formal identities
+$A_{v,\,h}(x)=C_{v,\,h}(x)D_{h}(x),\ B_{h}(x)=B_{v,\,h}(x)+C_{v,\,h}(x)^{2}B_{h}(x)\,\text{ and}$
+$C_{v,\,h}(x)=B_{v,\,h}(x)E_{v,\,h}(x)$
+in Proposition 4.2 are only sketched.
+For $v\neq 1$ one can introduce more auxiliary generating functions than for $v=1$, and in Proposition 4.3 we give another two formal identities for the generating function $F_{v,h}(x)$ of walks going from $1$ to $v$.
+In Theorem 4.4 we show that for any weight $h\colon\mathbb{N}_{2}\to\mathbb{R}$ and any vertex $v\neq 1$ we have, for every real $x\in(-1,1)$ and under appropriate assumptions, that if $B_{h}(x)=0$ then $B_{v,h}(x)=0$, and if $B_{h}(x)\neq 0$ then
+$\frac{B_{v,\,h}(x)}{B_{h}(x)}\leq 1\,.$
+This generalizes the trivial bound
+$0\leq B_{v,\,h}(x)\leq B_{h}(x)$
+for $x\in[0,1)$ and $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$.
+1.6 Perspectives and further research
+We already mentioned that in [17] we develop a semiformal theory of formal power series.
+Many isolated instances of such uses of formal power series can already be found in the literature, and in [17] we want to systematize them.
+For example, the classical identity
+$1-\frac{(\pi x)^{2}}{3!}+\frac{(\pi x)^{4}}{5!}-\dots=\prod_{n\geq 1}\Big(1-\frac{x^{2}}{n^{2}}\Big)\ \ (\in\mathbb{C}[[x]])$
+resulting from the representation of the function $\sin(\pi x)/(\pi x)$ by an infinite product can be understood as an identity for formal power series employing semiformal operations.
+In [19] we generalize Pólya’s theorem to $K_{\mathbb{N}}$ with edge weights in the ring $\mathbb{C}[[x_{1},\dots,x_{k}]]$.
+There the norm is non-Archimedean, and we are curious what results we obtain.
+Weights with values in polynomial or formal power series rings are common in enumerative combinatorics.
+It would be interesting to obtain general Pólya’s theorems for several (maybe infinitely many?) vertices $v\in\mathbb{N}$ to be visited by walks in $K_{\mathbb{N}}$ starting at $1$.
 2 Series and generating functions in $\mathbb{C}$
-We review series and generating functions in $\mathbb{C}$.
-A series is a map $h\colon X\to\mathbb{C}$ defined on a finite or countable set $X$; we write it as $\sum_{x\in X}h(x)$.
+In this section, we develop a theory of general complex series and obtain propositions on the conversion of formal identities for generating functions to identities at $x=1$.
+We use these results in the two following sections.
+2.1 General complex series
+series is a map $h\colon X\to\mathbb{C}$ defined on a finite or countable set $X$; we write it also as $\sum_{x\in X}h(x)$.
 We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
@@
   2. 2.
-    If the set $X$ is countable, then for every bijection $f\colon\mathbb{N}\to X$ the proper limit
+    If $X$ is countable, then for every bijection $f\colon\mathbb{N}\to X$, the limit
     $s=\lim_{n\to\infty}{\textstyle\sum_{i=1}^{n}h(f(i))\ \ (\in\mathbb{C})}$
     exists and does not depend on $f$.
-(In fact, the independence of $s$ on $f$ may be omitted from the formulation.) We call $s$ the sum of the series and denote it again by $\sum_{x\in X}h(x)$.
+In fact, the independence of $s$ on $f$ may be omitted from the condition.
+We call the number $s$ the sum of the series and denote it again by $\sum_{x\in X}h(x)$ or by $h(X)$.
 If the set $X$ is finite and nonempty, $X=\{x_{1},x_{2},\dots,x_{n}\}$, then every series $\sum_{x\in X}h(x)$ absolutely converges and has the sum
 $h(x_{1})+h(x_{2})+\dots+h(x_{n})\,.$
-If $X=\emptyset$, we define the sum as 0. If
+We set $h(\emptyset)=0$. If
 $U=U(x)=\sum_{n\geq 0}u_{n}x^{n}\ \ (\in\mathbb{C}[[x]])$}
-is a formal power series, so that $x$ is a formal variable, we denote for any $c\in\mathbb{C}$ by $F_{U}(c)$ the sum of the power series $\sum_{n\geq 0}u_{n}c^{n}$, if it absolutely converges. Let
+is a formal power series, so that $x$ is a formal variable, we denote for any $c\in\mathbb{C}$ by $U(c)$ the sum of the power series $\sum_{n\geq 0}u_{n}c^{n}$, if it absolutely converges. Let
 $H_{1}={\textstyle\{U(x)\in\mathbb{C}[[x]]\colon\;\text{$U(c)$ absolutely converges for every $c\in\mathbb{C}_{<1}$}\}\,,}$
-where $\mathbb{C}_{<1}=\{c\in\mathbb{C}\colon\;|c|<1\}$. Thus for any $U\in H_{1}$ the sums form the function $F_{U}\colon\mathbb{C}_{<1}\to\mathbb{C}$.
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. A walk $w$ in $K_{\mathbb{N}}$ is any $(n+1)$-tuple $w$ of vertices $v_{i}\in\mathbb{N}$,
-$w=\langle v_{0},\,v_{1},\,\dots,\,v_{n}\rangle\,$,
-such that $n\in\mathbb{N}_{0}$ and $v_{i-1}\neq v_{i}$ for every $i\in[n]$ ($\{1,2,\dots,n\}$, $0]=\emptyset$). The length of $w$ is $n$ ($\in\mathbb{N}_{0}$). We extend $h$ to walks. For $n>0$ we set
-$h(w)=\prod_{i=1}^{n}h(\{v_{i-1},\,v_{i}\})\,$,}
-and for $n=0$ we define $h(w)=1$. Recall that $W(n)$ is the set of walks $w$ in $K_{\mathbb{N}}$ with length $n$ such that $v_{0}=1$.
-Definition 2.1 (lightness)
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We call the weight $h$ light if for every $n\in\mathbb{N}_{0}$, the series
-$\sum_{w\in W(n)}h(w)$}
+where $\mathbb{C}_{<1}=\{c\in\mathbb{C}\colon\;|c|<1\}$. So for any formal power series $U(x)\in H_{1}$ the sums $U(c)$ form the function
+$U\colon\mathbb{C}_{<1}\to\mathbb{C},\ c\mapsto U(c)\,.$
+Next definition is basic in our approach. Recall that $D_{n}$ is the set of walks in $K_{\mathbb{N}}$ with length $n$ starting at $1$.
+Definition 2.1 (light weights)
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a weight. We call $h$ light if for every $n\in\mathbb{N}_{0}$, the series
+$\sum_{w\in D_{n}}h(w)$}
 absolutely converges.
@@
 If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$.
-Any subseries of an absolutely convergent series absolutely converges.
-However, we cannot compare the sizes of sums
-$\big|\sum_{x\in Y}h(x)\big|\,\text{ and }\,\big|\sum_{x\in X}h(x)\big|\,.$}
-Except when $h\colon X\to\mathbb{R}_{\geq 0}$: if $\sum_{x\in X}h(x)$ absolutely converges and $Y\subset X$, then the sums satisfy
-$0\leq\sum_{x\in Y}h(x)\leq\sum_{x\in X}h(x)$}
-— we call it the subseries bound principle.
-If $R=\sum_{x\in X}g(x)$ and $S=\sum_{x\in X}h(x)$ are series and $\alpha,\beta\in\mathbb{C}$, the series
+Any subseries of an absolutely convergent series absolutely converges, but we cannot compare the sizes of sums $h(Y)$ and $h(X)$; we may have $h(X)=0$ and $h(Y)\neq 0$.
+However, if $h\colon X\to\mathbb{R}_{\geq 0}$, $\sum_{x\in X}h(x)$ absolutely converges, and if $Y\subset X$, then obviously
+$0\leq h(Y)\leq h(X)\,.$
+We call it the subseries bound.
+2.2 Linear combination, grouping and product of series
+Suppose that $R=\sum_{x\in X}g(x)$ and $S=\sum_{x\in X}h(x)$ are series and that $\alpha,\beta\in\mathbb{C}$. Then the series
 $\alpha R+\beta S=\sum_{x\in X}(\alpha g(x)+\beta h(x))$}
@@
 If $R$ and $S$ absolutely converge and have sums $r$ and $s$, respectively, then $\alpha R+\beta S$ absolutely converges and has sum $\alpha r+\beta s$.
-The next approximation lemma is useful.
+This proposition easily extends to linear combinations with more terms. The next approximation lemma for series is useful.
 Lemma 2.3
-If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, then for every $\varepsilon>0$ there exists a finite set $Y\subset X$, denoted by $Y(S,\,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$,
-$\big|s-\sum_{x\in Z}h(x)\big|\leq\varepsilon\,.$}
+If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, then for every $\varepsilon>0$ there exists a finite set $Y\subset X$, denoted by $Y(S,\,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$ we have
+$|s-h(Z)|\leq\varepsilon\,.$
 Proof. Let an $\varepsilon>0$ be given. For finite $X$ we set $Y=X$. For countable $X$ we take any bijection $f\colon\mathbb{N}\to X$, take an $N\in\mathbb{N}$ such that
@@
 and set $Y=\{f(1),f(2),\dots,f(N)\}$. Then for every finite set $Z$ with $Y\subset Z\subset X$ we have
-$\big|s-\sum_{x\in Z}h(x)\big|\leq\big|s-\sum_{x\in Y}h(x)\big|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$}
+$|s-h(Z)|\leq|s-h(Y)|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$}
 $\Box$
-Two operations on series are grouping and product.
+Two important operations on series are grouping and product.
 A partition of a set $X$ is a set $P$ of nonempty and disjoint sets such that $\bigcup P=X$.
 If $X$ is at most countable, then so is $P$ and every set $Z\in P$.
-If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every $Z\in P$ the subseries $\sum_{x\in Z}h(x)$ absolutely converges and has sum $s_{Z}$, then the series
+If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every $Z\in P$ the subseries $S_{Z}=\sum_{x\in Z}h(x)$ absolutely converges and has sum $s_{Z}$, then the series
 $S_{P}=\sum_{Z\in P}s_{Z}$}
@@
 Proposition 2.4
-If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, and if $P$ is a partition of $X$, then $S_{P}$ is correctly defined, absolutely converges and has the same sum $s$.
+If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, and if $P$ is a partition of $X$, then the grouping series $S_{P}$ is correctly defined, absolutely converges and has the same sum $s$.
 Proof. For every $Z\in P$ the series $S_{Z}=\sum_{x\in Z}h(x)$ is a subseries of $S$ and absolutely converges. Thus the series $S_{P}$ is correctly defined. We show that it absolutely converges. Let
-$c=\sup(\{\sum_{x\in Z}|h(x)|\colon\;\text{$Z\subset X$ and $Z$ is finite}\})\ \ (<+\infty)\,.$}
-Let $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. We use Lemma 2.3 and take for $i\in[n]$ finite sets
-$Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i})\ \ (\subset Z_{i})\,.$
-We form the disjoint union $Z_{0}=Z_{1}^{\prime}\cup\dots\cup Z_{n}^{\prime}$ ($\subset X$). Then
-${\textstyle\sum_{i=1}^{n}\big|\sum_{x\in Z_{i}}h(x)\big|}$ $\leq$ ${\textstyle\sum_{i=1}^{n}\big|\sum_{x\in Z_{i}}h(x)-\sum_{x\in Z_{i}^{\prime}}h(x)\big|\,+}$
-$+$ ${\textstyle\sum_{x\in Z_{0}}|h(x)|\leq\sum_{i=1}^{n}2^{-i}+c\leq 1+c}$
-and $S_{P}$ absolutely converges.
+$c=\sup(\{\sum_{x\in Z}|h(x)|\colon\;\text{$Z\subset X$ and $Z$ is finite}\})\ \ (<+\infty)$}
+and $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. We use Lemma 2.3, take for $i=1,2,\dots,n$ finite sets
+$Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i})\ \ (\subset Z_{i})$
+and define the finite set $Z_{0}=Z_{1}^{\prime}\cup Z_{2}^{\prime}\cup\dots\cup Z_{n}^{\prime}$ ($\subset X$). Then
+${\textstyle\sum_{i=1}^{n}|h(Z_{i})|}$ $\leq$ ${\textstyle\sum_{i=1}^{n}|h(Z_{i})-h(Z_{i}^{\prime})|+\sum_{x\in Z_{0}}|h(x)|}$
+$\leq$ ${\textstyle\sum_{i=1}^{n}2^{-i}+c\leq 1+c}$
+and we see that the series $S_{P}$ absolutely converges.
 Let $t$ be the sum of $S_{P}$. We show that $|t-s|\leq\varepsilon$ for every $\varepsilon>0$; then $t=s$. Let an $\varepsilon>0$ be given. We use Lemma 2.3 and take finite sets
@@
 $P^{\prime}\subset\{Z_{1},\,\dots,\,Z_{n}\}\,\text{ and }\,X^{\prime}\subset\bigcup_{i=1}^{n}Z_{i}\,.$}
-We again use Lemma 2.3 and take for $i\in[n]$ finite sets
+We again use Lemma 2.3 and take for $i=1,2,\dots,n$ finite sets
 $Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i}\cdot\frac{\varepsilon}{3})\ \ (\subset Z_{i})\,.$}
-For $i\in[n]$ let
+For $i=1,2,\dots,n$ let
 $Z_{i}^{\prime\prime}=Z_{i}^{\prime}\cup(X^{\prime}\cap Z_{i})\ \ (\subset Z_{i})\,.$
@@
 $X_{0}=\bigcup_{i=1}^{n}Z_{i}^{\prime\prime}\ \ (\subset X)\,.$}
-Then $X_{0}$ is finite and $X^{\prime}\subset X_{0}$. Also, $Z_{i}^{\prime}\subset Z_{i}^{\prime\prime}$ for every $i\in[n]$. We have
-$|s-t|$ $\leq$ ${\textstyle|s-\sum_{x\in X_{0}}h(x)|+\sum_{i=1}^{n}|\sum_{x\in Z_{i}^{\prime\prime}}h(x)-\sum_{x\in Z_{i}}h(x)|\,+}$
-$+$ ${\textstyle|\sum_{i=1}^{n}\sum_{x\in Z_{i}}h(x)-t|\leq\frac{\varepsilon}{3}+\sum_{i=1}^{n}2^{-i}\cdot\frac{\varepsilon}{3}+\frac{\varepsilon}{3}\leq\varepsilon\,.}$
+Then $X_{0}$ is a finite set and $X^{\prime}\subset X_{0}$. Also, $Z_{i}^{\prime}\subset Z_{i}^{\prime\prime}$ for every $i=1,2,\dots,n$. We have
+$|s-t|$ $\leq$ ${\textstyle|s-h(X_{0})|+\sum_{i=1}^{n}|h(Z_{i}^{\prime\prime})-h(Z_{i})|+|\sum_{i=1}^{n}h(Z_{i})-t|}$
+$\leq$ ${\textstyle\frac{\varepsilon}{3}+\sum_{i=1}^{n}2^{-i}\cdot\frac{\varepsilon}{3}+\frac{\varepsilon}{3}\leq\varepsilon\,.}$
 $\Box$
-Let $X_{1}$, $X_{2}$ be sets and $X_{1}\times X_{2}$ be their Cartesian product. If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, are series, then the series
-$S_{1}\cdot S_{2}=\sum_{\langle x,\,y\rangle\in X_{1}\times X_{2}}h_{1}(x)h_{2}(y)$}
-is the product of $S_{1}$ and $S_{2}$.
+If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$ are series, then the series
+$S_{1}\times S_{2}=\sum_{\langle x,\,y\rangle\in X_{1}\times X_{2}}h_{1}(x)\cdot h_{2}(y)$}
+is called the product of the series $S_{1}$ and $S_{2}$.
 Proposition 2.5
-If $S_{1}$ and $S_{2}$ absolutely converge and have sums $s_{1}$ and $s_{2}$, respectively, then $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s_{1}s_{2}$.
+If $S_{1}$ and $S_{2}$ absolutely converge and have sums $s_{1}$ and $s_{2}$, respectively, then the product series $S_{1}\times S_{2}$ absolutely converges and has the sum $s_{1}s_{2}$.
 Proof.
-We show that $S_{1}\cdot S_{2}$ absolutely converges.
-Let $c\geq 0$ be such that for any finite sets $Y_{i}\subset X_{i}$, $i=1,2$, we have $\sum_{x\in Y_{i}}|h_{i}(x)|\leq c$.
+We show that the series $S_{1}\times S_{2}$ absolutely converges.
+Let $c\geq 0$ be such that for any finite sets $Y_{i}\subset X_{i}$, $i=1,2$ we have $|h_{i}|(Y_{i})\leq c$.
 Let $A\subset X_{1}\times X_{2}$ be a finite set.
-We take any finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
+We take finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
 $\sum_{\langle x,\,y\rangle\in A}|h_{1}(x)h_{2}(y)|\leq\sum_{x\in Y_{1}}|h_{1}(x)|\cdot\sum_{y\in Y_{2}}|h_{2}(y)|\leq c\cdot c=c^{2}\,.$}
-We show that the sum $s$ of $S_{1}\cdot S_{1}$ equals $s_{1}s_{2}$.
-It suffices to prove that for any $\varepsilon\in(0,1)$ we have $|s-s_{1}s_{2}|\leq\varepsilon$.
+We show that the sum $s$ of $S_{1}\times S_{2}$ equals $s_{1}s_{2}$.
+It suffices to prove that for every $\varepsilon\in(0,1)$ we have $|s-s_{1}s_{2}|\leq\varepsilon$.
 So let $\varepsilon\in(0,1)$.
@@
 and
-$Z=Y(S_{1}\cdot S_{2},\frac{\varepsilon}{4})\ \ (\subset X_{1}\times X_{2})\,.$}
+$Z=Y(S_{1}\times S_{2},\frac{\varepsilon}{4})\ \ (\subset X_{1}\times X_{2})\,.$}
 We take finite sets $X_{1}^{\prime\prime}$ and $X_{2}^{\prime\prime}$ such that $X_{1}^{\prime}\subset X_{1}^{\prime\prime}\subset X_{1}$, $X_{2}^{\prime}\subset X_{2}^{\prime\prime}\subset X_{2}$ and $Z\subset X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}$.
-Then $|s-s_{1}s_{2}|$ is at most
-${\textstyle\big|s-\sum_{\langle x,\,y\rangle\in X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}}h_{1}(x)h_{2}(y)\big|+\big|\sum_{x\in X_{1}^{\prime\prime}}h_{1}(x)\sum_{y\in X_{2}^{\prime\prime}}h_{2}(y)-s_{1}s_{2}\big|}$
-${\textstyle\leq\frac{\varepsilon}{4}+|(s_{1}+t_{1})(s_{2}+t_{2})-s_{1}s_{2}|\,,}$
+Then
+$|s-s_{1}s_{2}|$ $\leq$ ${\textstyle\big|s-\sum_{\langle x,\,y\rangle\in X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}}h_{1}(x)h_{2}(y)\big|+|h_{1}(X_{1}^{\prime\prime})h_{2}(X_{2}^{\prime\prime})-s_{1}s_{2}|}$
+$\leq$ ${\textstyle\frac{\varepsilon}{4}+|(s_{1}+t_{1})(s_{2}+t_{2})-s_{1}s_{2}|\,,}$
 where $|t_{1}|\leq\frac{\varepsilon}{4(1+|s_{2}|)}$ and $|t_{2}|\leq\frac{\varepsilon}{4(1+|s_{1}|)}$.
@@
 $\Box$
-Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i\in[k]$, be $k\in\mathbb{N}$ series. We define their product as the series
-$\prod_{i=1}^{k}S_{i}=\sum_{\overline{x}\in\overline{X}}h_{1}(x_{1})h_{2}(x_{2})\dots h_{k}(x_{k})\,$,}
-where $\overline{x}=\langle x_{1},\dots,x_{k}\rangle\in X_{1}\times\dots\times X_{k}=\overline{X}$. Iterating Proposition 2.5 we obtain the following result.
+Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2,\dots,k$ be series. Their product is the series
+$S_{1}\times S_{2}\times\dots\times S_{k}=\sum_{\overline{x}\in\overline{X}}h_{1}(x_{1})\cdot h_{2}(x_{2})\cdot\ldots\cdot h_{k}(x_{k})\,$,}
+where $\overline{x}=\langle x_{1},\dots,x_{k}\rangle$ and $\overline{X}=X_{1}\times\dots\times X_{k}$. By iterating Proposition 2.5 we get the following result.
 Proposition 2.6
-If series $S_{1}$, $\dots$, $S_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
+Let $S_{1}$, $S_{2}$, $\dots$, $S_{k}$ be absolutely convergent series with respective sums $s_{1}$, $s_{2}$, $\dots$, $s_{k}$.
+Then the series $S_{1}\times S_{2}\times\dots\times S_{k}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
 Interestingly, a reversal of Proposition 2.5 holds.
 Corollary 2.7
-If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if their product $S_{1}\cdot S_{2}$ absolutely converges and has sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
+Let $S_{1}$ and $S_{2}$ be nonempty series that are not identically zero.
+If $S_{1}\times S_{2}$ absolutely converges and has sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge, and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
 Proof.
@@
 $\sum_{\langle x,y\rangle\in X}|h_{1}(x)h_{2}(y)|\leq c\,.$}
-We prove absolute convergence of $S_{1}$. We take any $y\in X_{2}$ with $h_{2}(y)\neq 0$. Then for every finite set $X\subset X_{1}$ we have
-$\sum_{x\in X}|h_{1}(x)|=\frac{1}{|h_{2}(y)|}\sum_{\langle x,y\rangle\in X\times\{y\}}|h_{1}(x)h_{2}(y)|\leq\frac{c}{|h_{2}(y)|}\,.$}
-Thus $S_{1}$ absolutely converges. For $S_{2}$ we argue similarly. $\Box$
-It is clear that the assumptions on $S_{1}$ and $S_{2}$ cannot be omitted.
+We show that $S_{1}$ absolutely converges. Let $y_{0}\in X_{2}$ be such that $h_{2}(y_{0})\neq 0$. Then for every finite set $X\subset X_{1}$ we have
+$|h_{1}|(X)=\frac{1}{|h_{2}(y_{0})|}\sum_{\langle x,y\rangle\in X\times\{y_{0}\}}|h_{1}(x)h_{2}(y)|\leq\frac{c}{|h_{2}(y_{0})|}\,.$}
+Thus $S_{1}$ absolutely converges. For the series $S_{2}$ we argue similarly. $\Box$
+It is clear that the assumptions on series $S_{1}$ and $S_{2}$ cannot be omitted.
+2.3 Slim weights
 Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$.
 We obtain a sufficient condition for lightness of the weight $h$ in terms of its values.
-We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk with nonzero weight and length at most $n$.
-Definition 2.8 (slimness)
-A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is slim if for every $n\in\mathbb{N}_{0}$ there is a constant $c=c(n)>0$ such that for every vertex $u\in V(h,n)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
+We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk with length at most $n$ and nonzero weight.
+Definition 2.8 (slim weights)
+A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is slim if for every $n$ in $\mathbb{N}_{0}$ there is a constant $c=c(n)>0$ such that for every vertex $u\in V(h,n)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
 $\sum_{v\in X}|h(\{u,\,v\})|\leq c\,.$}
@@
 Every slim weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is light.
-Proof. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be slim. We show by induction on $n\in\mathbb{N}_{0}$ that the series
-$\sum_{w\in W(n)}h(w)$}
-absolutely converges. For $n=0$ it is trivial because $W(0)=\{\langle 1\rangle\}$. Let $n>0$ and let $c>0$ be a constant such that
+Proof.
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be slim.
+We show by induction on $n\in\mathbb{N}_{0}$ that the series $\sum_{w\in D_{n}}h(w)$ absolutely converges.
+For $n=0$ it is trivial because $D_{0}=\{\langle 1\rangle\}$.
+Let $n>0$ and let $c>0$ be a constant such that
 $\sum_{w\in X}|h(w)|\leq c\,\text{ and }\,\sum_{v\in X^{\prime}}|h(\{u,\,v\})|\leq c$}
-for every finite set $X\subset W(n-1)$, every $u\in V(h,n-1)$ and every finite set $X^{\prime}\subset\mathbb{N}\setminus\{u\}$.
-Let $X\subset W(n)$ be finite.
+for every finite set $X\subset D_{n-1}$, every $u\in V(h,n-1)$ and every finite set $X^{\prime}\subset\mathbb{N}\setminus\{u\}$.
+Let $X\subset D_{n}$ be a finite set.
 We may assume that all walks in $X$ have nonzero weights.
-We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in W(n-1)$ and $\ell(w)$ is the last vertex.
+We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in D_{n-1}$ and $\ell(w)$ is the last vertex.
 Let $Y$ be the set of walks $w^{\prime}$.
@@
 ${\textstyle\sum_{w\in X}|h(w)|}$ $=$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot\sum_{v\in X(w^{\prime})}|h(\{\ell(w^{\prime}),\,v\})|}$
-$\leq$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot c\leq c^{2}\,.}$
+$\leq$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot c\leq c\cdot c=c^{2}\,.}$
 $\Box$
@@
 $G(h)=\langle V(h),\,E(h)\rangle$
-as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. It follows that if $G(h)$ has finite degrees then $h$ is slim and hence light.
-Definition 2.10 ($\xi$-convexity)
-Let $\xi\in\mathbb{C}$. We call a light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ $\xi$-convex if for every vertex $u\in V(h)$ the sum
+as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight.
+If $G(h)$ has finite vertex degrees, then every series $\sum_{w\in D_{n}}h(w)$ has only finitely many nonzero summands and $h$ is therefore light.
+2.4 $\xi$-convex weights
+We generalize the concept of convexity.
+Definition 2.10 ($\xi$-convex weights)
+Suppose that $\xi\in\mathbb{C}$. We say that a light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $\xi$-convex if for every vertex $u\in V(h)$, the sum
 $\sum_{e\in E(h),\,e\ni u}h(e)=\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,\,v\})=\xi\,.$}
-We say that $h$ is convex if it is $1$-convex.
-We show that the sums in the definition exist.
+We say that a light weight is convex if it is $1$-convex.
+Note that 0-convex weights are the ones that satisfy Kirchhoff’s second law. We show that the sums in the definition exist.
 Proposition 2.11
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be light and $u\in V(h)$. Then the series
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and let $u\in V(h)$. Then the series
 $\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})$}
@@
 Let $c=h(w_{0})$ ($\neq 0$) and $W$ be the set of walks $\langle u_{0},\dots,u_{n},u_{n+1}\rangle$ in $K_{\mathbb{N}}$.
-Thus vertices $u_{0}$, $\dots$, $u_{n}$ are fixed and $u_{n+1}$ runs in $\mathbb{N}\setminus\{u_{n}\}=\mathbb{N}\setminus\{u\}$.
-The series $\sum_{w\in W}h(w)$ is a subseries of $\sum_{w\in W(n+1)}h(w)$ and therefore absolutely converges.
-By Proposition 2.2,
+Thus, the vertices $u_{0}$, $\dots$, $u_{n}$ are fixed, and $u_{n+1}$ runs in $\mathbb{N}\setminus\{u_{n}\}=\mathbb{N}\setminus\{u\}$.
+The series $\sum_{w\in W}h(w)$ is a subseries of $\sum_{w\in D_{n+1}}h(w)$ and therefore absolutely converges.
+By Proposition 2.2, the series
 $\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})=\frac{1}{c}\sum_{w\in W}h(w)$}
 absolutely converges. $\Box$
-We survey results on generating functions.
-We use univariate formal power series $U=U(x)=\sum_{n\geq 0}u_{n}x^{n}$ with complex coefficients.
-They form an integral domain denoted by $\mathbb{C}[[x]]$.
-For $U$ and $V$ in $\mathbb{C}[[x]]$,
-$U+V=\sum_{n\geq 0}(u_{n}+v_{n})x^{n}\,\text{ and }\,U\cdot V=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
-Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}u_{n}x^{n}$ with $u_{0}\neq 0$.
-Let $\mathbb{C}_{1}=\{x\in\mathbb{C}\colon\;|x|=1\}$ and $\xi\in\mathbb{C}_{1}$. We define
+We proceed to generating functions.
+2.5 Sums $U(1)_{1}$ and $U(1)_{2}$ for a generating function $U(x)$
+We use univariate formal power series
+$U=U(x)=\sum_{n\geq 0}u_{n}x^{n}$}
+with complex coefficients. They form an integral domain denoted by $\mathbb{C}[[x]]$. For $U=\sum_{n\geq 0}u_{n}x^{n}$ and $V=\sum_{n\geq 0}v_{n}x^{n}$ in $\mathbb{C}[[x]]$,
+$U+V=\sum_{n\geq 0}(u_{n}+v_{n})x^{n}\,\text{ and }\,U\cdot V=UV=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
+Neutral elements are $0=0x^{0}+0x^{1}+0x^{2}+\dots$ and $1=1x^{0}+0x^{1}+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}u_{n}x^{n}$ with $u_{0}\neq 0$.
+Let $\mathbb{C}_{1}=\{x\in\mathbb{C}\colon\;|x|=1\}$ and $\xi\in\mathbb{C}_{1}$. We define the set
 $\xi\mathbb{R}_{\geq 0}=\{\xi x\colon\;x\in\mathbb{R}_{\geq 0}\}\ \ (\subset\mathbb{C})\,.$
-It is the ray in $\mathbb{C}$ going from the origin in the direction $\xi$.
-Clearly, $\xi\mathbb{R}_{\geq 0}$ is closed under addition.
-If $\alpha_{j}\in\xi_{j}\mathbb{R}_{\geq 0}$, $j=1,2$, then $\alpha_{1}\alpha_{2}\in(\xi_{1}\xi_{2})\mathbb{R}_{\geq 0}$.
+It is the ray in $\mathbb{C}$ emanating from the origin in the direction $\xi$.
+It is closed under addition.
+If $\alpha_{j}\in\xi_{j}\mathbb{R}_{\geq 0}$, $j=1,2$ then $\alpha_{1}\alpha_{2}\in(\xi_{1}\xi_{2})\mathbb{R}_{\geq 0}$.
 If
 $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$
-is a light weight and $n\in\mathbb{N}_{0}$ then $h(W(n))\in\xi^{n}\mathbb{R}_{\geq 0}$.
+is a light weight, $n\in\mathbb{N}_{0}$ and $X\subset D_{n}$, then $h(X)\in\xi^{n}\mathbb{R}_{\geq 0}$.
 The lines $\xi\mathbb{R}$ in $\mathbb{C}$ going through the origin in the direction $\xi$ have the same properties.
@@
     $\sum_{n\geq 0}u_{n}1^{n}=\sum_{n\geq 0}u_{n}$}
-    absolutely converges, we denote the sum by $U(1)_{2}$ and say that $U(1)_{2}$ exists.
+    absolutely converges, we denote its sum by $U(1)_{2}$ and say that $U(1)_{2}$ exists.
 If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$. The opposite implication is, in general, not true. For example, if
@@
 $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]$}
-and $U(1)_{1}$ exists, then $U(1)_{2}$ exists and $U(1)_{1}=U(1)_{2}$. The following is a corollary of the well known Hadamard’s formula for the radius of convergence of a formal power series.
+and $U(1)_{1}$ exists, then $U(1)_{2}$ exists and $U(1)_{1}=U(1)_{2}$. The following is a corollary of Hadamard’s formula for the radius of convergence.
 Proposition 2.13
@@
 Dividing by $\xi^{n}$ and taking the limit for $n\to\infty$, we get the stated formula. $\Box$
-We manipulate the quantities $U(1)_{1}$ by means of three Abelian theorems, which we state only for the radius of convergence $R\geq 1$.
-The first theorem is classical, see [1, 2] and [29, Chapter II.7].
+2.6 Three Abelian theorems
+We manipulate the quantities $U(1)_{1}$ by means of three Abelian theorems.
+We state them only for the radius of convergence at least $1$.
+The first theorem is classical, see [1, 2] and [30, Chapter II.7].
 Theorem 2.16 (Abel 1)
 Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. If $U(1)_{1}$ exists, so that $U(x)\in H_{1}$ by Proposition 2.13, then
-$\lim_{x\nearrow 1}F_{U}(x)=U(1)_{1}\ \ (\in\mathbb{C})\,.$
-The second Abelian theorem is rarely mentioned, but see [2] for a simple proof. It is given only for the case $U(1)_{1}=+\infty$, but it works for the sum and limit -$\infty$ as well.
+$\lim_{x\nearrow 1}U(x)=U(1)_{1}\ \ (\in\mathbb{C})\,.$
+The second theorem is seldom mentioned, but see [2] for a simple proof. It is given there only in the case $U(1)_{1}=+\infty$, but it works for the sum and limit -$\infty$ as well.
 Theorem 2.17 (Abel 2)
 Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}[[x]]\cap H_{1}$. If $U(1)_{1}=\pm\infty$ then
-$\lim_{x\nearrow 1}F_{U}(x)=\pm\infty\,$,
-with equal signs.
-The third Abelian theorem allows only nonnegative real coefficients.
-It has the advantage over the previous two theorems that it is an equivalence.
-Now one can go not only from the sum to the functional limit, but also in the reverse.
+$\lim_{x\nearrow 1}U(x)=U(1)_{1}=\pm\infty\,.$
+The third theorem admits only nonnegative real coefficients.
+It has two advantages over the previous Abelian theorems: it works for every $U(x)$ in $\mathbb{R}_{\geq 0}[[x]]$ and it is an equivalence, one can go not only from the sum to the limit, but also in the reverse.
 We omit the easy proof.
 Theorem 2.18 (Abel 3)
-Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]\cap H_{1}$. Then the following limit and sum always exist (and are possibly $\infty$) and are equal,
-$\lim_{x\nearrow 1}F_{U}(x)=U(1)_{1}\ \ (\in[0,\,+\infty)\cup\{+\infty\})\,.$
+Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]\cap H_{1}$. Then the following limit and sum always exist, are possibly $\infty$ and are equal,
+$\lim_{x\nearrow 1}U(x)=U(1)_{1}\ \ (\in[0,\,+\infty)\cup\{+\infty\})\,.$
+2.7 Conversion of identities in conditional convergence
+We believe that the precise statements and proofs of conversions of formal identities to identities at $x=1$, given in this and the next subsection, is our original contribution to the subject.
 Proposition 2.19
@@
 $\sum_{j=0}^{n}u_{j}=\sum_{j=0}^{n}(\alpha v_{j}+\beta w_{j})=\alpha\sum_{j=0}^{n}v_{j}+\beta\sum_{j=0}^{n}w_{j}\,.$}
-Taking the limit for $n\to\infty$ we get that that $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$. $\Box$
-Here the formulation “then $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$” abbreviates “then $U(1)_{1}$ exists and $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$”.
-We will use such abbreviated formulations.
-This result can be easily generalized to linear combinations with more than two terms.
-Next we consider products of formal power series.
+Taking the limit for $n\to\infty$ we get $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$. $\Box$
+In this proposition, the formulation “then $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$” abbreviates the formulation “then $U(1)_{1}$ exists and $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$”.
+We freely use such abbreviations.
+This result easily generalizes to linear combinations with more terms.
 Proposition 2.20
-Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$ be $n+1$, $n\in\mathbb{N}$, formal power series.
-If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
+Let $U(x)$, $V_{1}(x)$, $\dots$, $V_{n}(x)\in\mathbb{C}[[x]]$ be $n+1\geq 2$ formal power series.
+If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and if $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
 $U(1)_{1}=\prod_{i=1}^{n}V_{i}(1)_{1}\,.$}
 Proof. By Propositions 2.4, 2.6 and 2.13, and by the assumptions we have
-$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}_{<1}$
+$U(x)=\prod_{i=1}^{n}V_{i}(x)$ for every number $x\in\mathbb{C}_{<1}$
 (see proofs of Propositions 2.24 and 2.27 for details of similar arguments). Using Theorem 2.16 and properties of limits of functions we get
-$U(1)_{1}$ $=$ $\lim_{x\nearrow 1}F_{U}(x)=\lim_{x\nearrow 1}{\textstyle\prod_{i=1}^{n}F_{V_{i}}(x)}$
-$=$ ${\textstyle\prod_{i=1}^{n}\lim_{x\nearrow 1}F_{V_{i}}(x)=\prod_{i=1}^{n}V_{i}(1)_{1}}\,.$
+$U(1)_{1}$ $=$ $\lim_{x\nearrow 1}U(x)=\lim_{x\nearrow 1}{\textstyle\prod_{i=1}^{n}V_{i}(x)}$
+$=$ ${\textstyle\prod_{i=1}^{n}\lim_{x\nearrow 1}V_{i}(x)=\prod_{i=1}^{n}V_{i}(1)_{1}}\,.$
 $\Box$
@@
 $U(1)_{1}=\frac{1}{V(1)_{1}}\,.$
-Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. Proposition 2.20 gives that $U(1)_{1}V(1)_{1}=1$. Thus $V(1)_{1}\neq 0$ and dividing by $V(1)_{1}$ gives the stated formula. $\Box$
+Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. Proposition 2.20 gives that $U(1)_{1}V(1)_{1}=1$. Thus $V(1)_{1}\neq 0$, and by dividing by it we get the stated identity. $\Box$
 Proposition 2.22
@@
 Proof.
-In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$.
+Formally $U(x)^{2}\cdot W(x)=W(x)-V(x)$.
 Let T(x)=W(x)V(x)T(x)=W(x)-V(x).
@@
 $\Box$
-We obtain analogs of the previous results in absolute convergence.
+2.8 Conversion of identities in absolute convergence
+We obtain analogs of the results in the previous subsection for absolute convergence.
 Proposition 2.23
@@
 Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$.
-Since $V(1)_{2}$ and $W(1)_{2}$ are sums of absolutely convergent series $\sum_{n\in\mathbb{N}_{0}}v_{n}$ and $\sum_{n\in\mathbb{N}_{0}}w_{n}$, respectively, and $u_{n}=\alpha v_{n}+\beta w_{n}$, the result follows from Proposition 2.2.
+Since $V(1)_{2}$ and $W(1)_{2}$ are sums of absolutely convergent series $\sum_{n\in\mathbb{N}_{0}}v_{n}$ and $\sum_{n\in\mathbb{N}_{0}}w_{n}$, respectively, and $u_{n}=\alpha v_{n}+\beta w_{n}$, the stated formula follows from Proposition 2.2.
 $\Box$
@@
 Proof. 1. We consider the series
-$S=\sum_{\langle j,\,k\rangle\in\mathbb{N}_{0}^{2}}v_{j}w_{k}=\sum_{j\in\mathbb{N}_{0}}v_{j}\cdot\sum_{k\in\mathbb{N}_{0}}w_{k}\,.$}
-By Proposition 2.5 it absolutely converges and has sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping
-$\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}\,$,}
-where the inner $\sum$ means sum, the outer $\sum$ means series, and $\sum_{j=0}^{n}v_{j}w_{n-j}=u_{n}$.
-By Proposition 2.4, $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}$.
-Thus $U(1)_{2}=V(1)_{2}W(1)_{2}$.
+$S=\sum_{\langle j,\,k\rangle\in\mathbb{N}_{0}^{2}}v_{j}w_{k}=\sum_{j\in\mathbb{N}_{0}}v_{j}\times\sum_{k\in\mathbb{N}_{0}}w_{k}\,.$}
+By Proposition 2.5, the series $S$ absolutely converges and has the sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping
+$\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}=\sum_{n\geq 0}u_{n}\,.$}
+Here the inner $\sum$ means sum and the outer $\sum$ means series. By Proposition 2.4, the series $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}=V(1)_{2}W(1)_{2}$.
 2. This follows from the inequality ($n\in\mathbb{N}_{0}$)
@@
 $\Box$
-For nonnegative real coefficients we can go in the reverse, from the product to the factors.
+For nonnegative coefficients one can go from the product to the factors.
 Proposition 2.25
-Let $U(x)$, $V(x)$ and $W(x)$ be in $\mathbb{R}_{\geq 0}[[x]]$.
-Suppose that $U(x)=V(x)W(x)$, that $U(x)$, $V(x)$ and $W(x)$ are in $H_{1}$, $V(x)W(x)\neq 0$ and that $U(1)_{2}$ exists.
-Then $V(1)_{2}$ and $W(1)_{2}$ exist, and
+Let $U(x)$, $V(x)$ and $W(x)$ be in $\mathbb{R}_{\geq 0}[[x]]\cap H_{1}$. If $U(x)=V(x)W(x)$, $V(x)W(x)\neq 0$ and $U(1)_{2}$ exists, then $V(1)_{2}$ and $W(1)_{2}$ exist, and
 $U(1)_{2}=V(1)_{2}W(1)_{2}\,.$
-Proof. By Theorem 2.18 and the assumption $V(x)W(x)\neq 0$ we have $V(1)_{2}=V(1)_{1}=\lim_{x\nearrow 1}F_{V}(x)>0$ and the same for $W(x)$. Since $F_{U}=F_{V}F_{W}$ on $\mathbb{C}_{<1}$, we have
-$U(1)_{2}$ $=$ $U(1)_{1}=\lim_{x\nearrow 1}F_{U}(x)$
-$=$ $\lim_{x\nearrow 1}F_{V}(x)\cdot\lim_{x\nearrow 1}F_{W}(x)=V(1)_{1}W(1)_{1}=V(1)_{2}W(1)_{2}\,.$
-In the second equality we used Theorem 2.18. The third equality holds because the product of limits is neither $0\cdot(+\infty)$ nor $(+\infty)\cdot 0$. $\Box$
+Proof.
+By Theorem 2.18 and by the assumption $V(x)W(x)\neq 0$, we have $V(1)_{2}=V(1)_{1}=\lim_{x\nearrow 1}V(x)>0$, and the same applies for $W(x)$.
+Since $U(x)=V(x)W(x)$ for every $x\in\mathbb{C}_{<1}$, we have
+$(+\infty>)\ U(1)_{2}$ $=$ $U(1)_{1}=\lim_{x\nearrow 1}U(x)=\lim_{x\nearrow 1}(V(x)W(x))$
+$=$ $\lim_{x\nearrow 1}V(x)\cdot\lim_{x\nearrow 1}W(x)=V(1)_{1}W(1)_{1}=V(1)_{2}W(1)_{2}\,.$
+In particular, $V(1)_{2},W(1)_{2}<+\infty$.
+In the second and fifth equalities we used Theorem 2.18.
+In the fourth equality we used the arithmetic of functional limits.
+It holds because the product of limits is neither $0\cdot(+\infty)$ nor $(+\infty)\cdot 0$.
+$\Box$
 Proposition 2.20 for $n=2$ and item 1 of Proposition 2.24 are incomparable.
@@
 The latter assumes stronger absolute convergence, but only for two series.
-Another variant of the proposition is due to F.
+Another variant of this results is due to F.
 Mertens.
-Theorem 2.26 (Mertens’s)
+Theorem 2.26 (Mertens)
 Suppose that $U(x)$, $V(x)$ and $W(x)$ are in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$, and if $V(1)_{1}$ and $W(1)_{2}$ exist, then
@@
 Proposition 2.27
-Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$ and $v_{0}=1$. If $U(x)=\frac{1}{V(x)}$ and $c=|v_{1}|+|v_{2}|+\dots<1$, then the following holds.
+Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$ and $v_{0}=1$. If $U(x)=\frac{1}{V(x)}$ and $c=|v_{1}|+|v_{2}|+\dots<1$, then the following holds.
   1. 1.
-    $V(1)_{2}$ and $U(1)_{2}$ exist, $V(1)_{2}\neq 0$ and $U(1)_{2}=\frac{1}{V(1)_{2}}$.
+    $U(1)_{2}$ and $V(1)_{2}$ exist, $V(1)_{2}\neq 0$ and $U(1)_{2}=\frac{1}{V(1)_{2}}$.
   2. 2.
     The sum $\sum_{n\geq 0}|u_{n}|\leq\frac{1}{1-c}$.
-Proof. 1. From the assumption on $c$ it follows that $V(1)_{2}$ exists. From
+Proof. 1. It follows from the assumption on $c$ that $V(1)_{2}$ exists. From
 $|V(1)_{2}-1|\leq c<1$
@@
 We show that $S$ absolutely converges. This follows from the bound that the sum of absolute values of summands in any finite subseries of $S$ is at most
-$1+\sum_{m\geq 1}(|v_{1}|+|v_{2}|+\dots)^{m}=\frac{1}{1-c}\,.$}
+$1+\sum_{m\geq 1}(|v_{1}|+|v_{2}|+\dots)^{m}=1+\sum_{m\geq 1}c^{m}=\frac{1}{1-c}\,.$}
 Let $m\in\mathbb{N}$. We consider the series
-$S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}v_{j_{2}}\dots v_{j_{m}}\,.$}
-By Proposition 2.6, it absolutely converges and has sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has sum
+$S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,j_{2},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}v_{j_{2}}\dots v_{j_{m}}\,.$}
+$S_{m}$ absolutely converges by Proposition 2.6 and has sum $(V(1)_{2}-1)^{m}$. By Propositions 2.2 and 2.4, the series $S$ has sum
 $1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}\,.$}
@@
 $u_{n}=\sum_{m\geq 1}(-1)^{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=n\end{subarray}}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\,.$}
-Proposition 2.4 shows that $S$ also has sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$. The stated formula follows.
+Proposition 2.4 shows that $S$ has sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$. The stated formula follows.
 2. Indeed, the sum
@@
 $V(x)=vx^{l}\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}\,.$}
-Proof. Let $n\in\mathbb{N}_{0}$ and a,ba,\,b be formal variables. By Vandermonde’s identity ([30]),
+Proof. Let $n\in\mathbb{N}_{0}$ and a,ba,\,b be formal variables. Vandermonde’s identity [31] states that
 $\sum_{j=0}^{n}\binom{a}{j}\binom{b}{n-j}=\binom{a+b}{n}\ \ (\in\mathbb{Q}[a,\,b])\,.$}
-Let $V_{0}(x)=\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}$. For any $v_{0}\in\mathbb{C}$ such that $(v_{0})^{2}=t$ the instance $a=b=\frac{1}{2}$ of the identity gives
+Let $V_{0}(x)=\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}$. For any $v_{0}\in\mathbb{C}$ such that $(v_{0})^{2}=t$ the instance $a=b=\frac{1}{2}$ of Vandermonde’s identity gives
 $\big(v_{0}x^{l}V_{0}(x)\big)^{2}$ $=$ ${\textstyle tx^{2l}\sum_{n\geq 0}\big(\sum_{j=0}^{n}\binom{1/2}{j}\binom{1/2}{n-j}\big)(t_{1}x+t_{2}x^{2}+\dots)^{n}}$
@@
 We consider the series
-$S=0+0+\dots+0+v+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}v\binom{1/2}{m}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}$}
-beginning with $l$ summands 0.
+$S=0+0+\dots+0+v+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}v\binom{1/2}{m}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}\,$,}
+beginning with $l$ zero summands.
 We show that it absolutely converges.
@@
 Since $\big|\binom{1/2}{m}\big|\leq 1$ for every $m\in\mathbb{N}$, the sum of absolute values of summands in $T$ is at most
-$|v|+|v|\sum_{m\geq 1}(|t_{1}|+|t_{2}|+\dots)^{m}=\frac{|v|}{1-c}\,.$}
-Thus $S$ absolutely converges. By Proposition 2.4, the series $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.24, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
+$|v|+|v|\sum_{m\geq 1}(|t_{1}|+|t_{2}|+\dots)^{m}=|v|+|v|\sum_{m\geq 1}c^{m}=\frac{|v|}{1-c}\,.$}
+Thus $S$ absolutely converges. By Proposition 2.4, the series $S$ has sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.24, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
 Proposition 2.30
 Let $U(x),V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ be in $\mathbb{C}[[x]]$.
-Suppose that $V(1)_{2}$ exists, $w_{0}=1$, $U(x)^{2}=1-\frac{V(x)}{W(x)}$ and $c=|w_{1}|+|w_{2}|+\dots<1$.
+Suppose that $V(1)_{2}$ exists, $w_{0}=1$, $U(x)^{2}=1-\frac{V(x)}{W(x)}$ and that $c=|w_{1}|+|w_{2}|+\dots<1$.
 If $V(x)=W(x)$, then $U(x)=0$.
-Else, if $k\in\mathbb{N}_{0}$ is minimum such that $v_{k}\neq w_{k}$ and
+Else, if $k\in\mathbb{N}_{0}$ is minimum such that $v_{k}\neq w_{k}$ and if
 $\big(|w_{k}-v_{k}|\cdot(1-c)\big)^{-1}\sum_{n>k}|w_{n}-v_{n}|<1\,$,}
@@
 $\Box$
-In conclusion, we consider $\xi$-convex light weights.
+2.9 Generating functions of walks for $\xi$-convex weights
+We generalize the result in Section 1.2 that $h(D_{n})=1$ for every $n$.
 Proposition 2.31
 Let $\xi\in\mathbb{C}$. For every $\xi$-convex light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ we have
-$\sum_{n=0}^{\infty}h(W(n))x^{n}=\frac{1}{1-\xi x}\,.$}
-In particular, for $\xi=0$ we have $\sum_{n=0}^{\infty}h(W(n))x^{n}=1$.
+$\sum_{n\geq 0}h(D_{n})x^{n}=\frac{1}{1-\xi x}\ \ (\in\mathbb{C}[[x]])\,.$}
+Thus for $\xi=0$ we have $\sum_{n\geq 0}h(D_{n})x^{n}=1$ ($1x^{0}+0x^{1}+0x^{2}+\dots$).
 Proof.
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be light and $\xi$-convex.
-We prove by induction on $n\in\mathbb{N}_{0}$ that $h(W(n))=\xi^{n}$.
-For $n=0$ this holds as $W(0)=\{\langle 1\rangle\}$.
+Let $\xi\in\mathbb{C}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light and $\xi$-convex weight.
+We prove by induction on $n\in\mathbb{N}_{0}$ that $h(D_{n})=\xi^{n}$.
+For $n=0$ this holds, $h(D_{0})=1$, because $D_{0}=\{\langle 1\rangle\}$.
 Let $n>0$.
 Recall that $\ell(w)$ denotes the last vertex of a walk $w$.
-Using Propositions 2.2 and 2.4, the $\xi$-convexity assumption and induction, we get
-$h(W(n))$ $=$ ${\textstyle\sum_{w\in W(n)}h(w)}$
-$=$ ${\textstyle\sum_{\begin{subarray}{c}w^{\prime}\in W(n-1)\\ \ell(w^{\prime})\in V(h)\end{subarray}}h(w^{\prime})\sum_{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}}h(\{\ell(w^{\prime}),\,v\})}$
-$=$ ${\textstyle\sum_{w^{\prime}\in W(n-1)}h(w^{\prime})\cdot\xi=\xi^{n-1}\cdot\xi=\xi^{n}\,.}$
-We can add and remove the condition $\ell(w^{\prime})\in V(h)$ without affecting the sum because walks $w^{\prime}$ not satisfying it have zero weight. $\Box$
-3 Extensions of Pólya’s theorem with $v=1$
-We begin with the case $v=1$ when the vertex to be visited by walks coincides with their starting vertex. In Theorems 3.23.7 we establish the first group of generalizations of Pólya’s theorem.
-For that end, we need some generating functions.
-Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$, and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight.
-Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the subset of walks visiting $v$ in some step $i>0$.
-We are interested in the generating function
-$A_{h}(x)=\sum_{n\geq 0}a_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(1,\,n))x^{n}$}
-and define three more generating functions
-$B_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h}x^{n}=\sum_{n\geq 0}h(W_{b}(n))x^{n}\,,}$
-$C_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h}x^{n}=\sum_{n\geq 0}h(W_{c}(n))x^{n}\,\text{ and}}$
-$D_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}d_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(n))x^{n}\,,}$
-all in $\mathbb{C}[[x]]$.
-Here $W_{b}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and $W_{c}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and avoid $1$ between.
-The coefficients $a_{n}^{h}$, $\dots$, $d_{n}^{h}$ are correctly defined because they are sums of subseries of the absolutely convergent series
-$d_{n}^{h}=\sum_{w\in W(n)}h(w)\,.$}
-The constant coefficients are $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$.
+Using Propositions 2.2 and 2.4, the assumption of $\xi$-convexity and induction, we get
+$h(D_{n})$ $=$ ${\textstyle\sum_{\begin{subarray}{c}w\in D_{n-1}\\ \ell(w)\in V(h)\end{subarray}}h(w)\cdot\sum_{v\in\mathbb{N}\setminus\{\ell(w)\}}h(\{\ell(w),\,v\})}$
+$=$ ${\textstyle\sum_{w\in D_{n-1}}h(w)\cdot\xi=h(D_{n-1})\cdot\xi=\xi^{n-1}\cdot\xi=\xi^{n}\,.}$
+We can add and remove the condition $\ell(w)\in V(h)$ without affecting the sum because walks $w$ not satisfying it have zero weight. $\Box$
+3 General Pólya’s theorems for $v=1$
+In this section, we generalize the results in Section 1.2 to weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$.
+We first introduce three auxiliary generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$.
+In Sections 3.2 and 3.3, we obtain general Pólya’s theorems for $v=1$ with general and $\xi$-convex weights, respectively.
+3.1 Auxiliary generating functions
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight (Definition 2.1). As we know from Section 1.4, we are interested in the main generating function
+$A_{h}(x)=A_{1,\,h}(x)=\sum_{n\geq 0}h(A_{n})x^{n}\ \ (\in\mathbb{C}[[x]])$}
+for the set $A$ of walks in $K_{\mathbb{N}}$ starting at $1$ and revisiting it later; $h(A_{n})$ is the sum of the series
+$\sum_{w\in A_{n}}h(w)=\sum_{\begin{subarray}{c}w\in A\\ s(w)=n\end{subarray}}h(w)\,.$}
+As in Section 1.2, we consider the set $B$ of closed walks in $K_{\mathbb{N}}$ starting and ending at $1$, the set $C\subset B$ of closed walks with inner vertices different from $1$, and the set $D$ of all walks in $K_{\mathbb{N}}$ starting at $1$.
+We define three auxiliary generating functions
+$B_{h}(x)=\sum_{n\geq 0}h(B_{n})x^{n},\ C_{h}(x)=\sum_{n\geq 0}h(C_{n})x^{n},\ D_{h}(x)=\sum_{n\geq 0}h(D_{n})x^{n}\,.$}
+The coefficients $h(A_{n})$, $h(B_{n})$ and $h(C_{n})$, $n\in\mathbb{N}_{0}$, are correctly defined because we assume that the coefficients $h(D_{n})$ are defined ($h$ is a light weight).
+We have $h(A_{0})=h(C_{0})=0$ and $h(B_{0})=h(D_{0})=1$.
+Let $F\colon X\to Y$ be a map and $Z$ be any set.
+In the next proof we use the notation
+$F[Z]=\{F(x)\colon\;x\in X\cap Z\}\ \ (\subset Y)\,.$
 Proposition 3.1
-Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be the above generating functions. The following relations hold.
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be the above generating functions. The following identities hold in $\mathbb{C}[[x]]$.
   1. 1.
@@
 1.
-We show that $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}\cdot d_{n-j}^{h}$ for every $n\in\mathbb{N}_{0}$.
-For $n\leq 1$ it holds, then $a_{n}^{h}=c_{n}^{h}=0$.
-For $n\geq 2$ the set $W(1,n)$ is countable and we split every walk in it at the first revisit of $1$.
-We get a map
-$F\colon W(1,\,n)\to\bigcup_{j=0}^{n}W_{c}(j)\times W(n-j)$}
+We show that $h(A_{n})=\sum_{j=0}^{n}h(C_{j})\cdot h(D_{n-j})$ for every $n\in\mathbb{N}_{0}$.
+For $n\leq 1$ it holds because then $h(A_{n})=h(C_{n})=0$.
+For $n\geq 2$ the set $A_{n}$ is infinite and countable.
+We split every walk in it at the first revisit of $1$ and get a map
+$F\colon A_{n}\to\bigcup_{j=0}^{n}C_{j}\times D_{n-j}$}
 that is defined as follows. For any walk
-$w=\langle u_{0},\,u_{1},\,\dots,\,u_{j},\,\dots,\,u_{n}\rangle\in W(1,\,n)\,$,
-where $u_{0}=u_{j}=1$, $j>0$ and $u_{i}\neq 1$ for $i\in[j-1]$, we set
-$F(w)=\langle w_{1},\,w_{2}\rangle\ \ (\in W_{c}(j)\times W(n-j))\,$,
-with $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$.
+$w=\langle u_{0},\,u_{1},\,\dots,\,u_{j},\,\dots,\,u_{n}\rangle\in A_{n}\,$,
+where $u_{0}=u_{j}=1$, $j>0$ and $u_{i}\neq 1$ for every $i=1,2,\dots,j-1$, we set
+$F(w)=\langle w_{1},\,w_{2}\rangle\ \ (\in C_{j}\times D_{n-j})\,$,
+where $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$.
 It is easy to see that $F$ is a bijection.
 $F$ is also weight-preserving (WP): for every weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and value $F(w)=\langle w_{1},w_{2}\rangle$, we have $h(w)=h(w_{1})h(w_{2})$.
-If $X$ is a set of pairs of walks in $K_{\mathbb{N}}$, we write $h(X)$ for the sum of the series
+If $X$ is a set of pairs of walks in $K_{\mathbb{N}}$, we denote by $h(X)$ the sum of the series
 $\sum_{\langle w,\,w^{\prime}\rangle\in X}h(w)h(w^{\prime})\,$,}
-if it absolutely converges.
+if the series absolutely converges.
 Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and let $n\geq 2$. Using the map $F$, we get
-$a_{n}^{h}$ $=$ $h(W(1,\,n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W(1,\,n)])$
+$h(A_{n})$ $\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}$ $h(F[A_{n}])$
 […]
@@
 2. We show that
-$B_{h}(x)=\frac{1}{1-C_{h}(x)}=1+\sum_{j\geq 1}(C_{h}(x))^{j}\,.$}
-We recall that $b_{0}^{h}=1$ and $c_{0}^{h}=0$ and show that for every $n\in\mathbb{N}$,
-$b_{n}^{h}=\sum_{j=1}^{\infty}\sum_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}c_{n_{1}}^{h}c_{n_{2}}^{h}\dots c_{n_{j}}^{h}\,.$}
-For odd $n$ it is true: for odd $n$ we have $b^{h}_{n}=c^{h}_{n}=0$, and in every inner summand at least one index $n_{i}$ is odd.
-For even $n\geq 2$ the set $W_{b}(n)$ is countable.
-We split every walk in it at the visits of $1$ and get a map
-$F\colon W_{b}(n)\to\bigcup_{j=1}^{\infty}\bigcup_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}W_{c}(n_{1})\times\cdots\times W_{c}(n_{j})$}
+$B_{h}(x)=\frac{1}{1-C_{h}(x)}=1+\sum_{j\geq 1}(C_{h}(x))^{j}\ \ (\in\mathbb{C}[[x]])\,.$}
+We recall that $h(B_{0})=1$ and show that for every $n\in\mathbb{N}$,
+$h(B_{n})=\sum_{j\geq 1}\sum_{\begin{subarray}{c}n_{1},\,n_{2},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+n_{2}+\dots+n_{j}=n\end{subarray}}h(C_{n_{1}})h(C_{n_{2}})\dots h(C_{n_{j}})\,.$}
+For $n=1$ it holds because $h(B_{1})=h(C_{1})=0$. For $n\geq 2$ the set $B_{n}$ is infinite and countable. We split every walk in it at the visits of $1$ and get a map
+$F\colon B_{n}\to\bigcup_{j\geq 1}\bigcup_{\begin{subarray}{c}n_{1},\,n_{2},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+n_{2}+\dots+n_{j}=n\end{subarray}}C_{n_{1}}\times C_{n_{2}}\times\cdots\times C_{n_{j}}$}
 that is defined as follows. For any walk
-$w=\langle u_{0},\,\dots,\,u_{m_{1}},\,\dots,\,\dots,\,u_{m_{j-1}},\,\dots,\,u_{n}\rangle\in W_{b}(n)\,$,
-where $0=m_{0}<m_{1}<\dots<m_{j}=n$ and $u_{k}=1$ iff $k=0$ or $k=m_{i}$ for some $i\in[j]$, we set
-$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{1})\times W_{c}(n_{2})\times\dots\times W_{c}(n_{j}),\,n_{i}=m_{i}-m_{i-1})\,$,
-with $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i\in[j]$.
-It is easy to see that $F$ is a bijection and that it is weight-preserving, as in part 1.
-If $X$ is a set of tuples of walks, we write $h(X)$ for the sum of the series
-$\sum_{\langle w_{1},\,\dots,\,w_{k})\in X}h(w_{1})\dots h(w_{k})\,$,}
-if it absolutely converges.
-Let $h$ be a light weight and $n\geq 2$ be even. Using the map $F$, we get
-$b_{n}^{h}$ $=$ $h(W_{b}(n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W_{b}(n)])$
+$w=\langle u_{0},\,\dots,\,u_{m_{1}},\,\dots,\,\dots,\,u_{m_{j-1}},\,\dots,\,u_{n}\rangle\in B_{n}\,$,
+where $0=m_{0}<m_{1}<\dots<m_{j}=n$ and $u_{k}=1$ iff $k\in\{0,m_{1},m_{2},\dots,m_{j}\}$, we set
+$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in C_{n_{1}}\times C_{n_{2}}\times\dots\times C_{n_{j}},\,n_{i}=m_{i}-m_{i-1})\,$,
+where $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i=1,2,\dots,j$.
+It is easy to see that $F$ is a bijection and that it is weight-preserving (WP), as in part 1.
+If $X$ is a set of $k$-tuples of walks, we denote by $h(X)$ the sum of the series
+$\sum_{\langle w_{1},\,w_{2},\,\dots,\,w_{k})\in X}h(w_{1})h(w_{2})\dots h(w_{k})\,$,}
+if the series absolutely converges.
+Let $h$ be a light weight and $n\geq 2$. Using the map $F$, we get
+$h(B_{n})$ $\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}$ $h(F[B_{n}])$
 […]
@@
 $\Box$
-We label our fourteen generalizations of Pólya’s theorem for better orientation by the triples
-$\langle x,\,y,\,z\rangle\in\{v=1,\,v\neq 1\}\times\{\mathrm{gen},\,\mathrm{con}\}\times\{U(1)_{1},\,U(1)_{2},\,U(1)_{1}=\infty\}\,$,
-where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to $\xi$-convex ones.
-For $U(1)_{1}$ and $U(1)_{2}$ see Definition 2.12, and for $U(1)_{1}=\infty$ see Definition 2.14.
-The generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$, and $D_{h}(x)$ are defined at the beginning of this section.
-Theorem 3.2 (1: $v=1$, gen, $U(1)_{1}$)
-If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a light weight such that $A_{h}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
+We already mentioned in Section 1.5 that the previous proof differs from all related proofs in the first phase of the symbolic method because we work with infinite sets of combinatorial objects,
+3.2 Pólya’s theorems for $v=1$ and general weights
+We obtain general Pólya’s theorems of the form indicated in the subsection heading.
+For the labels $\exists\,U(1)_{1}$ and $\exists\,U(1)_{2}$ of the theorems see Definition 2.12, and for the label $U(1)_{1}=\infty$ see Definition 2.14.
+The generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$, and $D_{h}(x)$ are defined in the previous subsection.
+Theorem 3.2 ($\exists\,U(1)_{1}$)
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. If $A_{h}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
 $A_{h}(1)_{1}=\Big(1-\frac{1}{B_{h}(1)_{1}}\Big)D_{h}(1)_{1}\,.$
@@
 The stated formula follows. $\Box$
-Theorem 3.3 (2: $v=1$, gen, $U(1)_{2}$)
-If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a light weight such that $\sum_{n\geq 1}|b_{n}^{h}|<1$ and $D_{h}(1)_{2}$ exists, then $B_{h}(1)_{2}\neq 0$, and
+Theorem 3.3 ($\exists\,U(1)_{2}$)
+Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. If $\sum_{n\geq 1}|h(B_{n})|<1$ and $D_{h}(1)_{2}$ exists, then $B_{h}(1)_{2}\neq 0$ and
 $A_{h}(1)_{2}=\Big(1-\frac{1}{B_{h}(1)_{2}}\Big)D_{h}(1)_{2}\,.$
 Proof. By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}\neq 0$ and
-$(1-C_{h}(1))_{2}=\frac{1}{B_{h}(1)_{2}},\,\text{ so that }\,C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$}
+$(1-C_{h})(1)_{2}=\frac{1}{B_{h}(1)_{2}},\,\text{ so that }\,C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$}
 By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{h}(1)_{2}=C_{h}(1)_{2}D_{h}(1)_{2}$ . The stated formula follows. $\Box$
-Theorem 3.4 (3: $v=1$, gen, $U(1)_{1}=\infty$)
-If $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a light weight such that $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
+Theorem 3.4 ($U(1)_{1}=\infty$)
+Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a light weight. If $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
 $(A_{h})_{1/\xi}(1)_{1}=+\infty\,.$
 Proof.
-From $(D_{h})_{1/\xi}(1)_{1}=+\infty$ it follows that $\xi^{-n}d_{n}^{h}>0$ for some $n\in\mathbb{N}$.
-Hence there is a vertex $u\in\mathbb{N}\setminus\{1\}$ with $\xi^{-1}h(\{1,u\})>0$.
-The walk $\langle 1,u,1\rangle$ shows by the subseries bound principle that $\xi^{-2}c_{2}^{h}>0$.
-Since by item 1 of Proposition 3.1 we have $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}d_{n-j}^{h}$, we get that
-$\xi^{-2-n}a_{n+2}^{h}\geq\xi^{-2}c_{2}\cdot\xi^{-n}d_{n}^{h}$
+From $(D_{h})_{1/\xi}(1)_{1}=+\infty$ it follows that $\xi^{-n}h(D_{n})>0$ for some $n$ in $\mathbb{N}$.
+Hence there is a vertex $u\in\mathbb{N}\setminus\{1\}$ such that $\xi^{-1}h(\{1,u\})>0$.
+The walk $\langle 1,u,1\rangle$ shows by the subseries bound that $\xi^{-2}h(C_{2})>0$.
+Since by item 1 of Proposition 3.1 we have $h(A_{n})=\sum_{j=0}^{n}h(C_{j})h(D_{n-j})$, we get that
+$\xi^{-2-n}h(A_{n+2})\geq\xi^{-2}h(C_{2})\cdot\xi^{-n}h(D_{n})$
 for every $n\in\mathbb{N}_{0}$. Hence
-$\sum_{j=0}^{n}\xi^{-2-j}a_{j+2}^{h}\geq\xi^{-2}c_{2}^{h}\cdot\sum_{j=0}^{n}\xi^{-j}d_{j}^{h}\to+\infty$}
+$\sum_{j=0}^{n}\xi^{-2-j}h(A_{j+2})\geq\xi^{-2}h(C_{2})\cdot\sum_{j=0}^{n}\xi^{-j}h(D_{j})\to+\infty$}
 for $n\to\infty$, and $(A_{h})_{1/\xi}(1)_{1}=+\infty$. $\Box$
-We proceed to $\xi$-convex weights.
-Theorem 3.5 (4: $v=1$, con, $U(1)_{1}$)
+3.3 Pólya’s theorems for $v=1$ and $\xi$-convex weights
+We proceed to $\xi$-convex weights and begin with twisted nonnegative ones. The next theorem generalizes Section 1.2.
+Theorem 3.5 ($\exists\,U(1)_{1}$ and $h\geq 0$)
+Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $\xi$-convex light weight. Then, with $(B_{h})_{1/\xi}(1)_{1}\in[1,+\infty)\cup\{+\infty\}$, we have
+$\lim_{n\to\infty}\xi^{-n}h(A_{n})=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}\,.$
+For $(B_{h})_{1/\xi}(1)_{1}=+\infty$ the limit is $1$, and else it lies in $(0,1)$.
+Proof.
+Proposition 2.31 gives $D_{h}(x)=\frac{1}{1-\xi x}$.
+Since $B_{n},C_{n}\subset D_{n}$ for every $n\in\mathbb{N}_{0}$, by the subseries bound we have $\xi^{-n}h(B_{n}),\xi^{-n}h(C_{n})\in[0,1]$.
+Now we argue as in Section 1.2.
+The last claim that $(B_{h})_{1/\xi}(1)_{1}>1$ follows from the assumption of $\xi$-convexity: there is a $u\in\mathbb{N}\setminus\{1\}$ such that $\xi^{-1}h(\{1,u\})>0$, and hence $\xi^{-2}h(B_{2})>0$.
+$\Box$
+Theorem 3.6 ($\exists\,U(1)_{1}$)
 1.
-If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $\xi$-convex light weight and if $(B_{h})_{1/\xi}(1)_{1}$ and $(C_{h})_{1/\xi}(1)_{1}$ exist, then $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}\,.$
-2. If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a 0-convex light weight, and if $B_{h}(1)_{1}$ and $C_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
+Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $\xi$-convex light weight.
+If $(B_{h})_{1/\xi}(1)_{1}$ and $(C_{h})_{1/\xi}(1)_{1}$ exist, then $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
+$\lim_{n\to\infty}\xi^{-n}h(A_{n})=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}\,.$
+2. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a 0-convex light weight. If $B_{h}(1)_{1}$ and $C_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
 $A_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}\,.$
 Proof. 1. By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}\,.$
+$\lim_{n\to\infty}\xi^{-n}h(A_{n})=(C_{h})_{1/\xi}(1)_{1}\,.$
 By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By Proposition 2.21 we have $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
@@
 The stated formula follows.
-2. Now, by item 1 of Proposition 3.1 and by Proposition 2.31, we have
-$A_{h}(x)=C_{h}(x)$
-and argue as in the previous case for $\xi=1$. $\Box$
-This theorem generalizes the case $d\geq 3$ of Pólya’s theorem with $\overline{v}=\overline{0}$. The case $d\leq 2$ is generalized in Theorem 3.7.
-Theorem 3.6 (5: $v=1$, con, $U(1)_{2}$)
-1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $\xi$-convex light weight and $\sum_{n\geq 1}\big|b_{n}^{h}\big|<1$, then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$
-2. If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a 0-convex light weight and $\sum_{n\geq 1}\big|b_{n}^{h}\big|<1$, then $B_{h}(1)_{2}\neq 0$ and
+2. By item 1 of Proposition 3.1 and by Proposition 2.31, $A_{h}(x)=C_{h}(x)$. We use the displayed formula for $(C_{h})_{1/\xi}(1)_{1}$ with $\xi=1$. $\Box$
+This theorem generalizes the case $d\geq 3$ of the formula (1) for $\overline{v}=\overline{0}$. The case $d\leq 2$ is generalized in Theorem 3.8.
+Theorem 3.7 ($\exists\,U(1)_{2}$)
+1. Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $\xi$-convex light weight. If $\sum_{n\geq 1}\big|h(B_{n})\big|<1$, then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
+$\lim_{n\to\infty}\xi^{-n}h(A_{n})=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$
+2. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a 0-convex light weight. If $\sum_{n\geq 1}|h(B_{n})|<1$, then $B_{h}(1)_{2}\neq 0$ and
 $A_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$
@@
 $\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}=(C_{h})_{1/\xi}(1)_{2}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
-2. In the previous case we have for $\xi=1$ that $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. By item 1 of Proposition 3.1 and by Proposition 2.31, $A_{h}(x)=C_{h}(x)$. The stated formula follows. $\Box$
-Theorem 3.7 (6: $v=1$, con, $U(1)_{1}=\infty$)
-1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $\xi$-convex light weight and if $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1\,.$
-2. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ is a 0-convex light weight, $B_{h},C_{h}\in H_{1}$, and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
-$\lim_{x\nearrow 1}F_{(A_{h})_{1/\xi}}(x)=1\,.$
-Proof. 1. By item 1 of Proposition 3.1 and by Proposition 2.31 we have
-$A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}\,.$}
-By Proposition 2.15,
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}\,.$
-By Proposition 2.31 and the subseries bound principle,
-$0\leq\xi^{-n}b_{n}^{h},\,\xi^{-n}c_{n}^{h}\leq\xi^{-n}d_{n}^{h}=1$
-for every $n\in\mathbb{N}_{0}$. Hence $B_{h},C_{h}\in H_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. Thus
-$(C_{h})_{1/\xi}(1)_{1}=\lim_{x\nearrow 1}F_{(C_{h})_{1/\xi}}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1\,,}$
-where in the first and third equality we used Theorem 2.18. The stated formula follows.
-2. By item 1 of Proposition 3.1 and by Proposition 2.31, we have $A_{h}(x)=C_{h}(x)$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. Thus
-$\lim_{x\nearrow 1}F_{(C_{h})_{1/\xi}}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1-\frac{1}{\pm\infty}=1\,,}$
-where in the second equality we used Theorem 2.17. The stated formula follows at once. $\Box$
-4 Extensions of Pólya’s theorem with $v\neq 1$
-In Theorems 4.34.10 we establish the second group of generalizations of Pólya’s theorem, in the more complicated case when the vertex to be visited by walks differs from their starting vertex.
-The grid graph $\mathbb{Z}^{d}$ is vertex-transitive, which has so far been irrelevant.
-For $\overline{v}\neq\overline{0}$, vertex-transitivity becomes relevant.
-We generalize it to $K_{\mathbb{N}}$ as follows.
-Let $v\in\mathbb{N}\setminus\{1\}$.
-We say that a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $v$-transitive if there is a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that $f(1)=v$ and $h(f[e])=h(e)$ for every $e\in\mathbb{N}_{2}$.
-We again introduce some generating functions.
-Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight.
-Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the subset of walks visiting the vertex $v$ ($\neq 1$); necessarily in a step $i>0$.
-We are interested in the generating function
-$A_{h,v}(x)=\sum_{n\geq 0}a_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W(v,\,n))x^{n}\,.$}
-In the previous section we introduced the generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$. Now we define
-$B_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{b}(v,\,n))x^{n}\,,}$
-$C_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{c}(v,\,n))x^{n}\,\text{ and}}$
-$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e}(v,\,n))x^{n}\,,}$
-all in $\mathbb{C}[[x]]$.
-Here $W_{b}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$.
-$W_{c}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding $v$.
-Finally, $W_{e}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding both $1$ and $v$.
-The constant coefficients are $a_{0}^{h,v}=c_{0}^{h,v}=c_{0}^{h}=e_{0}^{h,v}=0$ and $b_{0}^{h}=b_{0}^{h,v}=d_{0}^{h}=1$.
-We obtain an analog of Proposition 3.1. Since the proof is similar, we omit details and argue only on the level of semiformal method.
+2. By item 1 of Proposition 3.1 and by Proposition 2.31, $A_{h}(x)=C_{h}(x)$. We use the displayed formula for $(C_{h})_{1/\xi}(1)_{2}$ with $\xi=1$. $\Box$
+Theorem 3.8 ($U(1)_{1}=\infty$)
+1. Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $\xi$-convex light weight. If $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
+$\lim_{n\to\infty}\xi^{-n}h(A_{n})=1\,.$
+2. Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ be a 0-convex light weight. If $B_{h}(x),C_{h}(x)\in H_{1}$ and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
+$\lim_{x\nearrow 1}(A_{h})_{1/\xi}(x)=1\,.$
+Proof. 1. We proved this in Theorem 3.5.
+2. By item 1 of Proposition 3.1 and by Proposition 2.31, we have $A_{h}(x)=C_{h}(x)$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By Theorem 2.17,
+$\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)=(B_{h})_{1/\xi}(1)_{1}=\pm\infty\,.$
+It follows from this and the assumptions that
+$\text{$(C_{h})_{1/\xi}(x)=1-\frac{1}{(B_{h})_{1/\xi}(x)}$ for every $x\in(1-\delta,\,1)$}\,.$}
+We compute that
+$\lim_{x\nearrow 1}(C_{h})_{1/\xi}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1-\frac{1}{\pm\infty}=1\,.}$
+In the second equality we used Theorem 2.17. The stated formula follows. $\Box$
+4 General Pólya’s theorems for $v\neq 1$
+In this last section, we generalize the results in Section 1.3 to weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$.
+We first introduce another three auxiliary generating functions $B_{v,h}(x)$, $C_{v,h}(x)$, and $E_{v,h}(x)$.
+The second identity in Proposition 4.2 contains a square, and in Theorem 4.4 we deduce a result on monotonicity from this.
+In Sections 4.2 and 4.3, we obtain general Pólya’s theorems for $v\neq 1$ with general and $\xi$-convex weights, respectively.
+4.1 Auxiliary generating functions
+Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. As we know from Section 1.4, we are interested in the main generating function
+$A_{v,\,h}(x)=\sum_{n\geq 0}h(A_{v,\,n})x^{n}\ \ (\in\mathbb{C}[[x]])$}
+for the set $A_{v}$ of walks in $K_{\mathbb{N}}$ starting at $1$ and visiting $v$; $h(A_{v,n})$ is the sum of the series
+$\sum_{w\in A_{v,\,n}}h(w)=\sum_{\begin{subarray}{c}w\in A_{v}\\ s(w)=n\end{subarray}}h(w)\,.$}
+Besides the sets $A_{v}$, $B$, and $D$ of walks in $K_{\mathbb{N}}$, we use the set $B_{v}$ of walks in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$, the set $C_{v}$ of walks in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and with the inner vertices different from $v$, and the set $E_{v}$ of walks in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and with the inner vertices different from both $1$ and $v$.
+We define three additional auxiliary generating functions
+$B_{v,\,h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}h(B_{v,\,n})x^{n},\ C_{v,\,h}(x)=\sum_{n\geq 0}h(C_{v,\,n})x^{n}\,\text{ and}}$
+$E_{v,\,h(x)}$ $=$ ${\textstyle\sum_{n\geq 0}h(E_{v,\,n})x^{n}\,.}$
+We have $h(A_{v,0})=h(C_{v,0})=h(E_{v,0})=0$ and $h(B_{0})=h(B_{v,0})=h(D_{0})=1$.
+In the case $\overline{v}=\overline{0}$ the symmetry of the grid graph $G_{d}=\langle\mathbb{Z}^{d},E_{d}\rangle$, namely its vertex-transitivity, played no role, but in the case $\overline{v}\neq\overline{0}$ this symmetry is crucial; it ensures that the enumeration of walks does not depend on the starting vertex.
+The same holds for the general weighted case, and we introduce the weighted version of vertex transitivity.
+Definition 4.1 ($v$-transitivity)
+Let $v\in\mathbb{N}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We say that the weight $h$ is $v$-transitive if there exists a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that
+$f(1)=v\,\text{ and }\,\forall e\in\mathbb{N}_{2}:\;h(e)=h(f[e])\,.$
+We obtain an analog of Proposition 3.1. The proofs are similar, and we only sketch them.
 Proposition 4.2
-Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ be the above generating functions.
-The following relations hold.
+Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight, and $A_{v,h}(x)$, $B_{h}(x)$, $B_{v,h}(x)$, $C_{v,h}(x)$, $D_{h}(x)$, and $E_{v,h}(x)$ be the above generating functions.
+The following identities hold in $\mathbb{C}[[x]]$.
   1. 1.
-    $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$.
+    $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)$.
   2. 2.
-    $B_{h}(x)=B_{h,v}(x)+C_{h,v}(x)^{2}B_{h}(x)$, equivalently, $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
+    $B_{h}(x)=B_{v,h}(x)+C_{v,h}(x)^{2}B_{h}(x)$, equivalently, $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}$.
   3. 3.
-    $C_{h,v}(x)=B_{h,v}(x)E_{h,v}(x)$.
-Proof.
-1.
-Walks $w$ weight-counted by $A_{h,v}(x)$ split at the first visit of $v$ as $w=w_{1}w_{2}$.
-Walks $w_{1}$ are weight-counted by $C_{h,v}(x)$ and $w_{2}$ are arbitrary walks starting at $v$.
-By the $v$-transitivity of $h$, walks $w_{2}$ are weight-counted by $D_{h}(x)$.
-2.
-It suffices to prove the first equality.
-$B_{h}(x)$ weight-counts walks $w$ starting and ending at $1$.
-Those avoiding $v$ are weight-counted by $B_{h,v}(x)$.
-If $w$ visits $v$, it uniquely splits at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$.
-Walks $w_{1}$ start at $1$, end at $v$ and between avoid $v$.
-Walks $w_{2}$ start and end at $v$.
-Walks $w_{3}$ start at $v$, end at $1$ and between avoid $v$.
-We reverse walks $w_{3}$ and see that both $w_{1}$ and $w_{3}$ are weight-counted by $C_{h,v}(x)$.
-Due to the $v$-transitivity of $h$, the middle walks $w_{2}$ are weight-counted by $B_{h}(x)$.
-3.
-We consider walks $w$ weight-counted by $C_{h,v}(x)$.
-They start at $1$, end at $v$, and between avoid $v$.
-These walks uniquely split at the last visit of $1$ as $w=w_{1}w_{2}$.
-Walks $w_{1}$ are weight-counted by $B_{h,v}(x)$, and $w_{2}$ by $E_{h,v}(x)$.
-$\Box$
+    $C_{v,h}(x)=B_{v,h}(x)E_{v,h}(x)$.
+Proof. 1. We split every walk $w\in A_{v}$ at the first visit of $v$ as $w=w_{1}w_{2}$. This splitting yields, when the $v$-transitivity of $h$ is applied to $w_{2}$, the stated identity.
+2. It suffices to prove the first identity. We have the partition
+$B=B_{v}\cup B_{v}^{\prime}$
+where the set $B_{v}^{\prime}$ consists of the walks in $K_{\mathbb{N}}$ that start and end at $1$ and visit $v$.
+We split every walk $w\in B_{v}^{\prime}$ at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$.
+The above partition, the reversal of walks (applied to $w_{3}$), and the $v$-transitivity of $h$ (applied to $w_{2}$) yield the stated identity.
+3. We split every walk $w\in C_{v}$ at the last visit of $1$ as $w=w_{1}w_{2}$. This splitting yields, without help of the $v$-transitivity of $h$, the stated identity. $\Box$
+One would expect to see for $v\neq 1$ the generating function of the set $F_{v}$ of walks in $K_{\mathbb{N}}$ going from $1$ to $v$.
+We managed for a long time without it, but at last we include two identities involving $F_{v}$.
+So we define another auxiliary generating function
+$F_{v,\,h}(x)=\sum_{n\geq 0}h(F_{v,\,n})x^{n}\,.$}
+Proposition 4.3
+Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight, and $B_{h}(x)$, $B_{v,h}(x)$, $C_{v,h}(x)$, and $F_{v,h}(x)$ be the above generating functions.
+The following identities hold in $\mathbb{C}[[x]]$.
+  1. 1.
+    $F_{v,h}(x)=C_{v,h}(x)B_{h}(x)$.
+  2. 2.
+    $B_{v,h}(x)=B_{h}(x)-\frac{F_{v,h}(x)^{2}}{B_{h}(x)}$.
+Proof. 1. We split every walk $w\in F_{v}$ at the first visit of $v$ as $w=w_{1}w_{2}$. This splitting yields, when the $v$-transitivity of $h$ is applied to $w_{2}$, the identity.
+$F_{v,\,h}(x)=C_{v,\,h}(x)+C_{v,\,h}(x)(B_{h}(x)-1)=C_{v,\,h}(x)B_{h}(x)\,.$
+2. Using the reversal of walks and the $v$-transitivity of $h$, we get
+$B_{h}(x)-\frac{F_{v,\,h}(x)^{2}}{B_{h}(x)}=B_{h}(x)-F_{v,\,h}(x)^{2}\sum_{j\geq 1}(-1)^{j-1}(B_{h}(x)-1)^{j-1}=B_{v,\,h}(x)\,.$}
+Indeed, if $w$ is a walk in $K_{\mathbb{N}}$ from $1$ to $1$ with $k\in\mathbb{N}_{0}$ visits of the vertex $v$, then $w$ is counted exactly
+$1-\sum_{j=1}^{k}\binom{k}{j}(-1)^{j-1}=\sum_{j=0}^{k}\binom{k}{j}(-1)^{j}=(1-1)^{k}\ \ (\in\{0,\,1\})$}
+times in the left-hand side $\cdots$ of the displayed equality $\dots=B_{v,h}(x)$, which agrees with the count of $w$ in the right-hand side $B_{v,h}(x)$. $\Box$
+Item 2 in the previous proposition and item 1 in this one provide two expressions for $C_{v,h}(x)$:
+$C_{v,h}(x)^{2}=1-\frac{B_{v,\,h}(x)}{B_{h}(x)}\,\text{ and }\,C_{v,h}(x)=\frac{F_{v,\,h}(x)}{B_{h}(x)}\,.$}
+Although the latter expression is simpler than the former, both in appearance and factually ($F_{v}$ has simpler definition than $B_{v}$), we use the former: due to inclusions $B_{v,n}\subset B_{n}$ we easily show the membership in $(0,1)$ in Theorem 4.9.
+In the simple latter expression we have disjoint sets $F_{v,n}$ and $B_{n}$, and it is not clear how to deduce that $(F_{v,h})_{1/\xi}(1)_{1}<(B_{h})_{1/\xi}(1)_{1}$.
 Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $v\in\mathbb{N}\setminus\{1\}$.
-If $v\not\in V(h)$ then $a_{n}^{h,v}=\sum_{w\in W(v,n)}h(w)=\sum_{w\in W(v,n)}0=0$ for every $n\in\mathbb{N}_{0}$ because every walk starting at $1$ and visiting $v$ contains an edge with zero weight.
-Henceforth, we therefore assume that $v\in V(h)$.
-For $w,z\in\mathbb{C}$ we define the set
+If $v\not\in V(h)$ then $h(w)=0$ for every walk $w\in A_{v}$ because every walk starting at $1$ and visiting $v$ contains an edge with zero weight.
+Henceforth, we therefore always assume that $v\in V(h)$.
+For $w,z\in\mathbb{C}$, we define the set
 $w\cdot\mathrm{sqrt}(z)=\{w\alpha\colon\;\alpha\in\mathbb{C},\,\alpha^{2}=z\}\ \ (\subset\mathbb{C})\,.$
 For $w=0$ or $z=0$ it equals $\{0\}$. Else it has two non-zero elements differing only by sign. If $w=1$, we write just $\mathrm{sqrt}(z)$.
-Before we turn to generalizations of Pólya’s theorem in the case $v\neq 1$, we utilize the square in item 2 of Proposition 4.1.
-If $h\colon X\to\mathbb{R}$ is an absolutely convergent real series and $Y\subset X$, then in general the sum $\sum_{x\in Y}h(x)$ of the subseries need not be smaller than the sum $\sum_{x\in X}h(x)$.
-However, due to the mentioned square, we obtain the next monotonicity result for the sums
-$b_{n}^{h,v}=\sum_{w\in W_{b}(v,\,n)}h(w)\,\text{ and }\,b_{n}^{h}=\sum_{w\in W_{b}(n)}h(w)\,.$}
-It is expressed in terms of the generating functions $B_{h,v}(x)=\sum_{n\geq 0}b_{n}^{h,v}x^{n}$ and $B_{h}(x)=\sum_{n\geq 0}b_{n}^{h}x^{n}$.
+Before we proceed to general Pólya’s theorems for $v\neq 1$, we utilize the square in item 2 of Proposition 4.2.
+If $h\colon X\to\mathbb{R}$ is an absolutely convergent real series and $Y\subset X$, then the sum $h(Y)$ of the subseries need not be smaller than the sum $h(X)$.
+However, in the case of inclusions $B_{v,n}\subset B_{n}$, sums $h(B_{v,n})$ are, in a sense, indeed smaller than sums $h(B_{n})$, as shown in the following theorem.
 Theorem 4.4
-Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{R}$ be a $v$-transitive light weight and let $B_{h,v},B_{h},C_{h,v}\in H_{1}$.
-Then for every real $c\in(-1,1)$ the following holds.
+Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{R}$ be a $v$-transitive light weight and let $B_{v,h}(x),B_{h}(x),C_{v,h}(x)\in H_{1}$.
+Then for every number $x\in(-1,1)$ the following holds.
   1. 1.
-    If $B_{h}(c)=0$ then also $B_{h,v}(c)=0$.
+    If $B_{h}(x)=0$ then also $B_{v,h}(x)=0$.
   2. 2.
-    If $B_{h}(c)\neq 0$ then
-    $\frac{B_{h,v}(c)}{B_{h}(c)}\leq 1\,.$
-Proof. This follows from relations 2 in Proposition 4.1. $\Box$
-We proceed to the remaining generalizations of Pólya’s theorem.
-Theorem 4.3 (7: $v\neq 1$, gen, $U(1)_{1}$)
-If $v$ is in $\mathbb{N}\setminus\{1\}$, if $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$, and if $A_{h,v}(1)_{1}$, $B_{h}(1)_{1}\neq 0$, $B_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then
-$A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,.$
+    If $B_{h}(x)\neq 0$ then
+    $\frac{B_{v,\,h}(x)}{B_{h}(x)}\leq 1\,.$
+Proof. By item 2 of Proposition 4.2, the assumptions, and Propositions 2.2, 2.4 and 2.6 we have for every number $x\in(-1,1)$ that
+$B_{h}(x)=B_{v,h}(x)+C_{v,h}(x)^{2}B_{h}(x)\,.$
+The claim 1 is clear. If $B_{h}(x)\neq 0$ then
+$\frac{B_{v,\,h}(x)}{B_{h}(x)}=1-C_{v,\,h}(x)^{2}\leq 1\,.$}
+$\Box$
+4.2 Pólya’s theorems for $v\neq 1$ and general weights
+We obtain general Pólya’s theorems of the form indicated in the subsection heading.
+Theorem 4.5 ($\exists\,U(1)_{1}$)
+Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$.
+If $A_{v,h}(1)_{1}$, $B_{h}(1)_{1}$, $B_{v,h}(1)_{1}$, $C_{v,h}(1)_{1}$ and $D_{h}(1)_{1}$ exist and $B_{h}(1)_{1}\neq 0$, then
+$A_{v,\,h}(1)_{1}\in D_{h}(1)_{1}\cdot\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{1}}{B_{h}(1)_{1}}\Big)\,.$
 Proof.
-By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$.
-By Proposition 2.20, $A_{h,v}(1)_{1}=C_{h,v}(1)_{1}D_{h}(1)_{1}$.
-By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
-Thus by Proposition 2.22, $\big(C_{h,v}(1)_{1}\big)^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$.
+By item 1 of Proposition 4.2 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)$.
+By Proposition 2.20, $A_{v,h}(1)_{1}=C_{v,h}(1)_{1}D_{h}(1)_{1}$.
+By item 2 of Proposition 4.2 we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}$.
+Thus by Proposition 2.22, $\big(C_{v,h}(1)_{1}\big)^{2}=1-\frac{B_{v,h}(1)_{1}}{B_{h}(1)_{1}}$.
 The stated formula follows.
 $\Box$
-If we replace the assumption of non-vanishing of $B_{h}(1)_{1}$ with the existence of $E_{h,v}(1)_{1}$, we obtain the following variant of the theorem.
-Theorem 4.4 (7: $v\neq 1$, gen, $U(1)_{1}$)
-If $v$ is in $\mathbb{N}\setminus\{1\}$, if $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$, and if $A_{h,v}(1)_{1}$, $B_{h}(1)_{1}$, $B_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$, $D_{h}(1)_{1}$ and $E_{h,v}(1)_{1}$ exist, then
-$A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,$,
-where for $B_{h}(1)_{1}=0$ we have $B_{h,v}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then $A_{h,v}(1)_{1}=0$.
+We trade the non-vanishing of $B_{h}(1)_{1}$ for the existence of $E_{v,h}(1)_{1}$ and obtain the following variant of the last theorem.
+Theorem 4.6 ($\exists\,U(1)_{1}$)
+Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$.
+If $A_{v,h}(1)_{1}$, $B_{h}(1)_{1}$, $B_{v,h}(1)_{1}$, $C_{v,h}(1)_{1}$, $D_{h}(1)_{1}$ and $E_{v,h}(1)_{1}$ exist, then
+$A_{v,\,h}(1)_{1}\in D_{h}(1)_{1}\cdot\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{1}}{B_{h}(1)_{1}}\Big)\,$,
+where for $B_{h}(1)_{1}=0$ we have $B_{v,h}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then $A_{v,h}(1)_{1}=0$.
 Proof.
 If $B_{h}(1)_{1}\neq 0$, we use the previous proof.
-If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1, and Propositions 2.20 and 2.19 that also $B_{h,v}(1)_{1}=0$.
-Item 3 of Proposition 4.1 and Proposition 2.20 give $C_{h,v}(1)_{1}=0$.
-Finally, item 1 of Proposition 4.1 and Proposition 2.20 give $A_{h,v}(1)_{1}=0$.
+If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.2, and Propositions 2.20 and 2.19 that also $B_{v,h}(1)_{1}=0$.
+Item 3 of Proposition 4.2 and Proposition 2.20 give $C_{v,h}(1)_{1}=0$.
+Finally, item 1 of Proposition 4.2 and Proposition 2.20 give $A_{v,h}(1)_{1}=0$.
 $\Box$
-Theorem 4.5 (8: $v\neq 1$, gen, $U(1)_{2}$)
-Suppose that $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$,
-$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
-and that $B_{h,v}(1)_{2}$ and $D_{h}(1)_{2}$ exist.
-If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$.
-For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$.
-If
-$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
+Theorem 4.7 ($\exists\,U(1)_{2}$)
+Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. If
+$c=\sum_{n\geq 1}|h(B_{n})|<1$
+and $B_{v,h}(1)_{2}$ and $D_{h}(1)_{2}$ exist, then the following holds.
+If $B_{v,h}(x)=B_{h}(x)$ then $A_{v,h}(x)=0$.
+If $B_{v,h}(x)\neq B_{h}(x)$, $k\in\mathbb{N}_{0}$ is minimum such that $h(B_{v,k})\neq h(B_{k})$ and if
+$\big(|h(B_{v,\,k})-h(B_{k})|\cdot(1-c)\big)^{-1}\sum_{n>k}|h(B_{v,\,n})-h(B_{n})|<1\,$,
 then $B_{h}(1)_{2}\neq 0$ and
-$A_{h,\,v}(1)_{2}\in{\textstyle D_{h}(1)_{2}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
+$A_{v,\,h}(1)_{2}\in D_{h}(1)_{2}\cdot\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{2}}{B_{h}(1)_{2}}\Big)\,.$
 Proof.
-If $B_{h,v}(x)=B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=1-1=0$ and $C_{h,v}(x)=0$.
-Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
-Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold.
-By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
+If $B_{v,h}(x)=B_{h}(x)$, then by item 2 of Proposition 4.2 we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}=1-1=0$ and $C_{v,h}(x)=0$.
+Item 1 of Proposition 4.2 gives $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=0$.
+Let $B_{v,h}(x)\neq B_{h}(x)$, $k$ be as stated, and let the latter displayed bound hold.
+By item 2 of Proposition 4.2, we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}$.
 By Proposition 2.30, $B_{h}(1)_{2}\neq 0$ and
-$(C_{h,v}(1)_{2})^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}\,.$}
-By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{h,v}(1)_{2}=C_{h,v}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
-Theorem 4.6 (9: $v\neq 1$, gen, $U(1)_{1}=\infty$)
-If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $v$-transitive light weight with $v\in V(h)$ and if $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
-$(A_{h,v})_{1/\xi}(1)_{1}=+\infty\,.$
+$\big(C_{v,\,h}(1)_{2}\big)^{2}=1-\frac{B_{v,\,h}(1)_{2}}{B_{h}(1)_{2}}\,.$}
+By item 1 of Proposition 4.2 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{v,h}(1)_{2}=C_{v,h}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
+Theorem 4.8 ($U(1)_{1}=\infty$)
+Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $v$-transitive light weight such that $v\in V(h)$.
+If $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
+$(A_{v,\,h})_{1/\xi}(1)_{1}=+\infty\,.$
 Proof.
-We denote the minimum length of a walk joining $1$ and $v$ in $G(h)$ by $m$.
-The inner vertices of it differ from $v$.
-By the subseries bound principle we get $\xi^{-m}c^{h,v}_{m}>0$.
-By item 1 of Proposition 4.1 we have $a_{n}^{h,v}=\sum_{j=0}^{n}c_{j}^{h,v}d_{n-j}^{h}$ and we conclude that $\xi^{-n-m}a_{n+m}^{h,v}\geq\xi^{-m}c_{m}^{h,v}\cdot\xi^{-n}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$.
-Hence
-$\sum_{j=0}^{n}\xi^{-j-m}a_{j+m}^{h,v}\geq\xi^{-m}c_{m}^{h,v}\cdot\sum_{j=0}^{n}\xi^{-j}d_{j}^{h}\to+\infty$}
-for $n\to\infty$, and $(A_{h,v})_{1/\xi}(1)_{1}=+\infty$. $\Box$
-We proceed to $\xi$-convex weights.
-Theorem 4.7 (10: $v\neq 1$, con, $U(1)_{1}$)
+Let $m$ be the minimum length of a walk $w$ joining $1$ and $v$ in $G(h)$.
+Then $\xi^{-m}h(w)>0$ and $w\in C_{v,m}$.
+Using the subseries bound we get $\xi^{-m}h(C_{v,m})>0$.
+By item 1 of Proposition 4.2 we have $h(A_{v,n})=\sum_{j=0}^{n}h(C_{v,j})h(D_{n-j})$ and we conclude that
+$\xi^{-n-m}h(A_{v,\,n+m})\geq\xi^{-m}h(C_{v,\,m})\cdot\xi^{-n}h(D_{n})$
+for every $n\in\mathbb{N}_{0}$. Hence
+$\sum_{j=0}^{n}\xi^{-j-m}h(A_{v,\,j+m})\geq\xi^{-m}h(C_{v,\,m})\cdot\sum_{j=0}^{n}\xi^{-j}h(D_{j})\to+\infty$}
+for $n\to\infty$. So $(A_{v,h})_{1/\xi}(1)_{1}=+\infty$. $\Box$
+4.3 Pólya’s theorems for $v\neq 1$ and $\xi$-convex weights
+We proceed to $\xi$-convex weights and begin with twisted nonnegative ones. The next theorem generalizes Section 1.3.
+Theorem 4.9 ($\exists\,U(1)_{1}$ and $h\geq 0$)
+Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$.
+If $(B_{h})_{1/\xi}(1)_{1}$ exists, then
+$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=\sqrt{1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}}\in(0,\,1)\,.$
+Proof.
+By Proposition 2.31, $D_{h}(x)=\frac{1}{1-\xi x}$.
+Since $B_{n},B_{v,n},C_{v,n}\subset D_{n}$ for every $n\in\mathbb{N}_{0}$, we see by the subseries bound that $(B_{h})_{1/\xi}(x)$, $(B_{v,h})_{1/\xi}(x)$ and $(C_{v,h})_{1/\xi}(x)$ have all coefficients in $[0,1]$ and are in $H_{1}$.
+Since $B_{v,n}\subset B_{n}$ for every $n\in\mathbb{N}_{0}$, we see by the subseries bound that
+$0\leq(B_{v,\,h})_{1/\xi}(1)_{1}\leq(B_{h})_{1/\xi}(1)_{1}<+\infty\,.$
+By item 2 of Proposition 4.2 we have $(C_{v,h})_{1/\xi}(x)^{2}=1-\frac{(B_{v,h})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}$.
+Using the arithmetic of functional limits, Proposition 2.20 (more precisely, the version for nonnegative coefficients), the bound $(B_{h})_{1/\xi}(1)_{1}\geq 1>0$ and Theorem 2.18, we compute that
+$\big((C_{v,\,h})_{1/\xi}(1)_{1}\big)^{2}=((C_{v,\,h})_{1/\xi})^{2}(1)_{1}$ $=$ $\lim_{x\nearrow 1}((C_{v,\,h})_{1/\xi})^{2}(x)$
+$=$ ${\textstyle 1-\frac{\lim_{x\nearrow 1}(B_{v,\,h})_{1/\xi}(x)}{\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)}}$
+$=$ ${\textstyle 1-\frac{(B_{v,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\ \ (\in[0,\,1])\,.}$
+We used the third Abelian theorem on the first and third lines.
+Thus, due to nonnegativity, $(C_{v,\,h})_{1/\xi}(1)_{1}$ equals to the displayed root.
+By item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)\frac{1}{1-\xi x}$.
+Proposition 2.15 gives that
+$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=(C_{v,\,h})_{1/\xi}(1)_{1}$
+and the stated formula follows. The membership to $(0,1)$ follows from the inequality
+$1\leq(B_{v,\,h})_{1/\xi}(1)_{1}<(B_{h})_{1/\xi}(1)_{1}\,.$
+The strict inequality is witnessed by any walk in $G(h)$ from $1$ to $v$, which yields (by going back via the reversal) a walk in $G(h)$ from $1$ to $1$ visiting $v$. $\Box$
+Theorem 4.9 covers only the finite case $(B_{h})_{1/\xi}(1)_{1}<+\infty$; we move the case $(B_{h})_{1/\xi}(1)_{1}=+\infty$ to part 1 of Theorem 4.13.
+Theorem 4.10 ($\exists\,U(1)_{1}$)
 1.
-If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h,v})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and $(C_{h,v})_{1/\xi}(1)_{1}$ exist, then
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\big)\,.$}
+Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$.
+If $(B_{v,h})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}$ and $(C_{v,h})_{1/\xi}(1)_{1}$ exist and $(B_{h})_{1/\xi}(1)_{1}\neq 0$, then
+$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})\in\mathrm{sqrt}\Big(1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\Big)\,.$
 2.
-If $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, and if $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}\neq 0$ and $C_{h,v}(1)_{1}$ exist, then
-$A_{h,v}(1)_{1}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}\big)\,.$}
-Proof. 1. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}\,.$
-By item 2 of Proposition 4.1 we have $(C_{h,v})_{1/\xi}(x)^{2}=1-\frac{(B_{h,v})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}$.
-By Proposition 2.22, $\big((C_{h,v})_{1/\xi}(1)_{1}\big)^{2}=1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}$.
+Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive 0-convex light weight such that $v\in V(h)$.
+If $B_{v,h}(1)_{1}$, $B_{h}(1)_{1}$ and $C_{v,h}(1)_{1}$ exist and $B_{h}(1)_{1}\neq 0$, then
+$A_{v,\,h}(1)_{1}\in\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{1}}{B_{h}(1)_{1}}\Big)\,.$
+Proof. 1. By item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
+$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=(C_{v,\,h})_{1/\xi}(1)_{1}\,.$
+By item 2 of Proposition 4.2 we have $(C_{v,h})_{1/\xi}(x)^{2}=1-\frac{(B_{v,h})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}$. By Proposition 2.22,
+$\big((C_{v,\,h})_{1/\xi}(1)_{1}\big)^{2}=1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\,.$}
 The stated formula follows.
-2. Now, by item 1 of Proposition 4.1 and by Proposition 2.31, we have
-$A_{h,v}(x)=C_{h,v}(x)\,.$
-The rest of the argument is the same as in the previous case for $\xi=1$. $\Box$
-Again, if we replace the assumption of non-vanishing of $B_{h,v}(1)_{1}$ with the existence of $E_{h,v}(1)_{1}$, we obtain the following variant of the theorem.
-Theorem 4.8 (10: $v\neq 1$, con, $U(1)_{1}$)
+2. By item 1 of Proposition 4.2 and by Proposition 2.31, $A_{v,h}(x)=C_{v,h}(x)$. We use the displayed formula for $(C_{v,h})_{1/\xi}(1)_{1}$ with $\xi=1$. $\Box$
+We again trade the non-vanishing of $B_{h}(1)_{1}$ for the existence of $E_{v,h}(1)_{1}$ and obtain the following variant of the last theorem.
+Theorem 4.11 ($\exists\,U(1)_{1}$)
 1.
-If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h,v})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}$, $(C_{h,v})_{1/\xi}(1)_{1}$ and $(E_{h,v})_{1/\xi}(1)_{1}$ exist, then
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\big)\,$,}
-where for $(B_{h})_{1/\xi}(1)_{1}=0$ we have $(B_{h,v})_{1/\xi}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then the limit is zero.
+Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$.
+If $(B_{v,h})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}$, $(C_{v,h})_{1/\xi}(1)_{1}$ and $(E_{v,h})_{1/\xi}(1)_{1}$ exist, then
+$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})\in\mathrm{sqrt}\Big(1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\Big)\,$,
+where for $(B_{h})_{1/\xi}(1)_{1}=0$ we have $(B_{v,h})_{1/\xi}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then the limit is zero.
 2.
-If $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, and if $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h,v}(1)_{1}$ and $E_{h,v}(1)_{1}$ exist, then
-$A_{h,v}(1)_{1}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}\big)\,$,}
-where for $B_{h}(1)_{1}=0$ we have $B_{h,v}(1)_{1}=0$ and interpret the $\frac{0}{0}$ as $1$, so that then $A_{h,v}(1)_{1}=0$.
+Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive 0-convex light weight such that $v\in V(h)$.
+If $B_{v,h}(1)_{1}$, $B_{h}(1)_{1}$, $C_{v,h}(1)_{1}$ and $E_{v,h}(1)_{1}$ exist, then
+$A_{v,\,h}(1)_{1}\in\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{1}}{B_{h}(1)_{1}}\Big)\,$,
+where for $B_{h}(1)_{1}=0$ we have $B_{v,h}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then $A_{v,h}(1)_{1}=0$.
+Proof. If $(B_{h})_{1/\xi}(1)_{1}\neq 0$, resp. $B_{h}(1)_{1}\neq 0$, we use the previous proofs. We therefore assume that $(B_{h})_{1/\xi}(1)_{1}=0$, resp. $B_{h}(1)_{1}=0$.
+1. From the proof of part 1 in Theorem 4.10 we know that
+$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=(C_{v,\,h})_{1/\xi}(1)_{1}\,.$
+Using item 2 of Proposition 4.2 and Propositions 2.20 and 2.19, it follows from $(B_{h})_{1/\xi}(1)_{1}=0$ that $(B_{v,h})_{1/\xi}(1)_{1}=0$.
+Item 3 of Proposition 4.2 and Proposition 2.20 give $(C_{v,h})_{1/\xi}(1)_{1}=0$.
+So the displayed limit is 0.
+2.
+Since $\xi=0$, item 1 of Proposition 4.2 and Proposition 2.31 give that $A_{v,h}(x)=C_{v,h}(x)$.
+As we know, $B_{h}(1)_{1}=0$ implies that $B_{v,h}(1)_{1}=0$ and that $C_{v,h}(1)_{1}=0$.
+So $A_{v,h}(1)_{1}=0$.
+$\Box$
+The last two theorems generalize the case $d\geq 3$ of the formula (1) for $\overline{v}\neq\overline{0}$. The case $d\leq 2$ is generalized in Theorem 4.13.
+Theorem 4.12 ($\exists\,U(1)_{2}$)
+1. Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$. If
+$c=\sum_{n\geq 1}|h(B_{n})|<1$
+and if $(B_{v,h})_{1/\xi}(1)_{2}$ exists, then the following holds.
+If $B_{v,h}(x)=B_{h}(x)$ then $A_{v,h}(x)=0$.
+If $B_{v,h}(x)\neq B_{h}(x)$, $k\in\mathbb{N}_{0}$ is minimum such that $h(B_{v,k})\neq h(B_{k})$ and if
+$\big(|h(B_{v,\,k})-h(B_{k})|\cdot(1-c)\big)^{-1}\sum_{n>k}|h(B_{v,\,n})-h(B_{n})|<1\,$,
+then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
+$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})\in\mathrm{sqrt}\Big(1-\frac{(B_{v,\,h})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\Big)\,.$
+2. Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive 0-convex light weight such that $v\in V(h)$. If
+$c=\sum_{n\geq 1}|h(B_{n})|<1$
+and if $B_{v,h}(1)_{2}$ exists, then the following holds.
+If $B_{v,h}(x)=B_{h}(x)$ then $A_{v,h}(x)=0$.
+If $B_{v,h}(x)\neq B_{h}(x)$, $k\in\mathbb{N}_{0}$ is minimum such that $h(B_{v,k})\neq h(B_{k})$ and if
+$\big(|h(B_{v,\,k})-h(B_{k})|\cdot(1-c)\big)^{-1}\sum_{n>k}|h(B_{v,\,n})-h(B_{n})|<1\,$,
+then $B_{h}(1)_{2}\neq 0$ and
+$A_{v,\,h}(1)_{2}\in\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{2}}{B_{h}(1)_{2}}\Big)\,.$
 Proof.
+Let $B_{v,h}(x)=B_{h}(x)$.
+By item 2 of Proposition 4.2 we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}=1-1=0$ and $C_{v,h}(x)=0$.
+Item 1 of Proposition 4.2 gives $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=0$.
 1.
-If $(B_{h})_{1/\xi}(1)_{1}\neq 0$, we use the previous proof.
-If $(B_{h})_{1/\xi}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.20 and 2.19 that also $(B_{h,v})_{1/\xi}(1)_{1}=0$.
-Item 3 of Proposition 4.1 and Proposition 2.20 give $(C_{h,v})_{\xi}(1)_{1}=0$.
-Thus, as in the previous theorem,
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}=0\,.$
-2. Now, by item 1 of Proposition 4.1 and by Proposition 2.31, we have
-$A_{h,v}(x)=C_{h,v}(x)\,.$
-The rest of the argument is the same as in the previous case for $\xi=1$. $\Box$
-These two theorems generalize the case $d\geq 3$ of Pólya’s theorem with $\overline{v}\neq\overline{0}$. The case $d\leq 2$ is generalized in Theorem 4.10.
-Theorem 4.9 (11: $v\neq 1$, con, $U(1)_{2}$)
-1. Suppose that $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$,
-$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
-and that $(B_{h,v})_{1/\xi}(1)_{2}$ exists. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
-$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
-then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}}\big)\,.$
-2. Suppose that $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$,
-$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
-and that $B_{h,v}(1)_{2}$ exists. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
-$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
-then $(B_{h})(1)_{2}\neq 0$ and
-$A_{h,v}(1)_{2}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
-Proof.
-If $B_{h,v}(x)=B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=0$ and $C_{h,v}(x)=0$.
-Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
+Let $h$ be as stated, $B_{v,h}(x)\neq B_{h}(x)$, $k$ be as stated, and let the latter displayed bound hold.
+By item 2 of Proposition 4.2, we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}$.
+By Proposition 2.30, $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
+$\big((C_{v,h})_{1/\xi}(1)_{2}\big)^{2}=1-\frac{(B_{v,h})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
+By item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)\frac{1}{1-\xi x}$. Proposition 2.15 gives the stated formula:
+$\lim_{n\to\infty}\xi^{-n}h(A_{v,n})={\textstyle(C_{v,\,h})_{1/\xi}(1)_{1}=(C_{v,\,h})_{1/\xi}(1)_{2}\in\mathrm{sqrt}\big(1-\frac{(B_{v,h})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\big)\,.}$
+2.
+Let $h$ be as stated, $B_{v,h}(x)\neq B_{h}(x)$, $k$ be as stated, and let the latter displayed bound hold.
+By part 1 for $\xi=1$ we have $\big(C_{v,h}(1)_{2}\big)^{2}=1-\frac{B_{v,h}(1)_{2}}{B_{h}(1)_{2}}$.
+By item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)$.
+The stated formula follows.
+$\Box$
+Theorem 4.13 ($U(1)_{1}=\infty$)
 1.
-Let $h$ be $\xi$-converx with $\xi\in\mathbb{C}_{1}$, $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold.
-By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
-By Proposition 2.30, $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
-$\big((C_{h,v})_{1/\xi}(1)_{2}\big)^{2}=1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
-By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$. Proposition 2.15 gives the stated formula:
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}={\textstyle(C_{h,\,v})_{1/\xi}(1)_{1}=(C_{h,\,v})_{1/\xi}(1)_{2}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\big)\,.}$
+Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$.
+If $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
+$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=1\,.$
 2.
-Let $h$ be 0-convex, $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold.
-The beginning of the previous case gives for $\xi=1$ that $\big(C_{h,v}(1)_{2}\big)^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$.
-Since by item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)$, the stated formula follows.
+Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ be a $v$-transitive 0-convex light weight such that $v\in V(h)$.
+If $B_{h}(x),C_{v,h}(x)\in H_{1}$, $(B_{v,h})_{1/\xi}(1)_{1}$ exists and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
+$\lim_{x\nearrow 1}(A_{v,\,h})_{1/\xi}(x)\in\{-1,\,1\}\,.$
+Proof. 1. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-\xi x}$. By the subseries bound,
+$0\leq\xi^{-n}h(B_{n}),\,\xi^{-n}h(B_{v,\,n}),\,\xi^{-n}h(C_{v,\,n}),\,\xi^{-n}h(E_{v,\,n})\leq\xi^{-n}h(D_{n})=1$
+for every $n\in\mathbb{N}_{0}$.
+Hence $B_{h}(x)$, $B_{v,h}(x)$, $C_{v,h}(x)$, and $E_{v,h}(x)$ are in $H_{1}$.
+Any shortest walk joining $1$ and $v$ in $G(h)$ shows that for some $m\in\mathbb{N}$ we have $\xi^{-m}h(E_{v,m})>0$.
+Thus the function $(E_{v,h})_{1/\xi}(x)$ is positive and increasing on $(0,1)$.
+Since $(C_{v,h})_{1/\xi}(x)$ is bounded on $[0,1)$ by item 2 of Proposition 4.2, $(B_{v,h})_{1/\xi}(x)$ is bounded by item 3.
+Since by item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)\frac{1}{1-\xi x}$, Proposition 2.15 gives
+$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=(C_{v,\,h})_{1/\xi}(1)_{1}=(C_{v,\,h})_{1/\xi}(1)_{2}\,.$
+By item 2 of Proposition 4.2, by the arithmetic of functional limits, and by Theorem 2.18, we have
+$((C_{v,\,h})_{1/\xi})^{2}(1)_{1}=\lim_{x\nearrow 1}(C_{v,\,h})_{1/\xi}(x)^{2}=1-\frac{\lim_{x\nearrow 1}(B_{v,\,h})_{1/\xi}(x)}{\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)}=1\,$,}
+because the function $(B_{v,h})_{1/\xi}(x)$ is bounded on $[0,1)$ and $\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)=B_{h}(1)_{1}=+\infty$.
+Hence, by nonnegativity, $(C_{v,h})_{1/\xi}(1)_{1}=(C_{v,h})_{1/\xi}(1)_{2}=1$ because
+$\big((C_{v,\,h})_{1/\xi}(1)_{2}\big)^{2}=\big((C_{v,\,h})_{1/\xi}\big)^{2}(1)_{2}$
+by (the version for nonnegative coefficients of) Proposition 2.20. The stated formula follows.
+2. By item 1 of Proposition 4.2 and by Proposition 2.31, $A_{v,h}(x)=C_{v,h}(x)$. By item 2 of Proposition 4.2 and by the arithmetic of functional limits,
+$\lim_{x\nearrow 1}(C_{v,\,h})_{1/\xi}(x)^{2}=1-\frac{\lim_{x\nearrow 1}(B_{v,\,h})_{1/\xi}(x)}{\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)}=1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{\pm\infty}=1-0=1\,.$}
+In the second equality we used Theorems 2.16 (in the numerator) and 2.17 (in the denominator).
+Since the involved real functions are continuous on $[0,1)$, the limit $\lim_{x\nearrow 1}(A_{v,h})_{1/\xi}(x)$ exists and
+$\lim_{x\nearrow 1}(A_{v,\,h})_{1/\xi}(x)=\lim_{x\nearrow 1}(C_{v,\,h})_{1/\xi}(x)\in\{-1,\,1\}\,.$
 $\Box$
-Theorem 4.10 (12: $v\neq 1$, con, $U(1)_{1}=\infty$)
-1.
-If $\xi\in\mathbb{C}_{1}$, $v$ is in $\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=1\,.$
-2.
-If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, $B_{h},C_{h,v}\in H_{1}$, $(B_{h,v})_{1/\xi}(1)_{1}$ exists and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
-$\lim_{x\nearrow 1}F_{(A_{h,v})_{1/\xi}}(x)\in\{-1,\,1\}\,.$
-Proof. 1. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-\xi x}$. By the subseries bound principle we have
-$0\leq\xi^{-n}b_{n}^{h},\,\xi^{-n}b_{n}^{h,v},\,\xi^{-n}c_{n}^{h,v},\,\xi^{-n}e_{n}^{h,v}\leq\xi^{-n}d_{n}^{h}=1$
-for every $n\in\mathbb{N}_{0}$.
-Hence $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$ and $E_{h,v}(x)$ are in $H_{1}$.
-Any shortest walk joining $1$ and $v$ in $G(h)$ shows that for some $m\in\mathbb{N}$ we have $\xi^{-m}e_{m}^{h,v}>0$.
-Thus the function $F_{(E_{h,v})_{1/\xi}}(x)$ is positive and increasing on $(0,1)$.
-Since $F_{(C_{h,v})_{1/\xi}}(x)$ is bounded on $[0,1)$ by item 2 of Proposition 4.1, $F_{(B_{h,v})_{1/\xi}}(x)$ is bounded by item 3.
-Since by item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$, Proposition 2.15 gives
-$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}=(C_{h,v})_{1/\xi}(1)_{2}\,.$
-By item 2 of Proposition 4.1, Propositions 2.2, 2.4 and 2.5, and by Theorem 2.18 we have
-$(C_{h,v}^{2})_{1/\xi}(1)_{1}=\lim_{x\nearrow 1}F_{(C_{h,v}^{2})_{1/\xi}}(x)=1-\frac{\lim_{x\nearrow 1}F_{(B_{h,v})_{1/\xi}}(x)}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1\,$,}
-because $F_{(B_{h,v})_{1/\xi}}(x)$ is bounded on $[0,1)$ and $\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)=B_{h}(1)_{1}=+\infty$. By Proposition 2.25, $(C_{h,v})_{1/\xi}(1)_{2}$ exists and
-$\big((C_{h,v})_{1/\xi}(1)_{2}\big)^{2}=\big((C_{h,v})_{1/\xi}\big)^{2}(1)_{2}=(C_{h,v}^{2})_{1/\xi}(1)_{2}=(C_{h,v}^{2})_{1/\xi}(1)_{1}=1\,.$
-Since $(C_{h,v})_{1/\xi}(1)_{2}\geq 0$, we have $(C_{h,v})_{1/\xi}(1)_{2}=1$ and the stated formula follows.
-2. By item 1 of Proposition 4.1 and by Proposition 2.31, $A_{h,v}(x)=C_{h,v}(x)$. By item 2 of Proposition 4.1 and by Theorems 2.16 and 2.17,
-$\lim_{x\nearrow 1}F_{(C_{h,v}^{2})_{1/\xi}}(x)=1-\frac{\lim_{x\nearrow 1}F_{(B_{h,v})_{1/\xi}}(x)}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{\pm\infty}=1\,.$}
-By the continuity of the functions $F_{\dots}$,
-$\lim_{x\nearrow 1}F_{(A_{h,v})_{1/\xi}}(x)=\lim_{x\nearrow 1}F_{(C_{h,v})_{1/\xi}}(x)\in\{-1,\,1\}\,.$
-$\Box$
-5 Concluding remarks
-In [17] we consider the same problem of determining formulas for quantities such as $A_{h,v}(1)=\sum_{n\geq 0}h(W(v,n))$ in the case of formal weights $h$. These are maps
-$h\colon\mathbb{N}_{2}\to\mathbb{C}[[x_{1},\,x_{2},\,\dots,\,x_{k}]]\,$,
-where the last integral domain of formal power series in $k$ variables is endowed with the usual non-Archimedean norm.
-The main novelty of our generalization is this. In equalities like
-$P_{d}(\overline{v},\,n)=h(W(v,\,n))$
-that was stated in the Introduction, the left-hand side involves, in the standard rigorous probabilistic treatment, an implicit probability space supporting the corresponding Markov chain.
-Construction of this probability space is far from trivial; see, for example, [6].
-In our generalization, the right-hand side is completely self-contained.
+We feel that part 2 of the theorem is not completely satisfactory.
+Perhaps the assumption that $(B_{v,h})_{1/\xi}(1)_{1}$ exists may be dropped, and it is true that if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then always
+$\lim_{x\nearrow 1}\frac{(B_{v,\,h})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}=0\,.$}
+5 Overview of general Pólya’s theorems
+Thm type assumption conclusion
+3.2 $v=1,{}_{1}$ $A(1),B(1),C(1),D(1)$ $A(1)=(1-\frac{1}{B(1)})D(1)$
+3.3 $v=1,{}_{2}$ $\sum|h(B_{n})|<1,D(1)$ $A(1)=(1-\frac{1}{B(1)})D(1)$
+3.4 $v=1,\infty$ $\xi\mathbb{R}_{\geq 0},D(1)=+\infty$ $A(1)=+\infty$
+3.5 $v=1,{}_{1}$, $\xi$-c. $\xi\mathbb{R}_{\geq 0}$
+$\lim h(A_{n})=1-\frac{1}{B(1)}\in$
+$\in(0,1]$
+3.6.1 $v=1,{}_{1}$, $\xi$-c. $B(1),C(1)$ $\lim h(A_{n})=1-\frac{1}{B(1)}$
+3.6.2 $v=1,{}_{1}$, 0-c. $B(1),C(1)$ $A(1)=1-\frac{1}{B(1)}$
+3.7.1 $v=1,{}_{2}$, $\xi$-c. $\sum|h(B_{n})|<1$ $\lim h(A_{n})=1-\frac{1}{B(1)}$
+3.7.2 $v=1,{}_{2}$, 0-c. $\sum|h(B_{n})|<1$ $A(1)=1-\frac{1}{B(1)}$
+3.8.1 $v=1,\infty$, $\xi$-c. $\xi\mathbb{R}_{\geq 0},B(1)=+\infty$ $\lim h(A_{n})=1$
+3.8.2 $v=1,\infty$, 0-c.
+$\xi\mathbb{R},B(x),C(x)\in H_{1}$,
+$B(1)=+\infty$
+$\lim A(x)=1$
+4.5 $v\neq 1,{}_{1}$
+$A(1),B(1)\neq 0,B_{v}(1)$,
+$C_{v}(1),D(1)$
+$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}D(1)$
+4.6 $v\neq 1,{}_{1}$
+$A(1),B(1),B_{v}(1)$,
+$C_{v}(1),D(1),E(1)$
+$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}D(1)$
+4.7 $v\neq 1,{}_{2}$
+$B_{v}(1),D(1)$,
+$\sum|h(B_{n})|,\,\sum<1$
+$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}D(1)$
+4.8 $v\neq 1,\infty$ $\xi\mathbb{R}_{\geq 0},D(1)=+\infty$ $A(1)=+\infty$
+4.9 $v\neq 1,{}_{1}$, $\xi$-c. $\xi\mathbb{R}_{\geq 0},B(1)$
+$\lim h(A_{n})=$
+$(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}\in(0,1)$
+4.10.1 $v\neq 1,{}_{1}$, $\xi$-c. $B(1)\neq 0,B_{v}(1),C_{v}(1)$ $\lim h(A_{n})\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
+4.10.2 $v\neq 1,{}_{1}$, 0-c. $B(1)\neq 0,B_{v}(1),C_{v}(1)$ $A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
+4.11.1 $v\neq 1,{}_{1}$, $\xi$-c.
+$B(1),B_{v}(1),C_{v}(1)$,
+$E(1)$
+$\lim h(A_{n})\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
+4.11.2 $v\neq 1,{}_{1}$, 0-c.
+$B(1),B_{v}(1),C_{v}(1)$,
+$E(1)$
+$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
+4.12.1 $v\neq 1,{}_{2}$, $\xi$-c.
+$B_{v}(1),\sum h(B_{n})<1$,
+$\sum<1$
+$\lim h(A_{n})\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
+4.12.2 $v\neq 1,{}_{2}$, 0-c.
+$B_{v}(1),\sum h(B_{n})<1$,
+$\sum<1$
+$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
+4.13.1 $v\neq 1,\infty$, $\xi$-c. $\xi\mathbb{R}_{\geq 0},B(1)=+\infty$ $\lim h(A_{n})=1$
+4.13.2 $v\neq 1,\infty$, 0-c.
+$\xi\mathbb{R},B(x),C_{v}(x)\in H_{1}$,
+$B_{v}(1),B(1)=\pm\infty$
+$\lim A(x)\in\{-1,1\}$
+The first column lists the theorems.
+The second column lists the theorem types, with 1 and 2 abbreviating $\exists\,U(1)_{1}$ and $\exists\,U(1)_{2}$, respectively.
+By $\infty$ we abbreviate that $U(1)_{1}=\infty$.
+The third and fourth columns list the assumptions and conclusions of the theorems.
+We always omit indices h, 1/ξ, n→∞ and x↗1, and factor $\xi^{-n}$.
+We omit indices 1, 2, and v whenever they can be deduced from the context.
+By writing $A(1)$ in the third column and first row, we mean the assumption that $A_{h}(1)_{1}$ exists, and similarly in other cases.
+By writing $\xi\mathbb{R}_{\geq 0}$ and $\xi\mathbb{R}$ we mean the assumptions $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$, respectively; else $h\colon\mathbb{N}_{2}\to\mathbb{C}$.
+$\sum|h(B_{n})|$ abbreviates $\sum_{n\geq 1}|h(B_{n})|$, and $\sum<1$ abbreviates the complicated bound assumed in the respective theorem.